### v2

# Convergence and Optimality of the EM Algorithm Under Multi-Component Gaussian Mixture Models

Abstract
Gaussian mixture models (GMMs) are fundamental statistical tools for modeling heterogeneous data. Due to the nonconcavity of the likelihood function, the Expectation-Maximization (EM) algorithm is widely used for parameter estimation of each Gaussian component. Existing analyses of the EM algorithm’s convergence to the true parameter focus on either the two-component case or multi-component settings with known mixing probabilities and isotropic covariance matrices.
In this work, we study the convergence of the EM algorithm for multi-component GMMs in full generality. The population-level EM is shown to converge to the true parameter when the smallest separation among all pairs of Gaussian components exceeds a logarithmic factor of the largest separation and the reciprocal of the minimal mixing probabilities. At the sample level, the EM algorithm is shown to be minimax rate-optimal, up to a logarithmic factor. We develop two distinct novel analytical approaches, each tailored to a different regime of separation, reflecting two complementary perspectives on the use of EM. As a byproduct of our analysis, we show that the EM algorithm, when used for community detection, also achieves the minimax optimal rate of misclustering error under milder separation conditions than spectral clustering and Lloyd’s algorithm, an interesting result in its own right. Our analysis allows the number of components, the minimal mixing probabilities, the separation between Gaussian components and the dimension to grow with the sample size. Simulation studies corroborate our theoretical findings.
Keywords: Community detection; Clustering; EM algorithm; Gaussian mixture models; Minimax optimal estimation.
1 Introduction
Finite mixture models are fundamental statistical tools for modeling heterogeneous data [lindsay1995mixture, mclachlan1988mixture] and have found wide-ranging applications in fields such as biology, economics, image processing, and natural language processing, to name just a few. These models assume that the observed data arise from a mixture of several latent sub-populations, each governed by its own distribution [mclachlan2000finite, titterington1985statistical]. Among them, the Gaussian mixture model stands out as one of the most extensively studied and applied, due to its mathematical tractability and empirical success in sub-population inference, clustering, density estimation, and outlier detection [dempster1977maximum, redner1984mixture].
In this paper, we focus on the $d$-dimensional Gaussian location mixture model with $L\geq 2$ components, hereafter referred to as the $L$-GMM. Specifically, let $X_{1},\ldots,X_{n}$ be i.i.d. samples of a random vector $X\in\mathbb{R}^{d}$ that follows, for some latent label $Y\in[L]:=\{1,\ldots,L\}$,
$X\mid Y=\ell~\sim~\mathcal{N}_{d}\left(\mu_{\ell},\Sigma\right),$ (1.1)
with $\mathbb{P}(Y=\ell)=\pi_{\ell}$ for each $\ell\in[L]$. The Gaussian components have distinct mean vectors $\mu_{1},\ldots,\mu_{L}\in\mathbb{R}^{d}$, and share a common covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$, which is assumed to be strictly positive definite. The mixing probabilities $\pi=(\pi_{1},\ldots,\pi_{L})^{\top}$ are all positive and sum up equal to one. By collecting the mean vectors into the $d\times L$ matrix $M:=(\mu_{1},\ldots,\mu_{L}),$ our main interest lies in estimating the parameter $\theta:=(\pi,M,\Sigma)$ based on $X_{1},\ldots,X_{n}$.
In the long history of Gaussian mixture models, the identifiability of $\theta$, up to label switching, under model (1.1) is well-known [teicher1963identifiability, yakowitz1968identifiability]. See also [titterington1985statistical] and [lindsay1995mixture], as well as the recent [allman2009identifiability] for general finite mixture models. Regarding estimation, a significant line of work analyzes the maximum likelihood estimation (MLE) for density estimation [ghosal2001entropies] and for estimating the mixing measure, $\rho=\sum_{\ell=1}^{L}\pi_{\ell}\delta_{\{\mu_{\ell},\Sigma\}},$ with $\delta$ being the Dirac measure under the exactly specified (in terms of the number of fitted mixture components) and over-specified settings [chen1995optimal, ho2016convergence, ho2016strong]. However, since the likelihood function under GMMs is non-concave, computing the MLE is challenging, especially in multivariate settings. To overcome this challenge, three main types of alternative approaches have been developed and studied.
  • Method of Moments (MoM)-based approaches. MoM estimation of $\theta$ has been developed in [LindsayBasak, hsu2013learning, belkin2010polynomial, hardt2015tight, moitra2010settling, kalai2010efficiently]. Although these methods require minimal separation between Gaussian components, they typically yield slow convergence rates and suboptimal sample complexity with respect to both the dimension $d$ and the number of components $L$. Recently, [WuYang2020, doss2023optimal] studied MoM-based estimation of the mixing measure and showed that it achieves the minimax optimal rate under the Wasserstein distance for bounded $L$. While their rates for estimating the mixing measure could be translated into rates for estimating $\theta$ with improved sample complexity, the resulting convergence is still exponentially slow in $L$, even in the univariate case, and corresponding results in the multivariate setting remain largely unknown.
  • Clustering-based approaches. Another substantial body of literature focuses on clustering $X_{1},\ldots,X_{n}$ by predicting their labels $Y_{1},\ldots,Y_{n}$ in model (1.1), a task also known as the community detection problem [dasgupta1999learning, sanjeev2001learning, vempala2002spectral, kannan2008spectral, awasthi2012improved, achlioptas2005spectral, lu2016statistical, chen2024achieving]. Once the labels are known, the parameter $\theta$ can be readily estimated from the labeled data. However, as we detail in Section 4.4, such methods require strong separation conditions between Gaussian components to achieve (nearly) perfect clustering.
  • The Expectation Maximization (EM) algorithm. The EM algorithm is designed in a broader context to approximate the MLE by iteratively maximizing a lower bound of the likelihood function through alternating between the expectation step (E-step) and the maximization step (M-step) [dempster1977maximum]. Under the $L$-GMM, both steps have closed-form expressions, leading to fast computation. In contrast to its empirical success, EM is only known to converge to a local stationary point in general, which may be far from the true parameter [wu1983convergence, redner1984mixture]. Over the past two decades, significant progress has been made in the theoretical understanding of EM’s convergence to the true parameter under the $L$-GMM.
The general positioning of the EM algorithm lies between the first two approaches: compared to MoM-based methods, EM estimation, when suitably initialized, typically requires more separation but yields faster and, oftentimes minimax optimal, rates of convergence; whereas compared to clustering-based methods, EM requires weaker separation condition to achieve optimal convergence rates. The primary goal of this work is to conduct a thorough theoretical analysis of the EM algorithm to formally demonstrate its adaptivity under the $L$-GMM in the presence of mild separation conditions. As explained below, existing analyses of EM are largely limited either to the case $L=2$ or to settings where the mixing probabilities $\pi$ are known and the covariance matrix $\Sigma$ are isotropic (see the summary in Table 1). Analyzing EM for mixtures with $L\geq 3$ presents significant challenges, as the likelihood landscape deteriorates with increasing $L$. Furthermore, allowing $\pi$ and $\Sigma$ to be unknown requires not only estimating these quantities but also adjusting the estimation of the mean vectors, as the E-steps (see Section 2) depend on their estimates as well. This work is the first to establish both the convergence of EM to the true parameter and its minimax optimality under the $L$-GMM with $L\geq 2$ and both $\pi$ and $\Sigma$ unknown.
1.1 Related literature
In this section, we review existing theoretical results on the convergence of the EM algorithm under Gaussian mixture models. The early work by [xu1996convergence] establishes a connection between EM iterates and the gradient of the likelihood function under general GMMs. Later, [dasgupta2007probabilistic] analyzed a two-step EM procedure equipped with a specially designed initialization scheme, under GMMs with $L\geq 2$ components and isotropic covariance matrices. Under a strong separation condition between Gaussian components, the two-step EM estimator is proved to converge to within a small neighborhood of the true parameter, although the size of this neighborhood may not shrink to zero as the sample size increases.
Recently, there has been renewed progress in understanding the convergence of EM to the true parameter under the $L$-GMM. In [EM2017], the authors develop a general framework for analyzing EM and apply it to the symmetric, isotropic 2-GMM with $\pi_{1}=\pi_{2}=1/2$, $\mu_{1}=-\mu_{2}=\mu$, and $\Sigma=\mathbf{I}_{d}$. They prove that the EM iterates $\widehat{\mu}^{(t)}$ converge to the true parameter $\mu$ in $\ell_{2}$-norm at the rate $\|\mu\|_{2}^{3}\sqrt{d/n}$, provided that the initialization $\widehat{\mu}^{(0)}$ satisfies $\|\widehat{\mu}^{(0)}-\mu\|_{2}\leq\|\mu\|_{2}/4$, and the separation between Gaussian components, $\|\mu\|_{2}$, exceeds a sufficiently large constant. Further improvements in this symmetric, isotropic $2$-GMM are made in [Xu_Hsu_Maleki] and [daskalakis2017ten], where the population level EM is shown to converge to the true parameter as long as the separation $\|\mu\|_{2}>0$ and the initialization $\widehat{\mu}^{(0)}$ is not orthogonal to $\mu$. In other words, the population level likelihood function in this special case has no spurious local maxima, so that any random initialization suffices for the sample level EM algorithm to converge to the ground truth as $n\to\infty$. More recently, still under the setting $\mu_{1}=-\mu_{2}=\mu$ and $\Sigma=\mathbf{I}_{d}$, it has been shown that the EM iterates with simple random initialization converge to the MLE and achieve adaptive minimax optimality for $\|\mu\|_{2}=\mathcal{O}(1)$, with equal mixing weights $\pi_{1}=\pi_{2}=1/2$ [wu2021randomly] and for unequal but known weights [weinberger2022algorithm]. Similar guarantees are also obtained for the case where $L=2$ is specified but the true model consists of a single Gaussian component [dwivedi2020singularity]. These strong global convergence results, however, rely on the assumptions of known $\pi$ and $\Sigma=\mathbf{I}_{d}$. When these conditions are not met, EM must also estimate both $\pi$ and $\Sigma$, and even for $L=2$, it remains unclear whether global convergence guarantees are possible. For the 2-GMM with both unknown $\pi$ and unknown $\Sigma$, [cai2019chime] established convergence of the EM algorithm to the true parameter, provided that a suitable initialization condition is met and the separation between Gaussian components is sufficiently large but remains bounded.
Despite of these works on the $2$-GMM, they can not extend to the $L$-GMM with $L\geq 3$ from several aspects. First, the landscape of the likelihood function for $L\geq 3$ differs drastically from the two-component case. As shown in [jin2016local], the population-level likelihood function under the $L$-GMM with $L\geq 3$ can exhibit poor local maxima even in the simplest case of equally weighted mixtures of well-separated, spherical Gaussians. Moreover, the EM algorithm, when initialized randomly from the data, converges with high probability to such bad local maxima. It is therefore crucial to establish, even at the population level, conditions under which EM converges to the true parameter. Specifically, one must characterize: (1) the region in which EM contracts; (2) the type of initialization that ensures the iterates remain within this region; and (3) how the required separation between Gaussian components depends on $L$, the minimal mixing proportion $\pi_{\min}=\min_{\ell}\pi_{\ell}$ as well as the dimension $d$. Second, these issues must also be addressed at the sample level. Third, once we allow the aforementioned quantities to diverge as $n\to\infty$, both the minimax optimal rate for estimating $\theta$, and whether EM provably achieves this rate, remain open problems.
For such reasons, much fewer results exist on the convergence of the EM algorithm for $L\geq 3$. When $\pi$ is known and $\Sigma=\mathbf{I}_{d}$, yan2017convergence derive the convergence rate for a gradient-based EM algorithm, while Zhao2020 analyze the standard EM. Their results require that $\min_{k\neq\ell}\|\mu_{k}-\mu_{\ell}\|_{2}$ satisfy a certain separation condition and that the initialization lies within a neighborhood of the true parameter. This separation condition is subsequently relaxed in [kwon2020algorithm, regev2017learning], and the rate analyses are refined by [segol2021improved]. However, none of these works apply when both $\pi$ and $\Sigma$ are unknown. Moreover, the convergence rates established in the aforementioned studies are not minimax-optimal, even under known $\pi$ and $\Sigma$, and can be improved in several respects. Table 1 summarizes a technical comparison with existing results for $L\geq 2$, while a full table including the $L=2$ case is deferred to Table 3 in Appendix A.
[…]
Settings Separation $\Delta_{\min}\gtrsim$ Initialization Convergence Rates Optimality
[dasgupta2007probabilistic] $\Sigma=\sigma^{2}\mathbf{I}_{d}$ $\sqrt{d\log(\pi_{\min}^{-1})}+\log(\pi_{\min}^{-1})$ A particular initialization $d(\widehat{\pi},\pi)+d(\widehat{M},M)\lesssim\sqrt{d/(n\pi_{\min})}+\exp(-c\sqrt{d})/\pi_{\min}$ $\times$
[kwon2020algorithm] $\Sigma=\sigma^{2}\mathbf{I}_{d}$ $\log(1/\pi_{\min})$ $d(\pi^{(0)},\pi)<1,~d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ A sample-splitting version of the EM $\checkmark$
[yan2017convergence] $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $(d\wedge L)\log(\Delta_{\max}/\pi_{\min})$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)\lessapprox(\pi_{\max}/\pi_{\min})\sqrt{\Delta_{\max}}(\sqrt{d}+L^{3}\Delta_{\max})\sqrt{d/n}$ $\times$
[Zhao2020] $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $(d\wedge L)\log(\Delta_{\max}/\pi_{\min})$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)\lessapprox(\sqrt{\Delta_{\max}}/\pi_{\min})\sqrt{dL/n}$ $\times$
[segol2021improved] $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $\log(1/\pi_{\min})$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)\lessapprox(1/\pi_{\min})\sqrt{dL/n}$ $\times$
This work Full generality $\log(\Delta_{\max}/\pi_{\min})$ $d(\pi^{(0)},\pi)<1,~d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{\pi},\pi)+d(\widehat{M},M)\lessapprox\sqrt{d/(n\pi_{\min})}$ $\checkmark$
$d(\Sigma^{(0)},\Sigma)\leq c$, or via preliminary labels $d(\widehat{\Sigma},\Sigma)\lessapprox\sqrt{d/n}$
1.2 Our contributions
We now summarize our main results which include analyses of the EM algorithm at both the population level and the sample level. For the former, we establish a contraction region for EM while for the latter, our analysis is based on two complementary approaches, reflecting different perspectives on EM: parameter estimation and clustering. The first approach is more suitable when the separation is small to moderate, while the second is particularly effective when the separation is moderate to large. To distinguish between the true model parameters and their iterative estimates, we denote the true parameter by $\theta^{*}=(\pi^{*},M^{*},\Sigma^{*})$.
1.2.1 Convergence of the population level EM algorithm
We review the population level EM algorithm under model (1.1) in Section 2.1, and establish its convergence to $\theta^{*}$ in Theorem 2 (for unknown $\Sigma^{*}$) and in 3 (for known $\Sigma^{*}$) of Section 3. Denote by $\Delta_{\min}$ (resp. $\Delta_{\max}$) the smallest (resp. largest) squared Mahalanobis distance among all pairs of Gaussian components. When
$\Delta_{\min}~\gtrsim~\log(\Delta_{\max}/\Delta_{\min})+\log(1/\pi_{\min}^{*}),$ (1.2)
the population level EM iterates $\theta^{(t)}=(\pi^{(t)},M^{(t)},\Sigma^{(t)})$ with $t\geq 0$, once initialized within a neighborhood of $\theta^{*}$, converges: for all $t\geq 1$ and some contraction rate $\kappa\in(0,1)$,
$d(\pi^{(t)},\pi^{*})+d(M^{(t)},M^{*})+d(\Sigma^{(t)},\Sigma^{*})\leq\kappa^{t}\left\{d(\pi^{(0)},\pi^{*})+d(M^{(0)},M^{*})+d(\Sigma^{(0)},\Sigma^{*})\right\}.$ (1.3)
The distances between the parameters in $\theta^{(t)}$ and $\theta^{*}$ are given in Eqs. 3.2 and 3.3. The convergence in (1.3) is at least linear as $t\to\infty$ and becomes superlinear once $\Delta_{\min}$ increases, since our results show that $\kappa$ decays exponentially in $\Delta_{\min}$ (see, (3.6)). To the best of our knowledge, this is the first convergence result of the population level EM algorithm under the $L$-GMM with $L\geq 3$, unknown $\pi^{*}$ and unknown $\Sigma^{*}$. Compared to the required separation when $\Sigma^{*}=\mathbf{I}_{d}$, we only pay the price of a logarithmic factor in $\Delta_{\max}$ for estimating $\Sigma^{*}$. Regarding the initialization, Theorem 2 shows that (1.3) requires
$d(\pi^{(0)},\pi^{*})<1,\quad d(M^{(0)},M^{*})\leq c_{\mu}\sqrt{\Delta_{\min}},\quad d(\Sigma^{(0)},\Sigma^{*})\leq c_{\Sigma}$ (1.4)
for some absolute constants $0\leq 2c_{\mu}+\sqrt{2}c_{\Sigma}<\sqrt{2}-1$. The existing initialization requirements on $M^{(0)}$ are similar to ours in (1.4), although derived under known $\pi^{*}$ and $\Sigma^{*}=\mathbf{I}_{d}$. For known $\Sigma^{*}$, we have $\Sigma^{(t)}=\Sigma^{*}$ for all $t\geq 0$ in (1.3) and $c_{\Sigma}=0$ in (1.4). Moreover, our 3 states that the contraction rate $\kappa$ for known $\Sigma^{*}$ becomes faster by a factor of $\Delta_{\max}$. For unknown $\Sigma^{*}$, we note in Section 2.1 that the EM iterates $\Sigma^{(t)}$ depend on $\mathbb{E}_{\theta^{*}}[XX^{\top}]$, $\pi^{(t)}$ and $M^{(t)}$, which suggests a simple way to choose $\Sigma^{(0)}$. For such a choice, 4 of Section 3 shows that $d(\Sigma^{(0)},\Sigma^{*})\leq c_{\Sigma}$ in (1.4) is satisfied, provided that the requirements on $d(\pi^{(0)},\pi^{*})$ and $d(M^{(0)},M^{*})$ in (1.4) are tightened by a multiplicative factor of $1/\Delta_{\max}$ on the right-hand side.
To the best of our knowledge, this is the first convergence result of the population level EM algorithm under the $L$-GMM with $L\geq 3$, unknown $\pi^{*}$, and either known or unknown $\Sigma^{*}$. As mentioned earlier, for known $\pi^{*}$ and $\Sigma^{*}=\mathbf{I}_{d}$, [yan2017convergence, Zhao2020] analyze the population level EM and its variant under a separation condition that is strictly stronger than ours in (1.2), often substantially so when $L$ and $d$ are moderate to large. Their initialization requirement on $M^{(0)}$ is similar to ours in (1.4), although derived under the stronger separation condition in addition to known $\pi^{*}$ and $\Sigma^{*}=\mathbf{I}_{d}$.
At the technical level, we prove (1.3) by viewing the M-step of population level EM as an operator: for all $t\geq 0$,
$\theta^{(t)}~\mapsto~(\bm{\pi}(\theta^{(t)}),\mathbf{M}(\theta^{(t)}),\bm{\Sigma}(\theta^{(t)}))=:(\pi^{(t+1)},M^{(t+1)},\Sigma^{(t+1)})=\theta^{(t+1)}.$
Throughout this paper, we use $\bm{\pi}$, $\mathbf{M}$, and $\bm{\Sigma}$ to denote operators, in order to distinguish them from their vector or matrix counterparts $\pi$, $M$, and $\Sigma$. By further using the self-consistency of EM in Section C.1, that is, $\theta^{*}=(\bm{\pi}(\theta^{*}),\mathbf{M}(\theta^{*}),\bm{\Sigma}(\theta^{*}))$, bounding $d(\pi^{(t+1)},\pi^{*})$ boils down to controlling $d(\bm{\pi}(\theta^{(t)}),\bm{\pi}(\theta^{*}))$. This, in turn, amounts to establishing the Lipschitz continuity of the operator $\theta\mapsto\bm{\pi}(\theta)$, and similarly for $\theta\mapsto\mathbf{M}(\theta)$ and $\theta\mapsto\bm{\Sigma}(\theta)$, which constitutes the main challenge in our analysis. Our proof first re-parametrizes $\theta$ by $\omega(\theta)=(\pi,M,J(\theta))$ with $J(\theta):=\Sigma^{-1}M$ as such reparametrization alleviates the difficulty of direct manipulation of $\Sigma^{-1}$. Then in Theorem 23 of Section D.2 we establish the Lipschitz continuity of $\omega(\theta)\mapsto\bm{\pi}(\omega(\theta))$ and $\omega(\theta)\mapsto\mathbf{M}(\omega(\theta))$, by quantifying the partial derivatives of $\bm{\pi}(\cdot)$ and $\mathbf{M}(\cdot)$ with respect to $\pi$, $M$, and $J(\theta)$ uniformly for all $\theta$ satisfying (1.4) in lieu of $\theta^{(0)}$. Finally, combined with the Lipschitz property of $\theta\mapsto J(\theta)$ established in Lemma 17, the conclusion in (1.3) follows. Allowing $L\geq 3$, unknown $\pi^{*}$ and $\Sigma^{*}$, and avoiding strong separation conditions significantly complicates the analysis. We refer to the end of Section 3 for a more detailed technical discussion.
1.2.2 Convergence of the sample level EM algorithm
The sample level EM algorithm is reviewed in Section 2.2. We establish its rate of convergence in Section 4, and show that it achieves the minimax optimal rate under model (1.1) and the separation condition (1.2). Due to the dependence of $\Delta_{\min}$ on $\Delta_{\max}$ and $\pi_{\min}^{*}$ in (1.2), and hence implicitly on $L$ and $d$, the analysis becomes significantly more challenging when aiming to derive the optimal rate of convergence with respect to all these quantities. To overcome this challenge, we adopt two complementary analytical approaches in Sections 4.1 and 4.2, each tailored to different regimes depending on the magnitude of $\Delta_{\min}$ and $\Delta_{\max}$.
[…]
$d(\widehat{h}^{(t+1)},h^{*})~=~d(\widehat{\bm{h}}(\widehat{\theta}^{(t)}),\bm{h}(\theta^{*}))~\leq~d(\widehat{\bm{h}}(\widehat{\theta}^{(t)}),\widehat{\bm{h}}(\theta^{*}))+d(\widehat{\bm{h}}(\theta^{*}),\bm{h}(\theta^{*})).$ (1.5)
We emphasize that this is a different decomposition from the one used in the literature, such as [EM2017, yan2017convergence, cai2019chime], and is crucial to obtain minimax optimal rates. We refer to the beginning of Section 4.1 for detailed explanations.
In 6 of Section 4.1, we control the first term on the right-hand side by establishing the Lipschitz properties of the sample level operators $\theta\mapsto\widehat{\bm{\pi}}(\theta)$, $\theta\mapsto\widehat{\mathbf{M}}(\theta)$, and $\theta\mapsto\bm{\widehat{\Sigma}}(\theta)$, which readily yield: with probability $1-\mathcal{O}(n^{-1})$, for all $\theta$ satisfying (1.4) in lieu of $\theta^{(0)}$,
$d(\widehat{\bm{\pi}}(\theta),\widehat{\bm{\pi}}(\theta^{*}))+d(\widehat{\mathbf{M}}(\theta),\widehat{\mathbf{M}}(\theta^{*}))+d(\bm{\widehat{\Sigma}}(\theta),\bm{\widehat{\Sigma}}(\theta^{*}))\leq\kappa_{n}\left\{d(\pi,\pi^{*})+d(M,M^{*})+d(\Sigma,\Sigma^{*})\right\}.$
Its proof follows the road-map of proving (1.3). By the reparametrization $\omega(\theta)$, the key is to derive uniform convergence of certain empirical quantities of the form $\partial\widehat{\bm{\pi}}(\omega(\theta))$, $\partial\widehat{\mathbf{M}}(\omega(\theta))$ and $\partial\bm{\widehat{\Sigma}}(\omega(\theta))$, where the partial derivatives are taken with respect to $\pi$, $M$, and $J(\theta)$ separately. The established uniform convergence rate along with the population level contraction factor $\kappa$ in (1.3) determines $\kappa_{n}$. To ensure $\kappa_{n}<1$, we need
$\log{\Delta_{\max}\over\Delta_{\min}}+\log{1\over\pi_{\min}^{*}}~\lesssim~\Delta_{\min}~\lesssim~{\Delta_{\min}\over\Delta_{\max}}{n\pi_{\min}^{*}\over dL^{3/2}\log^{2}(n)}$ (1.6)
It is for this reason, this analysis is only suitable for small to moderate separation. We pause here to clarify that although our analysis significantly relaxes the upper bound restriction on $\Delta_{\max}$ compared to existing results [yan2017convergence, Zhao2020], the condition imposed in (1.6) remains inevitable, as it is rooted in directly proving the convergence of $\widehat{\theta}^{(t)}$ to $\theta^{*}$. Nevertheless, as we explain below, this dependence can be avoided by adopting a different analytical approach.
Regarding the second term $d(\widehat{\bm{h}}(\theta^{*}),\bm{h}(\theta^{*}))$ in (1.5), we derive in Theorem 5 of Section 4.1 concentration inequalities between the operators $\widehat{\bm{\pi}}$, $\widehat{\mathbf{M}}$, and $\bm{\widehat{\Sigma}}$ and their population-level counterparts, evaluated at $\theta^{*}$. By combining the bounds for both terms in (1.5), we establish in Theorem 8 of Section 4.1 that the EM iterates $\widehat{\theta}^{(t)}$, initialized with $\widehat{\theta}^{(0)}$ satisfying (1.4) in lieu of $\theta^{(0)}$, satisfy: with high probability, for all $t\geq C\log n$,
$d(\widehat{\pi}^{(t)},\pi^{*})+d(\widehat{M}^{(t)},M^{*})\lesssim\sqrt{d\log n\over n\pi_{\min}^{*}},\qquad d(\widehat{\Sigma}^{(t)},\Sigma^{*})\lesssim\sqrt{d\log n\over n}.$ (1.7)
For known $\Sigma^{*}$, Theorem 7 states that the first bound holds under a weaker condition on $\Delta_{\max}$.
In Theorem 13 of Section 4.3, we establish new minimax lower bounds for estimating $\theta^{*}$ under model (1.1) with the separation condition (1.2), thereby confirming that the bounds in (1.7) are minimax optimal up to logarithmic factors.
Optimal rate of convergence under moderate to large separation. Our second approach to analyzing the EM algorithm is more aligned with the clustering perspective, and is tailored to settings with moderate to large separation. Intuitively, when there is substantial separation between Gaussian components, it is easy to cluster data points according to their respective labels, and then estimate the model parameter from the clustered data. However, since EM is fundamentally designed to approximately maximize the likelihood for estimating $\theta^{*}$, its performance for predicting the labels $Y_{1},\ldots,Y_{n}$ remains unknown, even for the case $L=2$.
The main idea of our proof consists of two steps: in the first step we introduce a surrogate loss related to the community detection problem under model (1.1) and establish its convergence; in the second step we translate this convergence into a convergence rate for estimating $\theta^{*}$. The surrogate loss, $\phi(\theta)$ given by (4.9), characterizes the discrepancy between the posterior probabilities $\mathbb{P}_{\theta}(Y_{i}=\ell\mid X_{i})$ in the E-step at parameter value $\theta$, and the true labels $1\{Y_{i}=\ell\}$ for all $i\in[n]$ and $\ell\in[L]$. It also captures the degree of separation between the Gaussian components, thereby reflecting the benefit of increased separation. In Theorems 10 and 9 of Section 4.2, we establish the convergence of $\phi(\widehat{\theta}^{(t)})$ along the EM iterates $\widehat{\theta}^{(t)}$: with probability at least $1-4n^{-1}-\exp(-c\Delta_{\min})$ for some absolute constant $c>0$,
$\phi(\widehat{\theta}^{(t)})~\leq~n\exp(-c\Delta_{\min})+{1\over 2}\phi(\widehat{\theta}^{(t-1)}),\qquad\forall~t\geq 1$ (1.8)
for both known and unknown $\Sigma^{*}$. In addition to the initialization condition in (1.4), the above result admits an alternative requirement on $\widehat{\theta}^{(0)}$ in terms of $\phi(\widehat{\theta}^{(0)})$, which can be linked to preliminary estimates of the labels. On the other hand, unlike the contraction-based argument in our first approach, the convergence result in (1.8) requires only a lower bound on $\Delta_{\min}$ and, in fact, improves with larger values of $\Delta_{\min}$; see, details in Theorems 10 and 9.
For the estimation of $\theta^{*}$, the convergence in (1.8) further implies that the EM iterates $\widehat{\theta}^{(t)}$ themselves converge to $\widetilde{\theta}$, which is the MLE of $\theta^{*}$ under model (1.1) when the true labels are given. This, combined with a simple analysis of $\widetilde{\theta}$, leads to Theorem 11 in Section 4.2, which states that whenever $\Delta_{\min}\gtrsim\log(n)$, with probability at least $1-\mathcal{O}(n^{-1})$, the bounds in (1.7) hold for all $\widehat{\theta}^{(t)}$ with $t\gtrsim\log n$. The required sample complexity is $n\gtrsim d(L^{2}/(\pi_{\min}^{*}\Delta_{\min})^{2}+L^{2}+1/\pi_{\min}^{*})$, up to a logarithmic factor. Deriving such dependence on $d$ is nontrivial and relies on a leave-one-out analysis. We refer the reader to the discussion following Theorem 10 for further explanation of the proof technique.
Finally, by combining the results from both approaches, we have established that the EM iterates satisfy the bounds in (1.7) after $\mathcal{O}(\log n)$ iterations, and is also minimax optimal up to logarithmic factors, for all $\Delta_{\min}$ satisfying (1.2). Although our results are established for $L\geq 3$, it is worth noting that they also improve upon existing results in the case of $L=2$ with unknown $\pi^{*}$ and $\Sigma^{*}$; see, 14 and its subsequent discussion.
1.2.3 Optimality of the EM for community detection
Building on our convergence result in (1.8), we also establish the convergence rate of the EM algorithm’s misclustering error in the context of community detection. Specifically, let $\ell(\widehat{\theta}^{(t)})$ denote the proportion of misclustered labels among $Y_{1},\ldots,Y_{n}$ when using the parameter estimate $\widehat{\theta}^{(t)}$ (see (4.15) for details). Then, Theorem 15 in Section 4.4 shows that, when $\Delta_{\min}\gtrsim\log(1/\pi_{\min}^{*})+\log(\Delta_{\max}\wedge\log n)$ and initialized as above, we have, with probability at least $1-n^{-1}-\exp(-c\Delta_{\min})$, for all $t\geq 2\log n$,
$\ell(\widehat{\theta}^{(t)})~\leq~\exp(-c\Delta_{\min}).$
This rate is minimax optimal under model (1.1), up to the constant $c$. To the best of our knowledge, this is the first work to establish the rate optimality of the EM algorithm for community detection under the $L$-GMM, including the case $L=2$.
As mentioned earlier, a substantial body of literature has studied the community detection problem under model (1.1), with spectral clustering and Lloyd’s algorithm being the most well-studied methods. Compared to both approaches, the EM algorithm is rate-optimal while requiring weaker separation conditions. These differences stem from two key distinctions: (1) Spectral clustering and Lloyd’s algorithm estimate only the labels and the mean vectors $\mu_{\ell}$’s of the Gaussian components, without directly estimating the covariance matrix $\Sigma$; (2) They rely on hard assignments, assigning each data point to a single label. In contrast, EM performs soft assignments in its E-steps, assigning probabilities to all labels, which are then used to update the estimates of $\mu_{\ell}$’s and $\Sigma$ in the subsequent M-step. Our analysis shows that, while estimating $\Sigma$ requires a stronger initialization condition, it enables EM to achieve the optimal misclustering rate. Moreover, the soft-assignment nature of EM allows it to attain this optimal rate under milder separation conditions than those required by Lloyd’s algorithm. We refer the reader to Section 4.4 for a detailed discussion.
This paper is organized as follows: In Section 2 we review both the population level and sample level EM algorithms. Convergence of the population level EM is established in Section 3. We analyze the convergence of the sample level EM in Section 4.1 for small to moderate separation, and in Section 4.2 for moderate to large separation. Minimax lower bounds for parameter estimation are presented in Section 4.3, while the application of our EM analysis to quantifying its performance in community detection is discussed in Section 4.4. We give detailed discussion on practical choices of the initialization in Section 5.1 and conduct numerical studies to corroborate our theoretical findings in Section 5.2. All the proofs are deferred to the Appendix.
Notation.
For any vector $x\in\mathbb{R}^{p}$, denote its $\ell_{q}$ norm as $\|x\|_{q}$. For any $p\times p$ symmetric and invertible matrix $\Sigma$, we write $\|x\|_{\Sigma}^{2}=x^{\top}\Sigma^{-1}x$. For any matrix $A$, we write $\|A\|_{{\rm op}}$ and $\|A\|_{F}$ as its operator norm and Frobenius norm. Further write $\|A\|_{2,\infty}=\max_{j}\|A_{j\cdot}\|_{2}$. For any symmetric, positive semi-definite matrix $Q\in\mathbb{R}^{p\times p}$, we use $\lambda_{1}(Q)\geq\lambda_{2}(Q)\geq\dots\geq\lambda_{p}(Q)$ to denote its eigenvalues. We write $c\prec Q\prec C$ if $c\leq\lambda_{p}(Q)\leq\lambda_{1}(Q)\leq C$. For any two sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if there exists some constant $C$ such that $a_{n}\leq Cb_{n}$. The notation $a_{n}\asymp b_{n}$ stands for $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. For any $a,b\in\mathbb{R}$, we write $a\wedge b=\min\{a,b\}$ and $a\vee b=\max\{a,b\}$. We use $\mathbf{I}_{d}$ to denote the $d\times d$ identity matrix. Throughout this paper, we use $c,c^{\prime},C,C^{\prime}$ to denote positive and finite absolute constants that, unless otherwise indicated, can vary from line to line.
2 The EM algorithm under the $L$-GMM
In this section we review both the population level and the sample level EM algorithms under model (1.1). Recall that $\theta^{*}=(\pi^{*},M^{*},\Sigma^{*})$ denotes the true parameter.
2.1 The population level EM algorithm
Under model (1.1), the log-likelihood of $\theta$ at any $X=x\in\mathbb{R}^{d}$ is
$\log f(\theta\mid x)\\ \propto-\frac{1}{2}\log|\Sigma|+\log{\left\{\sum_{\ell=1}^{L}\pi_{\ell}\exp{\left[-\frac{1}{2}(x-\mu_{\ell})^{\top}\Sigma^{-1}(x-\mu_{\ell})\right]}\right\}}$
which is non-concave in $\theta$. The EM algorithm iteratively maximizes a lower bound of the log-likelihood function, which is known as the surrogate $Q$-function.
To state its definition, we first recall model (1.1). When $Y$ were observable, the complete log-likelihood of $\theta$ at $X=x$ and $Y=y$ is
$\log f(\theta\mid x,y)\propto-\sum_{\ell=1}^{L}1\{y=\ell\}\left\{(x-\mu_{\ell})^{\top}\Sigma^{-1}(x-\mu_{\ell})-2\log(\pi_{\ell})+\log|\Sigma|\right\},$
which is a concave function in $\theta$. Given any parameter $\theta$, the surrogate $Q$-function simply replaces the indicator $1\{y=\ell\}$ in the above display with
$\gamma_{\ell}(x;\theta)$ $:=~\mathbb{P}_{\theta}(Y=\ell\mid X=x)~=~\frac{\pi_{\ell}}{\sum_{k=1}^{L}\pi_{k}\exp\left\{(x-{1\over 2}(\mu_{\ell}+\mu_{k}))^{\top}\Sigma^{-1}(\mu_{k}-\mu_{\ell})\right\}}.$ (2.1)
Concretely, given any parameter value $\theta^{\prime}$, the population level $Q$-function is
$Q(\theta\mid\theta^{\prime})=-\mathbb{E}_{\theta^{*}}\left\{\sum_{\ell=1}^{L}\gamma_{\ell}(X;\theta^{\prime})\left[(X-\mu_{\ell})^{\top}\Sigma^{-1}(X-\mu_{\ell})-2\log(\pi_{\ell})+\log|\Sigma|\right]\right\}.$ (2.2)
For any $t\geq 0$ with the current parameter $\theta^{(t)}$, the E-step is to evaluate $\gamma_{\ell}(x;\theta^{(t)})$ in (2.1) while the M-step is to maximize $Q(\theta\mid\theta^{(t)})$ in (2.2) with respect to its first argument $\theta$. The latter turns out to have a closed-form solution: for all $\ell\in[L]$,
$\pi_{\ell}^{(t+1)}:=\bm{\pi}_{\ell}(\theta^{(t)})~=~\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta^{(t)})],$ (2.3)
$\mu_{\ell}^{(t+1)}:=\bm{\mu}_{\ell}(\theta^{(t)})~=~\frac{\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta^{(t)})X]}{\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta^{(t)})]}$ (2.4)
and, by collecting $\pi^{(t+1)}=(\pi_{1}^{(t+1)},\ldots,\pi_{L}^{(t+1)})^{\top}$ and $M^{(t+1)}=(\mu_{1}^{(t+1)},\ldots,\mu_{L}^{(t+1)})$,
$\Sigma^{(t+1)}:=\bm{\Sigma}(\theta^{(t)})$ $=\mathbb{E}_{\theta^{*}}\left[\sum_{\ell=1}^{L}\gamma_{\ell}(X;\theta^{(t)})\bigl(X-\mu_{\ell}^{(t+1)}\bigr)\bigl(X-\mu_{\ell}^{(t+1)}\bigr)^{\top}\right]$ (2.5)
$=\Sigma_{T}^{*}-M^{(t+1)}{\rm diag}(\pi^{(t+1)})M^{(t+1)\top}.$ (2.6)
Here we write $\Sigma_{T}^{*}:=\mathbb{E}_{\theta^{*}}[XX^{\top}]$ and ${\rm diag}(\pi^{(t+1)})$ denotes the diagonal matrix with diagonal elements equal to $\pi^{(t+1)}$. In Eqs. 2.3 and 2.4, we use bold fonts $\bm{\pi}_{\ell}(\cdot)$, $\bm{\mu}_{\ell}(\cdot)$, and $\bm{\Sigma}(\cdot)$ to denote the operators used in the population level EM algorithm. The updated parameter $\theta^{(t+1)}=(\pi^{(t+1)},M^{(t+1)},\Sigma^{(t+1)})$ is subsequently used to compute the E-step in the next iteration. The population level EM algorithm keeps iterating between the E-step in (2.1) and the M-step in Eqs. 2.3, 2.4 and 2.5 until convergence.
2.2 The sample level EM algorithm
When we have access to $X_{1},X_{2},\ldots,X_{n}$ from (1.1), the sample level EM maximizes the following sample level $Q$-function: for any given $\theta^{\prime}$,
$Q_{n}(\theta\mid\theta^{\prime})=-\mathbb{E}_{n}\left\{\sum_{\ell=1}^{L}\gamma_{\ell}(X;\theta^{\prime})\Bigl[(X-\mu_{\ell})^{\top}\Sigma^{-1}(X-\mu_{\ell})-2\log(\pi_{\ell})+\log|\Sigma|\Bigr]\right\}$ (2.7)
where the expectation $\mathbb{E}_{\theta^{*}}$ in (2.2) is replaced by the empirical average $\mathbb{E}_{n}$ over $X_{1},\ldots,X_{n}$. Correspondingly, for any $t\geq 0$ and given its current estimate $\widehat{\theta}^{(t)}$, the E-step at the $(t+1)$th iteration computes $\gamma_{\ell}(X_{i};\widehat{\theta}^{(t)})$ as in (2.1), for each $i\in[n]$, while the M-step maximizes $Q_{n}(\theta\mid\widehat{\theta}^{(t)})$ in (2.7) over $\theta$, yielding that: for all $\ell\in[L]$,
$\widehat{\pi}^{(t+1)}_{\ell}$ $:=~\widehat{\bm{\pi}}_{\ell}(\widehat{\theta}^{(t)})~:=~\mathbb{E}_{n}[\gamma_{\ell}(X;\widehat{\theta}^{(t)})],$ (2.8)
$\widehat{\mu}^{(t+1)}_{\ell}$ $:=~\widehat{\bm{\mu}}_{\ell}(\widehat{\theta}^{(t)})~:=~\frac{\mathbb{E}_{n}[\gamma_{\ell}(X;\widehat{\theta}^{(t)})X]}{\mathbb{E}_{n}[\gamma_{\ell}(X;\widehat{\theta}^{(t)})]}$ (2.9)
and, with $\widehat{\Sigma}_{T}=n^{-1}\sum_{i=1}^{n}X_{i}X_{i}^{\top}$,
$\widehat{\Sigma}^{(t+1)}:=\bm{\widehat{\Sigma}}(\theta^{(t)})$ $=\mathbb{E}_{n}\left[\sum_{\ell=1}^{L}\gamma_{\ell}(X;\widehat{\theta}^{(t)})(X-\widehat{\mu}^{(t+1)}_{\ell})(X-\widehat{\mu}^{(t+1)}_{\ell})^{\top}\right]$ (2.10)
$=\widehat{\Sigma}_{T}-\widehat{M}^{(t+1)}{\rm diag}(\widehat{\pi}^{(t+1)})\widehat{M}^{(t+1)\top}.$ (2.11)
[…]
One practical consideration is that computing $\gamma_{\ell}(X_{i};\widehat{\theta}^{(t)})$ in (2.1) requires to invert $\widehat{\Sigma}^{(t)}$ for each iteration $t\geq 0$. Using the expression in (2.11), the following lemma shows that $\widehat{\Sigma}^{(t)}$, for all $t\geq 0$, are invertible as long as $\widehat{\Sigma}_{T}$ is so. Its proof is stated in Appendix B.
Lemma 1.
Provided that $\widehat{\Sigma}_{T}$ is invertible, the matrices $\widehat{\Sigma}^{(t+1)}$ in (2.10) for all $t\geq 0$ are invertible.
We remark that the singularity of $\widehat{\Sigma}_{T}$ can be easily checked, and is guaranteed almost surely when $n\geq d$.
3 Convergence of the population level EM algorithm
In this section, we analyze the population level EM and establish its convergence to the true parameter. Denote the smallest and largest squared Mahalanobis distance between any two Gaussian components by
$\Delta_{\min}:=\min_{k\neq\ell}\left\|\mu_{k}^{*}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}^{2},\qquad\Delta_{\max}:=\max_{k,\ell}\left\|\mu_{k}^{*}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}^{2}.$ (3.1)
The smallest separation, $\Delta_{\min}$, is known as a key quantity of learning Gaussian location mixtures. Regarding $\Delta_{\max}$, we have $\Delta_{\max}=\Delta_{\min}$ when $L=2$. For $L\geq 3$, [jin2016local] finds that local search procedures may fail to find the global optimum of the likelihood function when one mixture component is significantly farther from the others. This suggests that $\Delta_{\max}$, or more precisely the ratio $\Delta_{\max}/\Delta_{\min}$, influences the difficulty of learning Gaussian mixtures. Another well-known quantity that affects the learning is the smallest mixing probabilities $\pi_{\min}^{*}=\min_{\ell\in[L]}\pi_{\ell}^{*}.$ Our results below characterize the effect of $\Delta_{\min}$, $\Delta_{\max}/\Delta_{\min}$ and $1/\pi_{\min}^{*}$ on the convergence of the EM algorithm. Finally, since we can always subtract the marginal mean of $X$, we assume $\mathbb{E}[X]=\sum_{\ell=1}^{L}\pi_{\ell}^{*}\mu_{\ell}^{*}=0$ throughout our analysis.
The following theorem states the convergence rate of the population level EM iterates, $\theta^{(t)}=(\pi^{(t)},M^{(t)},\Sigma^{(t)})$, given in Eqs. 2.3, 2.4 and 2.5. For any $\theta=(\pi,M,\Sigma)$ and $\theta^{\prime}=(\pi^{\prime},M^{\prime},\Sigma^{\prime})$, we define the distances between their mixing probabilities and mean vectors as
$d(\pi,\pi^{\prime})=\max_{\ell\in[L]}~{1\over\pi_{\ell}^{*}}|\pi_{\ell}-\pi^{\prime}_{\ell}|,\qquad d(M,M^{\prime})=\max_{\ell\in[L]}\left\|\mu_{\ell}-\mu_{\ell}^{\prime}\right\|_{\Sigma^{*}}$ (3.2)
and the distance between their covariance matrices as
$d(\Sigma,\Sigma^{\prime})=\|\Sigma^{*-1/2}(\Sigma-\Sigma^{\prime})\Sigma^{*-1/2}\|_{\rm op}.$ (3.3)
Theorem 2.
Assume there exists some large constant $C>0$ such that
$\Delta_{\min}\geq C\left(\log{\Delta_{\max}\over\Delta_{\min}}+\log{1\over\pi_{\min}^{*}}\right).$ (3.4)
Further assume the initialization $\theta^{(0)}=(\pi^{(0)},M^{(0)},\Sigma^{(0)})$ satisfies
$d(\pi^{(0)},\pi^{*})\leq{1\over 2},\quad d(M^{(0)},M^{*})\leq c_{\mu}\sqrt{\Delta_{\min}},\quad d(\Sigma^{(0)},\Sigma^{*})\leq c_{\Sigma}$ (3.5)
for some absolute positive constants $0\leq 2c_{\mu}+\sqrt{2}c_{\Sigma}<\sqrt{2}-1$. There exists some $\kappa\in(0,1)$ such that $\theta^{(t)}=(\pi^{(t)},M^{(t)},\Sigma^{(t)})$ for any $t\geq 0$, given by Eqs. 2.3, 2.4 and 2.5, satisfy
$d(\pi^{(t)},\pi^{*})+d(M^{(t)},M^{*})+d(\Sigma^{(t)},\Sigma^{*})$~ $\leq~\kappa^{t}\left\{d(\pi^{(0)},\pi^{*})+d(M^{(0)},M^{*})+d(\Sigma^{(0)},\Sigma^{*})\right\}.$
Proof.
The proof is given in Section C.3. ∎
Theorem 2 guarantees linear convergence of the population level EM iterates. In fact, our proof reveals that the contraction rate $\kappa$ satisfies
$\kappa={\Delta_{\max}^{2}\over\pi_{\min}^{*}}\exp(-c\Delta_{\min}),$ (3.6)
for some absolute constant $c>0$ depending only on $c_{\mu}$ and $c_{\Sigma}$ in (3.5). This implies the convergence rate of the EM algorithm can be superlinear, since $\kappa$ decays exponentially when $\Delta_{\min}\equiv\Delta_{\min}(d)\to\infty$ as $d=d(n)\to\infty$.
As highlighted in the Introduction, this is the first convergence result of the population level EM iterates for the $L$-GMM with $L\geq 2$ under full generality. In particular, it covers the case when $\Sigma^{*}$ is known, in which case the EM algorithm only updates $\pi^{(t)}$ and $M^{(t)}$ in the M-step as in Eqs. 2.3 and 2.4 while computes the E-step in (2.1) using $\theta^{(t)}=(\pi^{(t)},M^{(t)},\Sigma^{*})$. The following corollary states a faster convergence rate of such EM iterates.
Corollary 3 (Known $\Sigma^{*}$).
Grant condition (3.4). Assume the initialization $\pi^{(0)}$ and $M^{(0)}$ satisfy (3.5) with $c_{\Sigma}=0$. There exists some $\kappa\in(0,1)$ given by (3.6) such that $\theta^{(t)}=(\pi^{(t)},M^{(t)},\Sigma^{*})$ for all $t\geq 0$, computed in Eqs. 2.3 and 2.4, satisfy
$d(\pi^{(t)},\pi^{*})+d(M^{(t)},M^{*})$ $~\leq~(\kappa/\Delta_{\max})^{t}\left\{d(\pi^{(0)},\pi^{*})+d(M^{(0)},M^{*})\right\}.$
Remark 1 (Separation requirement).
Condition (3.4) states the required separation condition between Gaussian components for the population level EM iterates to converge. It characterizes the effect of both the smallest mixing probability $\pi_{\min}^{*}$ and the ratio between the largest and smallest separations among Gaussian components, $\Delta_{\max}/\Delta_{\min}$. We emphasize that such dependence is particularly informative for $L\geq 3$. In the case of $L=2$, we have $\Delta_{\max}=\Delta_{\min}$, and it is standard to assume that $\pi_{1},\pi_{2}\in(c,1-c)$ for some $c\in(0,1/2)$, under which condition (3.4) reduces to $\Delta_{\min}\geq C$. This is assumed in [EM2017] for symmetric, isotropic setting, and in [cai2019chime] when $\pi^{*}$ and $\Sigma^{*}$ are unknown. For $L\geq 3$, even when $\Delta_{\max}\asymp\Delta_{\min}$ and $\pi_{\min}^{*}\asymp 1/L$, condition (3.4) simplifies to $\Delta_{\min}\geq C\log(L)$, still reflecting the effect of $L$ on the required separation. Such logarithmic dependence on $L$ is known to be necessary for achieving polynomial sample complexity guarantees in the isotropic setting [regev2017learning]. Our condition in (3.4) further shows that when $\Sigma^{*}$ is unknown, the additional price to pay is only a logarithmic factor in $\Delta_{\max}$. We refer to Section 4.4 for comparison with the separation conditions of clustering-based algorithms.
Remark 2 (Initialization).
The requirement in (3.5) on $M^{(0)}$ becomes milder as the separation $\Delta_{\min}$ increases. This is consistent with existing results [EM2017] for $L=2$, and [yan2017convergence, Zhao2020, segol2021improved, kwon2020algorithm] for $L\geq 3$. Compared to the latter, their initialization requirement is comparable to ours, although derived under known / unknown $\pi^{*}$ and $\Sigma^{*}=\mathbf{I}_{d}$. Since we allow both $\pi^{*}$ and $\Sigma^{*}$ unknown, our initialization requirement in (3.5) also depends on $\pi^{(0)}$ and $\Sigma^{(0)}$. The number $1/2$ in $d(\pi^{(0)},\pi)$ can be replaced by any nonnegative value less than $1$. Later in Section 4.2 we introduce an alternative initialization scheme that also enjoys provable guarantees. Detailed discussion on practical choices of initialization is given in Section 5.1.
In the following we mention one simple choice of $\Sigma^{(0)}$ when $\Sigma^{*}$ is unknown. Recall that the EM iterates in Eqs. 2.3, 2.4 and 2.5 update $\Sigma^{(t)}$ by using $\Sigma_{T}^{*}$, $\pi^{(t)}$ and $M^{(t)}$ via (2.6). This suggests using
$\Sigma^{(0)}=\Sigma_{T}^{*}-M^{(0)}{\rm diag}(\pi^{(0)})M^{(0)\top}$ (3.7)
as the initialization for $\Sigma^{*}$, once $\pi^{(0)}$ and $M^{(0)}$ are available. The following corollary states the convergence of the EM algorithm for such choice.
Corollary 4 (A special initialization of $\Sigma^{*}$).
Grant condition (3.4). Assume that the initialization $\theta^{(0)}=(\pi^{(0)},M^{(0)},\Sigma^{(0)})$, with $\Sigma^{(0)}$ given in (3.7), satisfies
$d(\pi^{(0)},\pi^{*})~\leq~{c_{0}\over\Delta_{\max}},\qquad d(M^{(0)},M^{*})~\leq~{c_{0}\over\sqrt{\Delta_{\max}}},$ (3.8)
for some small, absolute constant $c_{0}>0$, There exists some $\kappa\in(0,1)$ given by (3.6) such that $\theta^{(t)}=(\pi^{(t)},M^{(t)},\Sigma^{(t)})$ for all $t\geq 0$, computed in Eqs. 2.3, 2.4 and 2.5, satisfy
$d(\pi^{(t)},\pi^{*})+d(M^{(t)},M^{*})+d(\Sigma^{(t)},\Sigma^{*})$ $~\leq~\kappa^{t}\left\{d(\pi^{(0)},\pi^{*})+d(M^{(0)},M^{*})\right\}.$
Proof.
In (C.14) of Appendix C we prove that for any $\Sigma=\Sigma_{T}^{*}-M{\rm diag}(\pi)M^{\top}$ with $d(M,M^{*})\leq\sqrt{\Delta_{\max}}$, one has
$d(\Sigma,\Sigma^{*})~\leq~2\Delta_{\max}~d(\pi,\pi^{*})+3\sqrt{\Delta_{\max}}~d(M,M^{*}).$ (3.9)
4 follows by invoking (3.9) for $\Sigma^{(0)}$ and noting that (3.5) holds under (3.8). ∎
In order to ensure the choice of $\Sigma^{(0)}$ in (3.7) satisfying (3.5), the requirement on $\pi^{(0)}$ and $M^{(0)}$ gets more restrictive than (3.5) by a multiplicative factor of $1/\Delta_{\max}$.
Proof sketch of Theorem 2.
The full proof can be found in Section C.3 and we only offer its sketch here. For any $\theta^{(t)}=(\pi^{(t)},M^{(t)},\Sigma^{(t)})$ and any $\ell\in[L]$, with $\pi_{\ell}^{(t+1)}$ and $\mu_{\ell}^{(t+1)}$ defined in Eqs. 2.3 and 2.4, we need to analyze the differences
$\pi_{\ell}^{(t+1)}-\pi^{*}_{\ell}$ $=\bm{\pi}_{\ell}(\theta^{(t)})-\bm{\pi}_{\ell}(\theta^{*})=\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta^{(t)})]-\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta^{*})]$
$\mu_{\ell}^{(t+1)}-\mu^{*}_{\ell}$ $=\bm{\mu}_{\ell}(\theta^{(t)})-\bm{\mu}_{\ell}(\theta^{*})={\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta^{(t)})X]\over\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta^{(t)})]}-{\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta^{*})X]\over\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta^{*})]}$
where the first equalities in both lines use the so-called self-consistency of the $Q$ function in (2.2), that is, $\theta^{*}=\operatorname*{argmax}_{\theta}Q(\theta\mid\theta^{*})$ such that $\bm{\pi}_{\ell}(\theta^{*})=\pi_{\ell}^{*}$ and $\bm{\mu}_{\ell}(\theta^{*})=\mu_{\ell}^{*}$. See, details in Section C.1. The first step of our analysis is to re-parametrize the function $\gamma_{\ell}(x;\theta)$ in (2.1) by using $\omega(\theta)=(\pi,M,J(\theta))$ with $J(\theta):=\Sigma^{-1}M$ so that
$\gamma_{\ell}(x;\theta)=\frac{\pi_{\ell}}{\sum_{k=1}^{L}\pi_{k}\exp\left\{(x-{1\over 2}(\mu_{\ell}+\mu_{k}))^{\top}J(\theta)({\bm{e}}_{k}-{\bm{e}}_{\ell})\right\}}=\gamma_{\ell}(x;\omega(\theta))$
as well as $\bm{\pi}_{\ell}(\theta)=\bm{\pi}_{\ell}(\omega(\theta))$ and $\bm{\mu}_{\ell}(\theta)=\bm{\mu}_{\ell}(\omega(\theta))$. This reparametrization avoids direct manipulation of $\Sigma^{-1}$ which alleviates the difficulties of subsequent analysis. In Theorem 19 of Section C.3 we establish the following Lipschitz continuity of the map $\omega(\theta)\mapsto\bm{\pi}_{\ell}(\omega(\theta))$ with the Lipschitz constant $\kappa_{w}\asymp(1/\pi_{\min}^{*})\exp{(-c\Delta_{\min})}$:
${1\over\pi_{\ell}^{*}}\left|\bm{\pi}_{\ell}(\omega(\theta))-\bm{\pi}_{\ell}(\omega(\theta^{*}))\right|$ $~\leq~\kappa_{w}\sqrt{\Delta_{\max}}\bigl\{d(\pi,\pi^{*})+d(M,M^{*})+d(J(\theta),J(\theta^{*}))\bigr\}$ (3.10)
for all $\omega(\theta)$ within a neighborhood of $\omega(\theta^{*})$, in particular, for $\omega(\theta^{(t)})$. We refer to (C.8) the definition of $d(J(\theta),J(\theta^{*}))$. The same bound in (3.10), up to a multiplicative factor $\sqrt{\Delta_{\max}}$, is also shown to hold for $\|\bm{\mu}_{\ell}(\theta)-\bm{\mu}_{\ell}(\theta^{*})\|_{\Sigma^{*}}$. In Lemma 17 of Section C.3 we further derive the following Lipschitz property of $\theta\mapsto J(\theta)$ for any $\theta$ satisfying (3.5):
$d(J(\theta),J(\theta^{*}))~\leq~2d(M,M^{*})+\sqrt{\Delta_{\max}}~d(\Sigma,\Sigma^{*}),$ (3.11)
which in conjunction with (3.10) gives
$d(\pi^{(t+1)},\pi^{*})+d(M^{(t+1)},M^{*})\leq\kappa_{w}\Delta_{\max}\left\{d(\pi^{(t)},\pi^{*})+3d(M^{(t)},M^{*})+\sqrt{\Delta_{\max}}~d(\Sigma^{(t)},\Sigma^{*})\right\}.$
Combining with (3.9), Theorem 2 thus follows.
From the technical perspective, establishing the Lipschitz continuity in (3.10) for both $\bm{\pi}_{\ell}(\omega(\theta))$ and $\bm{\mu}_{\ell}(\omega(\theta))$ is the most challenging part. It requires quantifying the partial derivatives of $\gamma_{\ell}(x;\omega(\theta))$ with respect to $\pi$, $M$, and $J(\theta)$ uniformly over a neighborhood of $\omega(\theta^{*})$. The proofs of (3.10) are provided in Section C.3.1 and Section C.3.2, along with the supporting technical lemmas in Lemmas 20, 22 and 21 of Section C.4.
4 Theoretical analysis of the sample level EM algorithm
We present our theoretical guarantees of the sample level EM algorithm in this section. The analysis is based on two different approaches of analyzing the EM depending on the magnitude of $\Delta_{\min}$. When $\Delta_{\min}$ is small to moderate, as considered in Section 4.1, we directly characterize the convergence of the EM iterates in Eqs. 2.8, 2.9 and 2.10, to the true parameter. On the other hand, when $\Delta_{\min}$ is moderate to large, our analysis in Section 4.2 first establishes the convergence of a surrogate loss related to community detection under model (1.1), and then translates this into a convergence between the EM iterates and $\theta^{*}$.
4.1 Convergence of the EM algorithm under small to moderate separation
In this section we directly analyze the convergence to $\theta^{*}$ of the EM iterates, $\widehat{\theta}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\widehat{\Sigma}^{(t)})$, for $t\geq 0$, given by Eqs. 2.8, 2.9 and 2.10. For each $h\in\{\pi,M,\Sigma\}$, we decompose
$\widehat{h}^{(t+1)}-h^{*}~=~\widehat{\bm{h}}(\widehat{\theta}^{(t)})-\bm{h}(\theta^{*})~=~\left(\widehat{\bm{h}}(\widehat{\theta}^{(t)})-\widehat{\bm{h}}(\theta^{*})\right)+\left(\widehat{\bm{h}}(\theta^{*})-\bm{h}(\theta^{*})\right).$ (4.1)
[…]
It is worth clarifying that the decomposition in (4.1) differs from that introduced by [EM2017] and subsequently followed in [cai2019chime, yan2017convergence, Zhao2020, segol2021improved]. Specifically, their approach decomposes
$\widehat{h}^{(t+1)}-h^{*}~=~\left(\widehat{\bm{h}}(\widehat{\theta}^{(t)})-\bm{h}(\widehat{\theta}^{(t)})\right)+\left(\bm{h}(\widehat{\theta}^{(t)})-\bm{h}(\theta^{*})\right)$
where the second term is relatively easy to handle using the population level contraction property of $\bm{h}$ in Section 3. The first term, on the other hand, characterizes the statistical error, which involves analyzing the empirical process $\sup_{\theta}\|\widehat{\bm{h}}(\theta)-\bm{h}(\theta)\|.$ This is a challenging task, especially when aiming for convergence rates with optimal dependence on the parameters $n$, $d$, $L$, $\pi_{\min}^{*}$, $\Delta_{\min}$, and $\Delta_{\max}$. In contrast, our decomposition in (4.1) only requires establishing that $\sup_{\theta}\|\widehat{\bm{h}}(\theta)-\widehat{\bm{h}}(\theta^{*})\|<1$, which is a relatively easier task to accomplish. Moreover, under (4.1), the statistical error is governed by the pointwise convergence $\widehat{\bm{h}}(\theta^{*})-\bm{h}(\theta^{*})$, for which the optimal convergence rate is comparatively easier to derive, as demonstrated in the next section.
4.1.1 Concentration inequalities of the M-steps at the true parameter
We derive concentration inequalities between the sample-level M-steps $\widehat{\bm{\pi}}_{\ell}(\theta^{*})$, $\widehat{\bm{\mu}}_{\ell}(\theta^{*})$, and $\bm{\widehat{\Sigma}}(\theta^{*})$ and their population-level counterparts. This requires to bound
$\widehat{\bm{\pi}}_{\ell}(\theta^{*})-\bm{\pi}_{\ell}(\theta^{*})$ $=\mathbb{E}_{n}[\gamma_{\ell}(X;\theta^{*})]-\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta^{*})],$
$\widehat{\bm{\mu}}_{\ell}(\theta^{*})-\bm{\mu}_{\ell}(\theta^{*})$ $={\mathbb{E}_{n}[\gamma_{\ell}(X;\theta^{*})X]\over\mathbb{E}_{n}[\gamma_{\ell}(X;\theta^{*})]}-{\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta^{*})X]\over\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta^{*})]}$
as well as $\bm{\widehat{\Sigma}}(\theta^{*})-\bm{\Sigma}(\theta^{*})$ which equals to
$\mathbb{E}_{n}[\gamma_{\ell}(X;\theta^{*})(X-\widehat{\bm{\mu}}_{\ell}(\theta^{*}))(X-\widehat{\bm{\mu}}_{\ell}(\theta^{*}))^{\top}]-\mathbb{E}[\gamma_{\ell}(X;\theta^{*})(X-\mu_{\ell}^{*})(X-\mu_{\ell}^{*})^{\top}].$
The following theorem states the rate of convergence of the above three terms. Recall the distance $d(\Sigma,\Sigma^{*})$ from (3.3).
Theorem 5.
Grant condition (3.4). Assume $n\geq d\log n$ and $n\pi_{\min}^{*}\geq C\log(n)(1\vee\log(n)/d).$ Then with probability at least $1-4n^{-1}$, one has that for all $\ell\in[L]$,
${1\over\pi_{\ell}^{*}}|\widehat{\bm{\pi}}_{\ell}(\theta^{*})-\bm{\pi}_{\ell}(\theta^{*})|$ $~\lesssim~\sqrt{\log n\over n\pi_{\ell}^{*}},$ (4.2)
$\|\widehat{\bm{\mu}}_{\ell}(\theta^{*})-\bm{\mu}_{\ell}(\theta^{*})\|_{\Sigma^{*}}$ $~\lesssim~\sqrt{d\log n\over n\pi_{\ell}^{*}}.$ (4.3)
If, additionally, $n\geq dL^{2}\log n$ and $nd\geq L^{2}\log^{3}n$ hold, then with the same probability,
$d(\bm{\widehat{\Sigma}}(\theta^{*}),\bm{\Sigma}(\theta^{*}))$ $\lesssim\sqrt{d\log n\over n}.$ (4.4)
Proof.
The proof can be found in Section D.1. ∎
4.1.2 Contraction of the sample level EM iterates
[…]
Proposition 6.
Under the separation condition (3.4), assume
$dL\log^{2}(n){\sqrt{\Delta_{\max}(L+\Delta_{\max})}}~\leq c~n\pi_{\min}^{*}$ (4.5)
for some small constant $c>0$. There exists some $\kappa_{n}\in(0,1)$ such that with probability at least $1-5n^{-1}$, the following holds uniformly for any $\theta=(\pi,M,\Sigma)$ satisfying (3.5):
$d(\widehat{\bm{\pi}}(\theta),\widehat{\bm{\pi}}(\theta^{*}))+d(\widehat{\mathbf{M}}(\theta),\widehat{\mathbf{M}}(\theta^{*}))$ $~\leq~\kappa_{n}\left\{d(\pi,\pi^{*})+d(M,M^{*})+\sqrt{\Delta_{\max}}~d(\Sigma,\Sigma^{*})\right\}$
$d(\bm{\widehat{\Sigma}}(\theta),\bm{\widehat{\Sigma}}(\theta^{*}))$ $~\leq~\kappa_{n}\sqrt{\Delta_{\max}}\left\{d(\pi,\pi^{*})+d(M,M^{*})+\sqrt{\Delta_{\max}}~d(\Sigma,\Sigma^{*})\right\}.$
6 is proved in Section D.2. Its proof follows the road-map of proving the population-level result in Theorem 2. By reparametrizing $\gamma_{\ell}(x;\theta)$ as $\gamma_{\ell}(x;\omega(\theta))$, where $\omega(\theta)=(\pi,M,J(\theta))$, the key technical difficulty lies in establishing the Lipschitz continuity of the mappings $\theta\mapsto\widehat{\bm{\pi}}_{\ell}(\omega(\theta))$ and $\theta\mapsto\widehat{\bm{\mu}}_{\ell}(\omega(\theta))$ for all $\theta$ satisfying (3.5). This reduces to deriving uniform convergence of certain empirical quantities of the form $\mathbb{E}_{n}[\partial\gamma_{\ell}(X;\omega(\theta))]$, $\mathbb{E}_{n}[\partial\gamma_{\ell}(X;\omega(\theta))X]$, and $\mathbb{E}_{n}[\partial\gamma_{\ell}(X;\omega(\theta))XX^{\top}]$, where the partial derivatives of $\gamma_{\ell}(X;\omega(\theta))$ are taken with respect to $\pi$, $M$, and $J(\theta)$ separately. The uniform convergence established in Lemma 24, together with the population-level contraction rate $\kappa$ defined in (3.6), determines the contraction rate $\kappa_{n}$. Under conditions (3.4) and (4.5), it is ensured that $\kappa_{n}<1$.
From 6, we also observe that estimating $\Sigma^{*}$ results in a slower algorithmic rate of convergence for the EM algorithm; specifically, it is slower by a factor of $\Delta_{\max}$ compared to the case when $\Sigma^{*}$ is known.
Remark 3 (Condition on $\Delta_{\min}$ and $\Delta_{\max}$).
Condition (4.5) puts a restriction on the largest separation $\Delta_{\max}$. Combined with condition (3.4), 6 requires
$C\left(\log{\Delta_{\max}\over\Delta_{\min}}+\log{1\over\pi_{\min}^{*}}\right)\leq\Delta_{\min}\leq{c~n\pi_{\min}^{*}\over dL(\Delta_{\max}/\Delta_{\min})\log^{2}n}\left(1+{n\pi_{\min}^{*}\over dL^{2}\log^{2}n}\right).$
Although this already allows for a wide range of $\Delta_{\min}$, it still imposes a restriction that prevents $\Delta_{\min}$ from growing too quickly. Since a large $\Delta_{\min}$ is expected to aid parameter estimation, this requirement is somewhat counterintuitive. Our analysis in the next section reveals that this constraint on $\Delta_{\min}$ arises from directly analyzing the contraction of the M-steps, and that it can be removed by adopting a different analytical approach. It is for this reason that the analysis presented in this section is only suitable for small to moderate separation.
Combining Theorem 5 with 6 yields the following guarantees of the sample level EM iterates. We state our results for known $\Sigma^{*}$ and unknown $\Sigma^{*}$ separately.
Theorem 7 (Known $\Sigma^{*}$).
Grant conditions (3.4) and (4.5). Assume $nd\geq L^{2}\log^{3}n$. Whenever the initialization $\widehat{\pi}^{(0)}$ and $\widehat{M}^{(0)}$ satisfy (3.5) with $c_{\Sigma}=0$, with probability at least $1-5n^{-1}$, the EM iterates $\widehat{\theta}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\Sigma^{*})$ in Eqs. 2.8 and 2.9 satisfy
$d(\widehat{\pi}^{(t)},\pi^{*})+d(\widehat{M}^{(t)},M^{*})$ $~\lesssim~\sqrt{d\log n\over n\pi_{\min}^{*}},\qquad\forall~t\geq C\log n.$ (4.6)
Proof.
The proof is given in Section D.3. ∎
Theorem 7 states that, with a suitable initialization, the EM iterates achieve a statistical precision of order $\sqrt{d\log n/(n\pi_{\min}^{*})}$ after only $\mathcal{O}(\log n)$ iterations. The next theorem extends such guarantees to the case when $\Sigma^{*}$ is unknown and needs to be estimated.
Theorem 8 (Unknown $\Sigma^{*}$).
Grant condition (3.4), $nd\geq L^{2}\log^{3}n$ and
$dL\log^{2}(n)\Delta_{\max}{\sqrt{(\Delta_{\max}+1/\pi_{\min}^{*})(L+\Delta_{\max})}}~\leq c~n\pi_{\min}^{*}.$ (4.7)
Whenever $\widehat{\theta}^{(0)}=(\widehat{\pi}^{(0)},\widehat{M}^{(0)},\widehat{\Sigma}^{(0)})$ satisfies (3.5), with probability $1-5n^{-1}$, the EM iterates $\widehat{\theta}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\widehat{\Sigma}^{(t)})$ in Eqs. 2.8, 2.9 and 2.10 satisfy: for all $t\geq C\log n$,
$d(\widehat{\pi}^{(t)},\pi^{*})+d(\widehat{M}^{(t)},M^{*})\lesssim\sqrt{d\log n\over n\pi_{\min}^{*}},\qquad d(\widehat{\Sigma}^{(t)},\Sigma^{*})\lesssim\sqrt{d\log n\over n}.$ (4.8)
Proof.
The proof is given in Section D.3. ∎
Comparing to (4.5), the required condition in (4.7) becomes stronger due to the slower rate of convergence for estimating $\Sigma^{*}$, as seen in 6.
Remark 4 (A special initialization of $\Sigma^{*}$).
Recall the expression of $\widehat{\Sigma}^{(t)}$ in (2.11). If we opt to initialize $\widehat{\Sigma}^{(0)}=\widehat{\Sigma}_{T}-\widehat{M}^{(0)}{\rm diag}(\widehat{\pi}^{(0)})\widehat{M}^{(0)\top}$ with $\widehat{\Sigma}_{T}$ being the marginal sample covariance matrix, then similar as 4, under condition (4.7) and provided that $\widehat{\pi}^{(0)}$ and $\widehat{M}^{(0)}$ satisfy (3.8), the conclusion in Theorem 8 remains. Its proof can be found in Section D.4.
4.2 Convergence of the EM algorithm under moderate to large separation
In this section, we establish the convergence result of the EM algorithm when $\Delta_{\min}$ is moderate to large. Our analysis is based on a surrogate loss function defined below. To provide some intuition, we note that for any data point $i\in[n]$ and label $\ell\in[L]$, when $\Delta_{\min}$ increases, one should expect $\mathbb{P}_{\theta^{*}}(Y_{i}=\ell\mid X_{i})=\gamma_{\ell}(X_{i};\theta^{*})$ in (2.1) to get closer to one if $Y_{i}=\ell$, and closer to zero otherwise. This is also expected for any $\theta$ that is close to $\theta^{*}$. Therefore, a useful metric to evaluate any given $\theta$ in this case should both characterize the discrepancy between $\gamma_{\ell}(X_{i};\theta)$ and the true labels $1\{Y_{i}=\ell\}$ for all $i\in[n]$ and $\ell\in[L]$, and reflect the separation magnitude.
In light of this, for any $\theta$ with $\gamma_{\ell}(\cdot;\theta)$ given in (2.1), we define the loss function
$\phi(\theta):=\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}\gamma_{\ell}(X_{i};\theta)\|\mu_{\ell}^{*}-\mu_{a}^{*}\|_{\Sigma^{*}}^{2}.$ (4.9)
[…]
where $n_{\ell}=\sum_{i=1}^{n}1\{Y_{i}=\ell\}$. Although loss functions analogous to (4.9) have been employed in the analysis of clustering methods based on hard assignments [chen2024achieving, gao2022iterative], (4.9) introduces a novel loss function tailored to the setting of soft assignments.
The following theorem states the convergence of $\phi(\widehat{\theta}^{(t)})$, for $t\geq 0$, with $\widehat{\theta}^{(t)}$ being the EM iterates in Eqs. 2.8, 2.9 and 2.10. We start with the case of known $\Sigma^{*}$.
Theorem 9 (Known $\Sigma^{*}$).
Assume $\Delta_{\min}\geq C\log(1/\pi_{\min}^{*})$ and
$n\geq Cd\left\{{\log n\over\pi_{\min}^{*}},~L^{2}\log n,~{L^{2}\log(n\Delta_{\max})\over\pi_{\min}^{*2}\Delta_{\min}^{2}}\right\}$
for some large constant $C>0$. For any $\widehat{\theta}^{(0)}=(\widehat{\pi}^{(0)},\widehat{M}^{(0)},\Sigma^{*})$ satisfying
  • (a)
    either condition (3.5) with $c_{\mu}\leq\min\{(\sqrt{2}-1)/2,~c~\pi_{\min}^{*}\Delta_{\min}/L\},$
  • (b)
    or $(1/\pi_{\min}^{*}\Delta_{\min})(1+{L/\pi_{\min}^{*}\Delta_{\min}})\phi(\widehat{\theta}^{(0)})\leq c~n,$
for some small constant $c>0$, with probability at least $1-4n^{-1}-\exp(-c^{\prime}\Delta_{\min})$,
$\phi(\widehat{\theta}^{(t)})~\leq~n\exp(-c^{\prime}\Delta_{\min})+{1\over 2}\phi(\widehat{\theta}^{(t-1)}),\qquad\forall~t\geq 1.$
Proof.
The proof can be found in Section E.2. ∎
Theorem 9 states that $\phi(\widehat{\theta}^{(t)})$ converges to $n\exp(-c^{\prime}\Delta_{\min})$ linearly as $t\to\infty$. When the latter is small, one should expect $\widehat{\theta}^{(t)}$ converges to $\widetilde{\theta}$ given in (4.10) which further converges to $\theta^{*}$ in the optimal rate. We state such convergence of $\widehat{\theta}^{(t)}$ to $\theta^{*}$ in Theorem 11.
The key difference from the convergence of $\widehat{\theta}^{(t)}$ established in 6 and 7 is that the convergence of $\phi(\widehat{\theta}^{(t)})$ in Theorem 9 only requires a lower bound on $\Delta_{\min}$, and benefits from larger values of $\Delta_{\min}$. This aligns with the intuition that recovering the labels $Y_{1},\ldots,Y_{n}$ becomes easier as $\Delta_{\min}$ increases.
Regarding initialization requirements, condition (a) is stronger than (3.5) in Theorem 2 only when $\Delta_{\min}\pi_{\min}^{*}\leq L/c$. This extra requirement is needed to have the factor $1/2$ in front of $\phi(\widehat{\theta}^{(t-1)})$. On the other hand, the convergence of $\phi(\widehat{\theta}^{(t)})$ allows for an alternative initialization scheme as given in (b) of Theorem 9. We refer to Section 5.1 for detailed discussion on initialization.
The following proposition extends Theorem 9 to unknown $\Sigma^{*}$ in which the required conditions become slightly stronger.
Theorem 10 (Unknown $\Sigma^{*}$).
For some large constant $C>0$, assume $n\pi_{\min}^{*}\geq Cd\log n$, $\Delta_{\min}\geq C\log(1/\pi_{\min}^{*})+C\log\log n$ and
$n\geq CdL^{2}\log(n\Delta_{\max})\left({1\over\pi_{\min}^{*2}\Delta_{\min}^{2}}+{\log n\over\Delta_{\min}}+1\right).$ (4.11)
For any $\widehat{\theta}^{(0)}=(\widehat{\pi}^{(0)},\widehat{M}^{(0)},\widehat{\Sigma}^{(0)})$ satisfying
  • (a)
    either (3.5) with $c_{\mu}+c_{\Sigma}\leq\min\{(\sqrt{2}-1)/2,c(1+\pi_{\min}^{*}\Delta_{\min})/L\},$
  • (b)
    or $({\Delta_{\max}/\Delta_{\min}}+L+{L/\pi_{\min}^{*2}\Delta_{\min}^{2}})\phi(\widehat{\theta}^{(0)})\leq c~n$
for some small constant $c>0$, with probability $1-4n^{-1}-\exp(-c^{\prime}\Delta_{\min})$,
$\phi(\widehat{\theta}^{(t)})~\leq~n\exp(-c^{\prime}\Delta_{\min})+{1\over 2}\phi(\widehat{\theta}^{(t-1)}),\qquad\forall~t\geq 1.$
Proof.
Its proof is deferred to Section E.1. ∎
Comparing to Theorem 9, both the condition in (4.11) and the initialization requirements in (a) and (b) become stronger due to estimating $\Sigma^{*}$. Although the required lower bound on $\Delta_{\min}$ has one additional $\log\log n$ term, the results in Theorem 10 still benefit from larger values of $\Delta_{\min}$, in contrast to Theorem 8.
Proof sketch of Theorems 9 and 10.
[…]
I $=\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}\gamma_{\ell}(X_{i};\theta^{*})\|\mu_{\ell}^{*}-\mu_{a}^{*}\|_{\Sigma^{*}}^{2},$
II $=\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}1\left\{\left|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|>c^{\prime}\|\mu_{\ell}^{*}-\mu_{a}^{*}\|_{\Sigma^{*}}^{2}\right\}\|\mu_{\ell}^{*}-\mu_{a}^{*}\|_{\Sigma^{*}}^{2},$
III $=\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}{n_{a}\over\|\mu_{\ell}^{*}-\mu_{a}^{*}\|_{\Sigma^{*}}^{2}}\left(\|\widehat{\mu}_{\ell}^{(t)}-\widetilde{\mu}_{\ell}\|_{\Sigma^{*}}^{2}+\|\widehat{\mu}_{a}^{(t)}-\widetilde{\mu}_{a}\|_{\Sigma^{*}}^{2}+\|(\widehat{\Sigma}^{(t)}-\widetilde{\Sigma})\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{\Sigma^{*}}^{2}\right).$
Here $N_{1},\ldots,N_{n}$ are i.i.d. from $\mathcal{N}_{d}(0,\mathbf{I}_{d})$ and $\widetilde{J}:=\widetilde{\Sigma}^{-1}\widetilde{M}$. The remaining steps analyze each term separately. The first term is relatively easy to analyze and the second step of our proof shows that $\textrm{I}\lesssim n\exp(-c^{\prime}\Delta_{\min})$ with probability at least $1-\exp(-c^{\prime}\Delta_{\min})$.
To bound II, one could apply the inequality
$|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})|\leq\|N_{i}\|_{2}\|(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{\Sigma^{*}}$
along with $\max_{i\in[n]}\|N_{i}\|_{2}=\mathcal{O}_{\mathbb{P}}(\sqrt{d+\log n})$ and $\|(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{\Sigma^{*}}=\mathcal{O}_{\mathbb{P}}(\sqrt{(d+\log n)/n\pi_{\min}^{*}})$, which follow from the analysis of $\widetilde{M}$ and $\widetilde{\Sigma}$ in (4.10). However, this approach leads to the suboptimal requirement $n\gg d^{2}$ in order to ensure that II is negligible. Instead, our analysis in Lemma 27 of Section E.3 employs a leave-one-out technique to control this term more carefully, thereby improving the requirement to $n\gg d$.
[…]
Using Lemma 28 in the fourth step above, the convergence of $\phi(\widehat{\theta}^{(t)})$ can be translated into the convergence of $\widehat{\theta}^{(t)}$ toward $\widetilde{\theta}$. Combined with the analysis of the estimation error of $\widetilde{\theta}$, this immediately yields the following convergence rate of the EM algorithm for estimating $\theta^{*}$.
Theorem 11.
Under conditions in Theorem 10, with probability at least $1-n^{-1}-\exp(-c^{\prime}\Delta_{\min})$, the following holds for all $t\geq C\log n$,
$d(\widehat{\pi}^{(t)},\pi^{*})+d(\widehat{M}^{(t)},M^{*})$ $~\lesssim~\sqrt{d\log n\over n\pi_{\min}^{*}}+\exp(-c^{\prime}\Delta_{\min}),$
$d(\widehat{\Sigma}^{(t)},\Sigma^{*})$ $~\lesssim~\sqrt{d\log n\over n}+\exp(-c^{\prime}\Delta_{\min}).$
As a result, when $\Delta_{\min}\geq\log(n)/(2c^{\prime})$, with probability at least $1-\mathcal{O}(n^{-1})$, for all $t\geq C\log n$,
$d(\widehat{\pi}^{(t)},\pi^{*})+d(\widehat{M}^{(t)},M^{*})~\lesssim~\sqrt{d\log n\over n\pi_{\min}^{*}},\qquad d(\widehat{\Sigma}^{(t)},\Sigma^{*})~\lesssim~\sqrt{d\log n\over n}.$
Proof.
See Section E.4. ∎
Compared to Theorem 8, Theorem 11 yields the same rates of convergence of the EM algorithm in the regime $\Delta_{\min}\gtrsim\log n$. In the next section, we summarize the theoretical guarantees of using both analytical approaches and demonstrate that the EM algorithm achieves minimax optimality under the separation condition (3.4).
4.3 Minimax optimality of the EM algorithm
In this section we establish the minimax optimality of the EM algorithm under the $L$-GMM. For obtaining the upper bounds, combining the results in Sections 4.1 and 4.2 gives two scenarios in Table 2 under either of which the EM estimator achieves the rates in (4.6) for known $\Sigma^{*}$ and (4.8) for unknown $\Sigma^{*}$.
Table 2: Settings in which the EM estimation achieves the rates in (4.6) and (4.8)
Known $\Sigma^{*}$ Unknown $\Sigma^{*}$
Analysis 1 $\log{\Delta_{\max}\over\pi_{\min}^{*}}\lesssim\Delta_{\min}\lesssim{\Delta_{\min}\over\Delta_{\max}}{n\pi_{\min}^{*}\over dL^{3/2}\log^{2}(n)}$ $\log{\Delta_{\max}\over\pi_{\min}^{*}}\lesssim\Delta_{\min}\lesssim{\Delta_{\min}\over\Delta_{\max}}\sqrt{n\pi_{\min}^{*3/2}\over dL^{3/2}\log^{2}(n)}$
$\widehat{\theta}^{(0)}$ satisfies (3.5) or (b) of Theorem 9 $\widehat{\theta}^{(0)}$ satisfies (3.5) or (b) of Theorem 10
Analysis 2 $\Delta_{\min}\gtrsim\log(n)\vee\sqrt{dL^{2}\log(n\Delta_{\max})\over n\pi_{\min}^{*2}}$ $\Delta_{\min}\gtrsim\log(n)\vee\sqrt{dL^{2}\log(n\Delta_{\max})\over n\pi_{\min}^{*2}}$
$dL^{2}\log n\lesssim n,\quad d\log n\lesssim n\pi_{\min}^{*}$ $dL^{2}\log(n\Delta_{\max})\lesssim n,\quad d\log n\lesssim n\pi_{\min}^{*}$
$\widehat{\theta}^{(0)}$ satisfies (a) or (b) of Theorem 9 $\widehat{\theta}^{(0)}$ satisfies (a) or (b) of Theorem 10
As illustrated in Fig. 1, the two analytical approaches require different sets of conditions on $\Delta_{\min}$, which overlap under mild assumptions on $\Delta_{\max}/\Delta_{\min}$. In this regime, EM achieves the rates in Sections 4.1 and 4.2 which are shown below to be minimax optimal.
Figure 1: $\Delta_{\min}$ requirement in which the EM estimation achieves the rates in (4.6) and (4.8). Left: known $\Sigma^{*}$; Right: unknown $\Sigma^{*}$.
To further facilitate understanding, we state a simplified corollary below which invokes the results in Section 4.1 for $\Delta_{\min}=\mathcal{O}(\log n)$ and states the guarantees of the EM algorithm for both known and unknown $\Sigma^{*}$.
Corollary 12.
Assume (3.4) and $dL^{2}\log(n)\leq cn\pi_{\min}^{*2}\Delta_{\min}^{2}$ for some small constant $c>0$.
  1. (i)
    For known $\Sigma^{*}$, if $(\Delta_{\max}/\Delta_{\min})dL^{3/2}\log^{3}(n)\leq cn\pi_{\min}^{*}$ holds, then with probability $1-\mathcal{O}(n^{-1})$, the bound in (4.6) holds for the EM iterates $\widehat{\theta}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\Sigma^{*})$ with $\widehat{\theta}^{(0)}$ satisfying either (a) or (b) of Theorem 9, after $\mathcal{O}(\log n)$ iterations.
  2. (ii)
    For unknown $\Sigma^{*}$, if $(\Delta_{\max}/\Delta_{\min})^{2}dL^{3/2}\log^{4}(n)\leq cn\pi_{\min}^{*3/2}$ holds, then with probability $1-\mathcal{O}(n^{-1})$, the bound in (4.6) holds for the EM iterates $\widehat{\theta}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\widehat{\Sigma}^{(t)})$ with $\widehat{\theta}^{(0)}$ satisfying either (a) or (b) of Theorem 10, after $\mathcal{O}(\log n)$ iterations.
To benchmark the rates in Eqs. 4.6 and 4.8, we establish the minimax lower bounds of estimating $\theta^{*}$ under model (1.1). We consider the following parameter space: for any $\alpha>0,\delta>0$ and $A\geq 1$,
$\Theta(\alpha,\delta,A)=\left\{(\pi,M,\Sigma):\min_{\ell\in[L]}\pi_{\ell}\geq\alpha,~\delta\leq\|\mu_{k}-\mu_{\ell}\|_{\Sigma}^{2}\leq A\delta,~\forall~k\neq\ell,~~\Sigma\succ 0\right\}.$
Note that for any $\theta^{*}\in\Theta(\alpha,\delta,A)$, one has $\pi_{\min}^{*}\geq\alpha$, $\Delta_{\min}\geq\delta$ and $\Delta_{\max}/\Delta_{\min}\leq A$. The following theorem states the minimax lower bounds of estimating $\theta^{*}$ under model (1.1) for fixed $\alpha$, $\delta$ and $A$. The magnitude of $A$ also depends on the relative size of $d$ and $L$. Recall $d(\widehat{M},M)$ and $d(\widehat{\Sigma},\Sigma)$ from (3.2) and (3.3).
Theorem 13.
Under model (1.1), fix any $\alpha$ and $\delta$ such that $d\leq n\alpha\delta$. Fix any $A\geq 4$ if $d\geq L-1$, and $A\geq CL^{2/d}$ with some absolute constant $C\geq 1$ if d<L1d<L-1. Then there exists some absolute constants $c_{0}\in(0,1)$ and $c_{1}>0$ such that
$\inf_{\widehat{\theta}}\sup_{\theta\in\Theta(\alpha,\delta,A)}\mathbb{P}_{\theta}\left\{d(\widehat{M},M)\geq c_{1}\sqrt{d\over n\alpha},~d(\widehat{\Sigma},\Sigma)\geq c_{1}\sqrt{d\over n}\right\}\geq c_{0},$
where the infimum $\widehat{\theta}$ is taken over all estimators.
Proof.
The proof can be found in Appendix F. ∎
Theorem 13 implies that under the conditions of 12, the EM algorithm is minimax optimal, up to a multiplicative logarithmic factor. The restriction on $\Delta_{\max}/\Delta_{\min}$ depends on the size $d$ and $L$. Specifically, for $d\geq L-1$, the most favorable case is $\Delta_{\max}/\Delta_{\min}\asymp 1$ whence 12 requires $dL^{3/2}\log^{4}(n)=\mathcal{O}(n\pi_{\min}^{*})$ for known $\Sigma^{*}$ and $dL^{3/2}\log^{4}(n)=\mathcal{O}(n\pi_{\min}^{*3/2})$ for unknown $\Sigma^{*}$. While for d<L1d<L-1, the most ideal case is $\Delta_{\max}/\Delta_{\min}\asymp L^{2/d}$ whence 12 requires $n\pi_{\min}^{*}\gg dL^{2/d}L^{3/2}\log^{3}(n)$ for known $\Sigma^{*}$ and $n\pi_{\min}^{*3/2}\gg dL^{4/d}L^{3/2}\log^{4}(n)$ for unknown $\Sigma^{*}$. Such requirement becomes the strongest in the univariate case $d=1$ whence one needs $n\pi_{\min}^{*}\gg L^{7/2}\log^{3}(n)$ and $n\pi_{\min}^{*3/2}\gg L^{11/2}\log^{4}(n)$ for known and unknown $\Sigma^{*}$, respectively. It is worth mentioning that in such univariate setting with $\Sigma^{*}$ known, even for fixed $L$ and $\Delta_{\min}$, the denoising method-of-moment approach in [WuYang2020] requires at least $(\Delta_{\max}/\Delta_{\min})L^{4}\leq\pi_{\min}^{*2}n^{1/(2L-1)}$ to achieve the $\sqrt{n}$-rate, which is stronger than what 12 requires. For either $L=L(n)$ and $\Delta_{\min}=\Delta_{\min}(n)$ satisfying (3.4), or unknown $\Sigma^{*}$, the rate obtained in WuYang2020 is not optimal.
Finally, to conclude this section, we state a simplified result for the $2$-GMM with both unknown $\pi^{*}$ and $\Sigma^{*}$, and compared it with the existing result.
Corollary 14 ($L=2$).
Assume $\Delta_{\min}\geq C\log(1/\pi_{\min}^{*})$, $Cd\log(n)\leq n\pi_{\min}^{*2}\Delta_{\min}^{2}$ and $Cd\log^{4}(n)\leq n\pi_{\min}^{*3/2}$ for some large constant $C>0$. Further assume $\widehat{\theta}^{(0)}$ satisfies either (3.5), or $\phi(\widehat{\theta}^{(0)})/n\leq c(1\wedge\pi_{\min}^{*2}\Delta_{\min}^{2})$. Then with probability $1-\mathcal{O}(n^{-1})$, (4.6) holds for all $t\geq C\log n$.
For the $2$-GMM with both $\pi^{*}$ and $\Sigma^{*}$ unknown, [cai2019chime] establishes convergence of the EM algorithm under
$C\leq\Delta_{\min}\leq C^{\prime},\quad C\leq\lambda_{d}(\Sigma^{*})\leq\lambda_{1}(\Sigma^{*})\leq C^{\prime},\quad c<\pi_{1}^{*},\pi_{2}^{*}\leq 1-c.$ (4.12)
Their initialization requirement is specified on $\pi^{(0)},M^{(0)}$ and $\beta^{(0)}$, with the last quantity being the initialization of $\beta^{*}:=\Sigma^{*-1}(\mu_{2}^{*}-\mu_{1}^{*})$. Concretely, their condition (C1) requires at least
$d(\pi^{(0)},\pi^{*})+d(M^{(0)},M^{*})+\|\beta^{(0)}-\beta^{*}\|_{2}=\mathcal{O}(\sqrt{\Delta_{\min}/d}).$ (4.13)
Since our initialization of $\beta^{*}$ is $\beta^{(0)}=\Sigma^{(0)-1}(\mu_{2}^{(0)}-\mu_{1}^{(0)})$ with $\Sigma^{(0)}$ given in Theorem 2, inspecting our proof reveals that condition (3.5) can be replaced by
$d(\pi^{(0)},\pi^{*})<1,\quad\|\beta^{(0)}-\beta^{*}\|_{2}+d(M^{(0)},M^{*})~\leq~c_{0}\sqrt{\Delta_{\min}},$
a much milder condition than (4.13) for large $d$. On the other hand, although our optimal rates in (4.6) coincide with those derived in [cai2019chime] under their condition (4.12), our analysis allows $\Delta_{\min}$ to grow and accommodates highly unbalanced mixing weights with $\pi_{\min}^{*}=o(1)$ as $n\to\infty$. In this more general setting, an inspection of the proof in [cai2019chime] reveals that their analysis yields $d(\widehat{M},M^{*})\lesssim(\pi_{\max}^{*}/\pi_{\min}^{*2})\Delta_{\max}^{3/2}\sqrt{{d\log(d)/n}}$ whereas our analysis gives the optimal rate in (4.8).
4.4 Application to community detection
As noted earlier, our second analytical approach in Section 4.2 is closely connected to the community detection problem, which concerns predicting the labels $Y_{1},\ldots,Y_{n}$. In this section, we discuss its implications and show that both of our analytical approaches to the EM algorithm can in fact be used to quantify its performance in community detection.
For any parameter $\theta$, the Bayes rule of predicting the label of any $x$ from (1.1) is
$g(x;\theta)=\operatorname*{argmax}_{\ell\in[L]}~\mathbb{P}_{\theta}(Y=\ell\mid X=x)=\operatorname*{argmax}_{\ell\in[L]}~\gamma_{\ell}(x;\theta).$ (4.14)
The misclustering error between the true labels $Y_{1},\ldots,Y_{n}$ and $g(X_{1};\theta),\ldots,g(X_{n};\theta)$ is commonly measured by the Hamming distance, given by
$\ell(\theta):={1\over n}\sum_{i=1}^{n}1\{Y_{i}\neq g(X_{i};\theta)\}.$ (4.15)
In Section E.5, we prove $\ell(\theta)\leq 2\phi(\theta)/(n\Delta_{\min}),$ so that invoking Theorem 10 readily yields the convergence rate of the misclustering error of EM when $\Delta_{\min}\gtrsim\log(1/\pi_{\min}^{*})+\log\log n$. Combined with Theorem 8 of using the first analysis in Section 4.1, we are able to extend this to $\Delta_{\min}\gtrsim\log(1/\pi_{\min}^{*})+\log(\Delta_{\max}\wedge\log n)$.
Theorem 15.
Grant $n\pi_{\min}^{*}\geq Cd\log n$ and $\Delta_{\min}\geq C\log(1/\pi_{\min}^{*})+C\log(\Delta_{\max}\wedge\log n)$ for some large constant $C>0$.
  • If $\Delta_{\max}\leq\log n$, assume (4.7) with $\widehat{\theta}^{(0)}$ satisfying either (3.5) or (b) of Theorem 10;
  • If $\Delta_{\max}\geq\log n$, assume (4.11) with $\widehat{\theta}^{(0)}$ satisfying either (a) or (b) of Theorem 10;
then with probability at least $1-n^{-1}-\exp(-c\Delta_{\min})$ for some absolute constant $0<c<1/8$,
$\ell(\widehat{\theta}^{(t)})~\leq~\exp(-c\Delta_{\min}),\qquad\forall~t\geq 2\log n.$ (4.16)
Furthermore, whenever $\Delta_{\min}>\log(n)/c$, with probability at least $1-\mathcal{O}(n^{-1})$, we have the exact recovery, that is, $\ell(\widehat{\theta}^{(t)})=0$ for all $t\geq 2\log n$.
Proof.
See Section E.5. ∎
It is worth comparing Theorem 15 with the existing literature on the community detection problem under GMMs. A large body of work focuses on exact label recovery, with spectral clustering being the primary driving approach [dasgupta1999learning, sanjeev2001learning, vempala2002spectral, kannan2008spectral, awasthi2012improved, achlioptas2005spectral]. Along this direction, the best known separation condition for exact recovery [achlioptas2005spectral] requires
$\min_{k\neq\ell}{\|\mu_{k}^{*}-\mu_{\ell}^{*}\|_{2}^{2}\over\lambda_{1}(\Sigma^{*})}~\gtrsim~{1\over\pi_{\min}^{*}}+L^{2}+{L\log(nL)}$ (4.17)
provided that $n\pi_{\min}^{*}\gtrsim L(d+\log L)\log n$. By contrast, Theorem 15 for exact recovery requires the separation condition $\Delta_{\min}\gtrsim\log(1/\pi_{\min}^{*})+\log n$. Since $\Delta_{\min}\geq\min_{k\neq\ell}\|\mu_{k}^{*}-\mu_{\ell}^{*}\|_{2}^{2}/\lambda_{1}(\Sigma^{*})$, ours is weaker than (4.17), especially when $\pi_{\min}^{*}$ is small or $L$ large. As shown in [Ndaoud2022], the requirement $\Delta_{\min}\gtrsim\log n$ is necessary for exact recovery even for the symmetric, isotropic 2-GMM. Regarding the sample complexity, under condition (4.17), our condition (4.11) simplifies to $dL^{2}\log(n\Delta_{\max})\lesssim n$ which is milder when $\log(\Delta_{\max}/\Delta_{\min})\ll 1/\pi_{\min}^{*}$.
The first rate of convergence for the misclustering error was established in lu2016statistical, where the authors showed that Lloyd’s algorithm, when initialized by spectral clustering and run for $\lceil 4\log n\rceil$ iterations, yields estimated labels $\widehat{Y}_{1},\ldots,\widehat{Y}_{n}$ that satisfy
${1\over n}\sum_{i=1}^{n}1\{Y_{i}\neq\widehat{Y}_{i}\}\leq\exp\left\{-(1+o(1))\min_{k\neq\ell}{\|\mu_{k}^{*}-\mu_{\ell}^{*}\|_{2}^{2}\over 8\lambda_{1}(\Sigma^{*})}\right\}$ (4.18)
with high probability. Their result requires the separation condition
$\min_{k\neq\ell}\frac{\|\mu_{k}^{*}-\mu_{\ell}^{*}\|_{2}^{2}}{\lambda_{1}(\Sigma^{*})}~\gg~\frac{L}{\pi_{\min}^{*}}+\frac{dL^{2}}{n\pi_{\min}^{*}},$ (4.19)
under the assumption that $L\log n\leq n\pi_{\min}^{*2}$. In comparison, Theorem 15 requires a weaker separation when $\min\{\log(\Delta_{\max}/\Delta_{\min}),\log\log n\}\ll L/\pi_{\min}^{*}$. This is due to the fact that the EM algorithm uses soft assignments in the E-steps, which influence the updates of the mean vectors in the M-steps. On the other hand, under condition (4.19), Theorem 15 still requires $dL\log^{4}(n)\leq n\pi_{\min}^{*}$, which is a stronger requirement on the sample size $n$ than that in [lu2016statistical]. This is attributable to the fact that the EM algorithm also estimates the covariance matrix $\Sigma^{*}$.
Precisely because Lloyd’s algorithm does not explicitly estimate $\Sigma^{*}$, the rates in (4.18) are only optimal in the special case where $\Sigma^{*}\propto\mathbf{I}_{d}$. chen2024achieving recently established the minimax lower bound of the misclustering error for a fixed, general $\Sigma^{*}$ under the $L$-GMM:
$\inf_{\widehat{Y}_{1},\ldots,\widehat{Y}_{n}}\sup_{Y_{1},\ldots,Y_{n}}\mathbb{E}_{X\mid Y}\left[{1\over n}\sum_{i=1}^{n}1\{Y_{i}\neq\widehat{Y}_{i}\}\right]\geq\exp\left\{-(1+o(1)){\Delta_{\min}\over 8}\right\}$
provided that $\Delta_{\min}/\log L\to\infty$. Here $\mathbb{E}_{X\mid Y}$ denotes the conditional expectation of $X_{1},\ldots,X_{n}$ under model (1.1) when $Y_{1},\ldots,Y_{n}$ are treated as deterministic quantities. In [chen2024achieving], they also show that a modified Lloyd’s algorithm that uses the hard assignments but also estimates $\Sigma^{*}$ provably achieves the above lower bound under conditions including $L=\mathcal{O}(1)$, $1/\pi_{\min}^{*}=\mathcal{O}(1)$, $d^{2}=\mathcal{O}(n)$, $\Delta_{\min}\to\infty$ and $c\leq\lambda_{d}(\Sigma^{*})\leq\lambda_{1}(\Sigma^{*})\leq C$, as well as an initialization such that $\sum_{i=1}^{n}1\{Y_{i}\neq\widehat{Y}^{(0)}\}=o(n)$. By contrast, our result in Theorem 15 shows that the EM algorithm also achieves the optimal rate of convergence under milder conditions, owing to the use of soft assignments.
5 Initialization and numerical studies
We discuss the practical choice of initializations in Section 5.1. Numerical studies are conducted in Section 5.2 to corroborate our theoretical findings.
5.1 Choice of the initialization
Our analysis in Sections 3 and 4 has focused on the convergent behavior of EM under the assumption that a suitable initialization is provided. In this section, we discuss practical procedures for choosing such initializations. Recall that our theoretical results accommodate two types of initialization schemes: (a) Initializing with $\theta^{(0)}$ that is close to $\theta^{*}$ in the sense of (3.5), or (b) Initializing the posterior probabilities $\gamma_{\ell}(X_{i};\theta^{(0)})$ for $i\in[n]$ and $\ell\in[L]$ such that $\phi(\theta^{(0)})$ in (4.9) is small, as required by (b) of Theorem 10.
For scheme (b), if some preliminary estimates of the labels can be obtained, say $\widehat{Y}_{1},\ldots,\widehat{Y}_{n}$, then one can initialize by setting $\gamma_{\ell}(X_{i};\theta^{(0)}):=1\{\widehat{Y}_{i}=\ell\}$ so that
$\phi(\theta^{(0)})=\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}1\{\widehat{Y}_{i}=\ell\}\|\mu_{a}^{*}-\mu_{\ell}^{*}\|_{\Sigma^{*}}^{2}\leq{\Delta_{\max}}~{1\over n}\sum_{i=1}^{n}1\{Y_{i}\neq\widehat{Y}_{i}\}.$
This leads to sufficient conditions on the proportion of misclassified data points under which (b) of Theorem 9 and (b) of Theorem 10 hold. For example, when the separation condition (4.19) is satisfied, one can use the Lloyd’s algorithm to obtain preliminary labels. Indeed, the rates of Lloyd’s algorithm given in (4.18) ensure that both (b) of Theorem 9 and (b) of Theorem 10 hold, provided the following balancing condition is met:
$\min_{k\neq\ell}{\|\mu_{k}^{*}-\mu_{\ell}^{*}\|_{2}^{2}\over\lambda_{1}(\Sigma^{*})}~\gg~\log{\max_{k,\ell}\|\mu_{k}^{*}-\mu_{\ell}^{*}\|_{2}^{2}\over\min_{k\neq\ell}\|\mu_{k}^{*}-\mu_{\ell}^{*}\|_{2}^{2}}+\log{\lambda_{1}(\Sigma^{*})\over\lambda_{p}(\Sigma^{*})}.$
This condition ensures that neither the cluster centers nor the covariance structure are too unbalanced relative to the separation.
On the other hand, when the separation between components is insufficient, scheme (a) can still be employed, provided that a consistent estimator of $\theta^{*}$ is available. In the context of Remark 4, it can be further simplified to find a consistent estimator of $\pi^{*}$ and $M^{*}$. A consistent estimator of $\theta^{*}$ can be obtained, for example, via the Method of Moments (MoM) approaches mentioned in the Introduction. When $d=1$, the denoising MoM estimator proposed in [WuYang2020] achieves a convergence rate of $(1/\pi_{\min}^{*}+1/\Delta_{\min})n^{-1/\mathcal{O}(L)}$. This exponential dependence on $L$ is unavoidable in the absence of any separation condition, and can be improved as the separation increases. For the multivariate $L$-GMM, a fast and consistent MoM estimation is proposed in [LindsayBasak], though no explicit convergence rates are provided. Recently, for higher dimensions ($d\geq 2$) and $L=2$, kalai2010efficiently proposed a MoM procedure that provably recovers the mixture parameters under the minimal separation. This framework was later extended to $L\geq 3$ by moitra2010settling. While their estimation errors exhibit only polynomial dependence on $d$, $1/\pi_{\min}^{*}$, and $1/\Delta_{\min}$, the degree of such polynomial is high, and the required sample complexity still grows exponentially in $L$. More recently, doss2023optimal developed an estimator for the mixing measure that is minimax optimal in the multivariate isotropic $L$-GMM setting. In principle, this estimator could be used to construct a consistent estimator of $\theta^{*}$ with lower-order dependence on $d$, but this direction is not pursued in their work. For general anisotropic $L$-GMMs with $L\geq 3$, designing a provably consistent and computationally efficient MoM estimator with low sample complexity remains an open and important direction for future research.
In practice, one could initialize EM with different preliminary label estimates obtained from other clustering algorithms, and then select the final estimator that yields the highest likelihood. In our simulation studies below, we adopt initialization scheme (a) for simplicity.
5.2 Simulation studies
In this section we first examine the decreasing behavior per iteration of the EM algorithm by tracking its optimization error and statistical error. Next, we evaluate how the convergence rate of the EM algorithm depends on the minimal mixing proportion $\pi_{\min}^{*}$, the smallest separation among all Gaussian components $\Delta_{\min}$, and whether the covariance matrix $\Sigma^{*}$ is known. Third, we track the empirical convergence rate by varying the sample size $n$, and then compare it with our theoretical minimax optimal rate. In all cases, the data are generated from model (1.1).
We initialize the EM algorithm as $\widehat{\pi}^{(0)}=0.7\pi^{*}+0.3\,\text{Dir}(5)$, $\widehat{\mu}_{\ell}^{(0)}=\mu_{\ell}^{*}+\nu_{\ell}$, for $\ell\in[L]$ and $\widehat{\Sigma}^{(0)}=\Sigma^{*}+(0.2\cdot{0.4^{2}}/d)\,AA^{\top}$ where $\text{Dir}(5)$ is the symmetric Dirichlet distribution, $\nu_{\ell}$ is drawn uniformly from the sphere of radius $r$ in $\mathbb{R}^{d}$, and $A\in\mathbb{R}^{d\times d}$ has i.i.d. entries from $\mathcal{N}(0,1)$. We set $r=0.4$ for the first two settings below while $r=0.2$ for the last setting.
Optimization error and statistical error
[…]
Convergence rate with parameters
We now investigate how the performance of EM algorithms depends on various parameters such as $\Delta_{\min}$. We fix $n=10,000,d=10$ and $L=3$. We set the mean vectors as $\mu_{1}=\Delta_{0}/\sqrt{2}{\bm{e}}_{1}$, $\mu_{2}=\Delta_{0}/\sqrt{2}{\bm{e}}_{7}$, and $\mu_{3}=\Delta_{0}/\sqrt{2}{\bm{e}}_{10}$, with $\Sigma^{*}=0.4^{2}\mathbf{I}_{d}$. Since the covariance is isotropic, this choice is equivalent to using $\{{\bm{e}}_{1},{\bm{e}}_{2},{\bm{e}}_{3}\}$, and we adopt it for numerical stability when varying $\Delta_{0}$ within $\{1.2,1.4,1.6,1.8,2.0\}$. We perform 10 independent trials for each case. The averages of $d(\widehat{M}^{(t)},M^{*})$ and $d(\widehat{\Sigma}^{(t)},\Sigma^{*})$ are plotted versus iterations. The left plots of Fig. 3 correspond to the balanced case $\pi=(1/3,1/3,1/3)$, while the right plots correspond to the imbalanced case $\pi=(0.6,0.2,0.2)$. We observe that the statistical error decreases initially and then reaches a plateau after several iterations. As expected, a larger $\Delta_{\min}$ yields a faster convergence rate. Comparing the left and right plots, a smaller $\pi_{\min}^{*}$ again results in a slower convergence rate. Finally, comparing the first row (known $\Sigma^{*}$) with the second row (unknown $\Sigma^{*}$), we see that the known $\Sigma^{*}$ case converges faster. Taken together, all of these findings are consistent with Theorem 2 and 3.
Figure 2: Plots of the iteration number versus log optimization error $)\log(d(\widehat{M}^{(t)},\widehat{M}$)) and log statistical error $\log(d(\widehat{M}^{(t)},M^{*}))$. Left: balanced mixing probability $\pi^{*}=(1/3,1/3,1/3)$. Right: imbalanced mixing probability $\pi^{*}=(0.6,0.2,0.2)$.
Figure 3: Convergence of EM on the mixture of $L=3$ Gaussians in $\mathbb{R}^{10}$. The first row (known $\Sigma^{*}$) and the second row (unknown $\Sigma^{*}$) show average mean estimation statistical error under different separations; the bottom row (unknown $\Sigma^{*}$) shows average covariance estimation statistical error. Left: balanced mixing probability $\pi^{*}=(1/3,1/3,1/3)$. Right: imbalanced mixing probability $\pi^{*}=(0.6,0.2,0.2)$.
Minimax Optimal Convergence Rate vs. Empirical Rate
In this section, we compare the empirical rates with the minimax optimal rates. Using similar data generation and initialization as before, we fix $L=5$ and $d=50$, set $\mu_{\ell}=2\sqrt{2}{\bm{e}}_{\ell}$ and $\pi_{\ell}=1/L$ for all $\ell\in[L]$, and vary $n$ over the sequence from 6,000 to 40,000 with equal increments of 2,000. We compare the isotropic model $\Sigma^{*}=0.4^{2}\mathbf{I}_{d}$ with the compound symmetry model $\Sigma^{*}=0.6\cdot\mathbf{I}_{d}+0.4\cdot\mathbf{1}\mathbf{1}^{\top}$. The optimal rates (see Table 1) are $d(\widehat{M},M^{*})\asymp\sqrt{d/(n\pi_{\min}^{*})}$ and $d(\widehat{\Sigma},\Sigma^{*})\asymp\sqrt{d/n}$. We then plot the empirical convergence rates against the optimal rates and fit a line through the origin. The empirical convergence rates under the two models exhibit similar behavior. The $R^{2}$ values for all cases are greater than $0.99$. The near-linear alignment in Fig. 4 thus indicates that the observed rates closely follow the minimax theoretical rates.
Figure 4: Empirical vs. optimal rates. Left: isotropic model; Right: compound symmetry model. Top row: mean; Bottom row: covariance.
Appendix A contains the full comparison table with existing theoretical works of EM. Lemma 1 is proved in Appendix B. The proofs of population-level convergence of the EM algorithm are stated in Appendix C. Appendix D contains the proofs of the sample-level results in Section 4.1 when the separation is small to moderate while Appendix E contains the proofs when the separation is moderate to large. Appendix F contains the proof of the minimax lower bounds. Auxiliary lemmas are collected in Appendix G.
Appendix A A full comparison table with existing theoretical works of the EM algorithm under the $L$-GMM
Table 3: Comparison with existing theoretical results for the EM algorithm under the $L$-GMM. Symmetric 2-GMM means $\pi_{1}=\pi_{2}=1/2$ and $\mu_{1}=-\mu_{2}=\mu$. In [cai2019chime], $\beta=\Sigma^{-1}(\mu_{2}-\mu_{1})$.
$L$ Parameter settings Separation $\Delta_{\min}$ Initialization Convergence Rates
[EM2017] 22 symmetric, $\Sigma=\mathbf{I}_{d}$ >C>C $\|\mu^{(0)}-\mu\|_{2}\leq\|\mu\|_{2}/4$ $d(\widehat{M},M)\lessapprox\|\mu\|_{2}^{3}\sqrt{d/n})$
[Xu_Hsu_Maleki] 22 symmetric, $\Sigma$ known >0>0 $\mu^{(0)\top}\mu\neq 0$ $d(\widehat{M},M)=o(1)$
[daskalakis2017ten] 22 symmetric, $\Sigma$ known $\widetilde{\Omega}((d/n)^{1/3})$ $\mu^{(0)\top}\mu\neq 0$ $d(\widehat{M},M)\lessapprox(\sqrt{d/n}\wedge\|\mu\|_{2}^{-2}\sqrt{d/n})$
[wu2021randomly] 22 symmetric, $\Sigma=\mathbf{I}_{d}$ $(0,C]$ $\mu^{(0)\top}\mu\neq 0$ $d(\widehat{M},M)\lessapprox\left((d/n)^{1/4}\wedge\|\mu\|_{2}^{-1}\sqrt{d/n}\right)$
[weinberger2022algorithm] 22
$\pi_{2}>\pi_{1}$ known, $\Sigma=\mathbf{I}_{d}$
$\mu_{2}=-\mu_{1}=\mu$
$(0,C]$ $\mu^{(0)\top}\mu\geq 0$ $d(\widehat{M},M)\lessapprox\left((d/n)^{1/4}\wedge\|\mu\|_{2}^{-1}\sqrt{d/n}\wedge(\pi_{2}-\pi_{1})^{-1}\sqrt{d/n}\right)$
[cai2019chime] 22
$c\leq\pi_{1},\pi_{2}\leq 1-c$
$c\prec\Sigma\prec C$
$C,C^{\prime}]$
$d(\pi^{(0)},\pi)+d(M^{(0)},M)$
$\|\beta^{(0)}-\beta\|_{2}\leq c\sqrt{\Delta_{\min}/d}$
$d(\widehat{M},M)+\|\widehat{\beta}-\beta\|_{2}\lesssim\sqrt{d/n}$
[dasgupta2007probabilistic] $\geq 2$ $\Sigma=\sigma^{2}\mathbf{I}_{d}$ $\Omega\Bigl(\sqrt{d\log\pi_{\min}^{-1}}+\log\pi_{\min}^{-1}\Bigr)$ A particular initialization $d(\widehat{\pi},\pi)+d(\widehat{M},M)\lesssim\left(\sqrt{d/(n\pi_{\min})}+{\exp(-c\sqrt{d})/\pi_{\min}}\right)$
[kwon2020algorithm] $\geq 2$ $\Sigma=\sigma^{2}\mathbf{I}_{d}$ $\log(1/\pi_{\min})$ $d(\pi^{(0)},\pi)<1,~d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ A sample-splitting version of the EM
[yan2017convergence] $\geq 2$ $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $\Omega\Bigl((d\wedge L)\log{(\Delta_{\max}/\pi_{\min})}\Bigr)$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)\lessapprox{(\pi_{\max}/\pi_{\min})}\sqrt{\Delta_{\max}}(\sqrt{d}+L^{3}\Delta_{\max})\sqrt{d/n}$
[Zhao2020] $\geq 2$ $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $\Omega\Bigl((d\wedge L)\log{(\Delta_{\max}/\pi_{\min})}\Bigr)$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)\lessapprox({\sqrt{\Delta_{\max}}/\pi_{\min}})\sqrt{dL/n}$
[segol2021improved] $\geq 2$ $\pi$ known, $\Sigma=\mathbf{I}_{d}$ $\log(1/\pi_{\min})$ $d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{M},M)\lessapprox(1/\pi_{\min})\sqrt{dL/n}$
This work $\geq 2$ Full generality $\Omega\Bigl(\log{(\Delta_{\max}/\pi_{\min})}\Bigr)$ $d(\pi^{(0)},\pi)<1,~d(M^{(0)},M)\leq c\sqrt{\Delta_{\min}}$ $d(\widehat{\pi},\pi)+d(\widehat{M},M)\lessapprox\sqrt{d/(n\pi_{\min})}$
$d(\Sigma^{(0)},\Sigma)\leq c$, or via preliminary labels
$d(\widehat{\Sigma},\Sigma)\lessapprox\sqrt{d/n}$
Appendix B Proof of Lemma 1
We first state and proof a simple lemma that is used in proving Lemma 1. Let $\Gamma$ be any $n\times L$ matrix with each row $\Gamma_{i\cdot}\in\mathcal{S}^{L}$ for all $i\in[n]$. Here $\mathcal{S}^{L}$ denotes the $L$-dimensional probability simplex. Let $P_{\mathbf{X}}\in\mathbb{R}^{n\times n}$ denote the orthogonal projection matrix onto the column space of ${\mathbf{X}}:=(X_{1},X_{2},\ldots,X_{n})^{\top}\in\mathbb{R}^{n\times d}$. Write $P_{\mathbf{X}}^{\perp}=\mathbf{I}_{n}-P_{\mathbf{X}}$.
Lemma 16.
If there exists at least one $i\in[n]$ such that $\Gamma_{i\ell}>0$ for all $\ell\in[L]$, then the matrix
${\rm diag}(\Gamma^{\top}1_{n})-\Gamma^{\top}P_{\mathbf{X}}~\Gamma$
is strictly positive definite.
Proof.
Start with
${\rm diag}(\Gamma^{\top}1_{n})-\Gamma^{{}^{\top}}P_{{\mathbf{X}}}\Gamma={\rm diag}(\Gamma^{\top}1_{n})-\Gamma^{{}^{\top}}\Gamma+\Gamma^{{}^{\top}}P_{{\mathbf{X}}}^{\perp}\Gamma.$
It remains to show that ${\rm diag}(\Gamma^{\top}1_{n})-\Gamma^{{}^{\top}}\Gamma$ is strictly positive definite. For any non-zero vector $v\in\mathbb{R}^{L}$, we have
$v^{\top}\left({\rm diag}(\Gamma^{\top}1_{n})-\Gamma^{{}^{\top}}\Gamma\right)v=$ $\sum_{i=1}^{n}v^{\top}\left({\rm diag}(\Gamma_{i\cdot})-\Gamma_{i\cdot}\Gamma_{i\cdot}^{\top}\right)v.$
Since $\Gamma_{i\cdot}\in\mathcal{S}^{L}$ for all $i\in[n]$, the above is always non-negative. Moreover, it is strictly greater than zero as long as there exists at least one $i\in[n]$ such that $\Gamma_{i\ell}>0$ for all $\ell\in[L]$. This completes the proof. ∎
Proof of Lemma 1.
For any $t\geq 0$ with any $\widehat{\theta}^{(t)}$, write
$\widehat{\Gamma}^{(t)}:=\Gamma({\mathbf{X}},\widehat{\theta}^{(t)}):=\begin{bmatrix}\gamma_{1}(X_{1};\widehat{\theta}^{(t)})&\cdots&\gamma_{L}(X_{1};\widehat{\theta}^{(t)})\\ \vdots&\ddots&\vdots\\ \gamma_{1}(X_{n};\widehat{\theta}^{(t)})&\cdots&\gamma_{L}(X_{n};\widehat{\theta}^{(t)})\end{bmatrix}\in\mathbb{R}^{n\times L}.$ (B.1)
Then using Eqs. 2.8 and 2.9 gives that
$\widehat{M}^{(t)}={\mathbf{X}}\widehat{\Gamma}^{(t)}{\rm diag}^{-1}(\widehat{\Gamma}^{(t)\top}1_{n}),\qquad\widehat{\pi}^{(t)}={1\over n}\widehat{\Gamma}^{(t)\top}1_{n}$
so that (2.11) yields
$\widehat{\Sigma}^{(t)}$ $=\widehat{\Sigma}_{T}-\widehat{M}^{(t)}{\rm diag}(\widehat{\pi}^{(t)})\widehat{M}^{(t)\top}$
$={1\over n}{\mathbf{X}}^{\top}{\mathbf{X}}-{1\over n}{\mathbf{X}}\widehat{\Gamma}^{(t)}{\rm diag}^{-1}(\widehat{\Gamma}^{(t)\top}1_{n})\widehat{\Gamma}^{(t)\top}{\mathbf{X}}^{\top}$
$={1\over n}{\mathbf{X}}^{\top}\left[\mathbf{I}_{n}-P_{{\mathbf{X}}}^{1/2}\widehat{\Gamma}^{(t)}{\rm diag}^{-1}(\widehat{\Gamma}^{(t)\top}1_{n})\widehat{\Gamma}^{(t)\top}P_{{\mathbf{X}}}^{1/2}\right]{\mathbf{X}}.$
Observe that the matrix $H:=P_{{\mathbf{X}}}^{1/2}\widehat{\Gamma}^{(t)}{\rm diag}^{-1}(\widehat{\Gamma}^{(t)\top}1_{n})\widehat{\Gamma}^{(t)\top}P_{{\mathbf{X}}}^{1/2}$ is positive semi-definite and has rank at most $L$. Moreover, for any $1\leq r\leq L$, its the $r$th largest eigenvalue satisfies
$0\leq\lambda_{r}(H)=\lambda_{r}\left({\rm diag}^{-1/2}(\widehat{\Gamma}^{(t)\top}1_{n})\widehat{\Gamma}^{(t)\top}P_{{\mathbf{X}}}\widehat{\Gamma}^{(t)}{\rm diag}^{-1/2}(\widehat{\Gamma}^{(t)\top}1_{n})\right)<1$
by using Lemma 16 and the fact that each row of $\widehat{\Gamma}^{(t)}$ has $L$ non-zero entries. Therefore $(\mathbf{I}_{n}-H)$ has rank $n$ and the invertibility of $\widehat{\Sigma}^{(t)}$ follows from that of $\widehat{\Sigma}_{T}$, completing the proof. ∎
Appendix C Proofs of Section 3: convergence of the population level EM algorithm
Notation. Throughout the proof, we write
$\Delta_{k\ell}=\|\mu_{k}^{*}-\mu_{\ell}^{*}\|_{\Sigma^{*}}^{2}=(\mu_{k}^{*}-\mu_{\ell}^{*})^{\top}\Sigma^{*-1}(\mu_{k}^{*}-\mu_{\ell}^{*}),\qquad\forall~k,\ell\in[L].$ (C.1)
Recall $\Delta_{\max}$ and $\Delta_{\min}$ from (3.1). Due to $\mathbb{E}[X]=0$, we have
${1\over 4}\Delta_{\max}\leq\max_{\ell\in[L]}~\mu_{\ell}^{*\top}\Sigma^{*-1}\mu_{\ell}^{*}\leq\Delta_{\max}.$ (C.2)
For any $\pi=(\pi_{1},\ldots,\pi_{L})^{\top}$, we write $\pi_{\min}=\min_{\ell\in[L]}\pi_{\ell}.$
Recall that given any parameter $\theta$, the population level M-step in Eqs. 2.3, 2.4 and 2.5 are
$\bm{\pi}_{\ell}(\theta)=\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta)],\qquad\bm{\mu}_{\ell}(\theta)={\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta)X]\over\mathbb{E}_{\theta^{*}}[\gamma_{\ell}(X;\theta)]},\qquad\text{for all $\ell\in[L]$,}$ (C.3)
and, by collecting $\bm{\pi}(\theta)=(\bm{\pi}_{1}(\theta),\ldots,\bm{\pi}_{L}(\theta))^{\top}$ and $\mathbf{M}(\theta)=(\bm{\mu}_{1}(\theta),\ldots,\bm{\mu}_{L}(\theta))$,
$\bm{\Sigma}(\theta)=\mathbb{E}_{\theta^{*}}[XX^{\top}]-\mathbf{M}(\theta){\rm diag}(\bm{\pi}(\theta))\mathbf{M}(\theta)^{\top}.$ (C.4)
Note that they all depend on $\gamma_{\ell}(X;\theta)$. We start by rewriting
$\gamma_{\ell}(x;\theta)=\frac{\pi_{\ell}}{\sum_{k=1}^{L}\pi_{k}\exp\left\{(x-{1\over 2}(\mu_{\ell}+\mu_{k}))^{\top}J(\theta)({\bm{e}}_{k}-{\bm{e}}_{\ell})\right\}}:=\gamma_{\ell}(x;\omega(\theta))$ (C.5)
for any $x\in\mathbb{R}^{d}$ and $\ell\in[L]$, where we introduce
$\omega(\theta)=(\pi,M,J(\theta))$ (C.6)
with $J(\theta):=\Sigma^{-1}M$.
C.1 Self-consistency of the (population level) EM iterates
The so-called self-consistency of the above update is easy to verify: for any $k\in[L]$, we have
$\bm{\pi}_{k}(\theta^{*})=\mathbb{E}_{\theta^{*}}[\gamma_{k}(X;\theta^{*})]=\mathbb{E}_{\theta^{*}}[\mathbb{P}_{\theta^{*}}(Y=k\mid X)]=\mathbb{P}_{\theta^{*}}(Y=k)=\pi_{k}^{*},$
$\bm{\mu}_{k}(\theta^{*})=\frac{\mathbb{E}_{\theta^{*}}[\gamma_{k}(X;\theta^{*})X]}{\mathbb{E}_{\theta^{*}}[\gamma_{k}(X;\theta^{*})]}=\frac{\mathbb{E}_{\theta^{*}}[\mathbb{P}_{\theta^{*}}(Y=k\mid X)X]}{\pi_{k}^{*}}=\frac{\mathbb{E}_{\theta^{*}}[1\{Y=k\}X]}{\pi_{k}^{*}}=\frac{\pi_{k}^{*}\mu^{*}_{k}}{\pi_{k}^{*}}=\mu_{k}^{*}.$
As a result,
$\bm{\Sigma}(\theta^{*})=\Sigma_{T}^{*}-M^{*}D_{\pi^{*}}M^{*\top}=\Sigma^{*}$ (C.7)
and $J(\theta^{*})=\Sigma^{*-1}M^{*}=:J^{*}.$ In the subsequent proofs, we drop the subscripts $\theta^{*}$ in $\mathbb{E}_{\theta^{*}}$ and $\mathbb{P}_{\theta^{*}}$ when there is no confusion.
C.2 A local neighborhood of $\omega^{*}$ and Lipschitz property of $J(\cdot)$ and $\Sigma(\cdot)$
For any $\omega=(\pi,M,J)$ and $\omega^{\prime}=(\pi^{\prime},M^{\prime},J^{\prime})$, by recalling $d(\pi,\pi^{\prime})$ and $d(M,M^{\prime})$ from (3.2), define
$d(J,J^{\prime}):={1\over 2}\max_{a,\ell\in[L]}\|\Sigma^{*1/2}(J-J^{\prime})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$ (C.8)
and the distance between $\omega$ and $\omega^{\prime}$ as
$d(\omega,\omega^{\prime}):=d(\pi,\pi^{\prime})+\sqrt{\Delta_{\max}}~d(M,M^{\prime})+\sqrt{\Delta_{\max}}~d(J,J^{\prime}).$ (C.9)
For some constant $0<c_{0}<(\sqrt{2}-1)/2$, denote the size-$c_{0}$ neighborhood of $\omega^{*}$ by
$\mathcal{B}_{d}(\omega^{*})\equiv\mathcal{B}_{d}(\omega^{*},c_{0})=$ $\left\{(\pi,M,J):d(\pi,\pi^{*})\leq{1\over 2},~{d(M,M^{*})\over\sqrt{\Delta_{\min}}}\leq c_{0},\right$.
$\hskip 28.45274pt\left.~\max_{a\neq\ell}{1\over\sqrt{\Delta_{a\ell}}}\left\|\Sigma^{*1/2}(J-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}\leq 2c_{0}\right\}.$ (C.10)
For future reference, Lemma 17 below ensures that for any $\theta$ satisfying (3.5),
${1\over\sqrt{\Delta_{a\ell}}}\left\|\Sigma^{*1/2}(J(\theta)-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}$ $\leq{\|\mu_{a}-\mu_{a}^{*}\|_{\Sigma^{*}}+\|\mu_{\ell}-\mu_{\ell}^{*}\|_{\Sigma^{*}}\over(1-c_{\Sigma})\sqrt{\Delta_{a\ell}}}+{c_{\Sigma}\over 1-c_{\Sigma}}$
$\leq{2c_{\mu}+c_{\Sigma}\over 1-c_{\Sigma}}$
holds for all $a\neq\ell$. Consequently, its corresponding $\omega(\theta)\in\mathcal{B}_{d}(\omega^{*},c_{0})$ with $c_{0}=(c_{\mu}+c_{\Sigma}/2)/(1-c_{\Sigma})$ and
$d(\omega(\theta),\omega^{*})\leq{1\over 2}+c_{0}\sqrt{\Delta_{\max}\Delta_{\min}}+c_{0}\Delta_{\max}\leq{1\over 2}+2c_{0}\Delta_{\max}.$ (C.11)
Lemma 17.
Fix any $\theta=(\pi,M,\Sigma)$ with $d(\Sigma,\Sigma^{*})<1$. For any $a\neq\ell\in[L]$, one has
$\|\Sigma^{*1/2}(J(\theta)-J(\theta^{*}))({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}\leq{\|\mu_{a}-\mu_{a}^{*}\|_{\Sigma^{*}}+\|\mu_{\ell}-\mu_{\ell}^{*}\|_{\Sigma^{*}}\over 1-d(\Sigma,\Sigma^{*})}+{d(\Sigma,\Sigma^{*})\over 1-d(\Sigma,\Sigma^{*})}\sqrt{\Delta_{a\ell}}.$
Moreover, if $d(\Sigma,\Sigma^{*})\leq 1/2$, then
$d(J(\theta),J(\theta^{*}))\leq 2d(M,M^{*})+\sqrt{\Delta_{\max}}~d(\Sigma,\Sigma^{*}).$
Proof.
Weyl’s inequality ensures that
$1-d(\Sigma,\Sigma^{*})\leq\lambda_{j}(\Sigma^{*-1/2}\Sigma\Sigma^{*-1/2})\leq 1+d(\Sigma,\Sigma^{*}),\quad\forall~1\leq j\leq d.$ (C.12)
Thus $\Sigma$ is invertible. The first result follows from
$\|\Sigma^{*1/2}(J(\theta)-J(\theta^{*}))({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$
$=\left\|\Sigma^{*1/2}\left(\Sigma^{-1}(M-M^{*})+\Sigma^{-1}(\Sigma^{*}-\Sigma)\Sigma^{*-1}M^{*}\right)({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}$
$\leq\|\Sigma^{*-1/2}\Sigma^{-1}\Sigma^{*-1/2}\|_{\rm op}\left\{\|(M-M^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{\Sigma^{*}}+d(\Sigma,\Sigma^{*})\|\mu_{a}^{*}-\mu_{\ell}^{*}\|_{\Sigma^{*}}\right\}$
$\leq{1\over 1-d(\Sigma,\Sigma^{*})}\left\{\|\mu_{a}-\mu_{a}^{*}\|_{\Sigma^{*}}+\|\mu_{\ell}-\mu_{\ell}^{*}\|_{\Sigma^{*}}+d(\Sigma,\Sigma^{*})\sqrt{\Delta_{a\ell}}\right\}$ (C.13)
where the last step uses (C.12). The second result follows trivially from the first one. ∎
For $\Sigma=\Sigma_{T}-M{\rm diag}(\pi)M^{\top}$ and $\Sigma^{\prime}=\Sigma_{T}^{\prime}-M^{\prime}{\rm diag}(\pi^{\prime})M^{{}^{\prime}\top},$ the following lemma relates $d(\Sigma,\Sigma^{\prime})$ to $d(\pi,\pi^{\prime})$, $d(M,M^{\prime})$ and $d(\Sigma_{T},\Sigma_{T}^{\prime})$ with
$d(\Sigma_{T},\Sigma_{T}^{\prime}):=\|\Sigma^{*-1/2}(\Sigma_{T}-\Sigma_{T}^{\prime})\Sigma^{*-1/2}\|_{\rm op}.$
Lemma 18.
Fix any $\theta=(\pi,M,\Sigma)$ with $\Sigma=\Sigma_{T}-M{\rm diag}(\pi)M^{\top}$ and $\theta^{\prime}=(\pi^{\prime},M^{\prime},\Sigma^{\prime})$ with $\Sigma^{\prime}:=\Sigma_{T}^{\prime}-M^{\prime}{\rm diag}(\pi^{\prime})M^{{}^{\prime}\top}.$
  1. (a)
    Taking $\theta^{\prime}=\theta^{*}$, if $d(M,M^{\prime})\leq\|\Sigma^{*-1/2}M^{\prime}\|_{2,\infty}$, then
    $d(\Sigma,\Sigma^{\prime})\leq 3\|\Sigma^{*-1/2}M^{\prime}\|_{2,\infty}d(M,M^{\prime})+2\|\Sigma^{*-1/2}M^{\prime}\|_{2,\infty}^{2}d(\pi,\pi^{\prime})+d(\Sigma_{T},\Sigma_{T}^{\prime}).$ (C.14)
  2. (b)
    If
    $6\sqrt{\Delta_{\max}}d(M,M^{*})+4\Delta_{\max}d(\pi,\pi^{*})+2d(\Sigma_{T},\Sigma_{T}^{*})\leq 1,$ (C.15)
    we have
    $d(J(\theta),J(\theta^{*}))\leq 8\Delta_{\max}d(M,M^{*})+4\Delta_{\max}^{3/2}d(\pi,\pi^{*})+2\sqrt{\Delta_{\max}}d(\Sigma_{T},\Sigma_{T}^{*}).$
Proof.
Part (a) follows by
$d(\Sigma,\Sigma^{\prime})$ $\leq\|\Sigma^{*-1/2}M^{\prime}{\rm diag}(\pi^{\prime})(M^{\prime}-M)^{\top}\Sigma^{*-1/2}\|_{\rm op}$
$\quad+\|\Sigma^{*-1/2}M^{\prime}({\rm diag}(\pi^{\prime})-{\rm diag}(\pi))M^{\top}\Sigma^{*-1/2}\|_{\rm op}$
$\quad+\|\Sigma^{*-1/2}(M^{\prime}-M){\rm diag}(\pi)M^{\top}\Sigma^{*-1/2}\|_{\rm op}+\|\Sigma^{*-1/2}(\Sigma_{T}-\Sigma_{T}^{\prime})\Sigma^{*-1/2}\|_{\rm op}$
$\leq\|\Sigma^{*-1/2}M^{\prime}\|_{2,\infty}d(M,M^{\prime})+d(\pi,\pi^{\prime})\|\Sigma^{*-1/2}M^{\prime}\|_{2,\infty}\left[\|\Sigma^{*-1/2}M^{\prime}\|_{2,\infty}+d(M,M^{\prime})\right]$
$\quad+d(M,M^{\prime})\left[\|\Sigma^{*-1/2}M^{\prime}\|_{2,\infty}+d(M,M^{\prime})\right]+d(\Sigma_{T},\Sigma_{T}^{\prime})$
$\leq 3\|\Sigma^{*-1/2}M^{\prime}\|_{2,\infty}d(M,M^{\prime})+2\|\Sigma^{*-1/2}M^{\prime}\|_{2,\infty}^{2}d(\pi,\pi^{\prime})+d(\Sigma_{T},\Sigma_{T}^{\prime}).$
The last step uses $d(M,M^{\prime})\leq\|\Sigma^{*-1/2}M^{\prime}\|_{2,\infty}$.
To prove part (b), since condition (C.15) ensures that $d(\Sigma,\Sigma^{*})\leq 1/2$ holds, the last claim then follows by plugging the above bound with $\theta^{\prime}=\theta^{*}$ into the result of Lemma 17. ∎
C.3 Proof of Theorem 2: contraction of the population level EM iterates
The following theorem states the one-step contraction property of the population level M-steps $\bm{\pi}(\omega)$, $\mathbf{M}(\omega)$ and $\bm{\Sigma}(\omega)$ for any $\omega\in\mathcal{B}_{d}(\omega^{*})$. Recall $d(\pi,\pi^{*})$ from (C.9).
Theorem 19.
Grant condition (3.4). For some absolute constants $c,C>0$, let
$\kappa_{\omega}=\frac{C}{\pi_{\min}^{*}}\exp{(-c\Delta_{\min})}.$ (C.16)
For any $\omega=(\pi,M,J)\in\mathcal{B}_{d}(\omega^{*})$, we have
$d(\bm{\pi}(\omega),\bm{\pi}(\omega^{*}))$ $~\leq~\kappa_{\omega}~d(\omega,\omega^{*}),$ (C.17)
$d(\mathbf{M}(\omega),\mathbf{M}(\omega^{*}))$ $~\leq~\kappa_{\omega}\sqrt{\Delta_{\max}}~d(\omega,\omega^{*}),$ (C.18)
$d(\bm{\Sigma}(\omega),\bm{\Sigma}(\omega^{*}))$ $~\leq~\kappa_{\omega}~\Delta_{\max}~d(\omega,\omega^{*}).$ (C.19)
Proof.
The proofs of (C.17) and (C.18) can be found in Section C.3.1 and Section C.3.2, respectively. The proof of (C.19) uses (C.14) of Lemma 18 with $\theta^{\prime}=\theta^{*}$ and $\Sigma_{T}=\Sigma_{T}^{*}$. ∎
Combining Theorem 19 and Lemma 17, we are ready to prove Theorem 2.
Proof of Theorem 2.
Fix any $\theta=(\pi,M,\Sigma)$ satisfying (3.5). By induction, it suffices to prove
$d(\bm{\pi}(\theta),\pi^{*})+d(\mathbf{M}(\theta),M^{*})+d(\bm{\Sigma}(\theta),\Sigma^{*})\leq\kappa\left\{d(\pi,\pi^{*})+d(M,M^{*})+d(\Sigma,\Sigma^{*})\right\}$ (C.20)
and the updated parameters $\bm{\pi}(\theta),\mathbf{M}(\theta),\bm{\Sigma}(\theta)$ still satisfy (3.5).
To prove (C.20), since condition (3.5) ensures $\omega=(\pi,M,J(\theta))\in\mathcal{B}_{d}(\omega^{*})$, one has
$d(\bm{\pi}(\theta),\pi^{*})$ $=d(\bm{\pi}(\omega(\theta)),\bm{\pi}(\omega(\theta^{*})))$
$\leq\kappa_{\omega}\left\{d(\pi,\pi^{*})+\sqrt{\Delta_{\max}}~d(M,M^{*})+\sqrt{\Delta_{\max}}~d(J(\theta),J^{*})\right\}$ by Theorem 19
$\leq\kappa_{\omega}\left\{d(\pi,\pi^{*})+3\sqrt{\Delta_{\max}}~d(M,M^{*})+{\Delta_{\max}}~d(\Sigma,\Sigma^{*})\right\}$ by Lemma 17
with $\kappa_{\omega}$ given by (C.16). Similarly, we also have
$d(\mathbf{M}(\theta),M^{*})$ $~\leq~\kappa_{\omega}\sqrt{\Delta_{\max}}\left\{d(\pi,\pi^{*})+3\sqrt{\Delta_{\max}}~d(M,M^{*})+{\Delta_{\max}}~d(\Sigma,\Sigma^{*})\right\}.$
By Lemma 18 with $\Sigma_{T}=\Sigma_{T}^{*}$, we further have
$d(\bm{\Sigma}(\theta),\Sigma^{*})$~ $\leq~3\sqrt{\Delta_{\max}}~d(\mathbf{M}(\theta),M^{*})+2\Delta_{\max}d(\bm{\pi}(\theta),\theta^{*})$
$~\leq~5\kappa_{\omega}~\Delta_{\max}\left\{d(\pi,\pi^{*})+3\sqrt{\Delta_{\max}}~d(M,M^{*})+{\Delta_{\max}}~d(\Sigma,\Sigma^{*})\right\},$
thereby proving (C.20) with $\kappa=21\Delta_{\max}^{2}\kappa_{w}$. Furthermore, since condition (3.4) ensures
$d(\bm{\pi}(\theta),\pi^{*})$ $\leq\kappa_{w}\left\{{1\over 2}+3c_{\mu}\sqrt{\Delta_{\max}\Delta_{\min}}+\Delta_{\max}c_{\Sigma}\right\}\leq{1\over 2},$
$d(\mathbf{M}(\theta),M^{*})$ $~\leq~\kappa_{\omega}\sqrt{\Delta_{\max}}\left\{{1\over 2}+3c_{\mu}\sqrt{\Delta_{\max}\Delta_{\min}}+{\Delta_{\max}}c_{\Sigma}\right\}\leq c_{\mu}\sqrt{\Delta_{\min}},$
$d(\bm{\Sigma}(\theta),\Sigma^{*})$ $\lesssim 5\kappa_{\omega}~\Delta_{\max}\left\{{1\over 2}+3c_{\mu}\sqrt{\Delta_{\max}\Delta_{\min}}+{\Delta_{\max}}c_{\Sigma}\right\}\leq c_{\Sigma},$
we conclude that $\bm{\pi}(\theta),\mathbf{M}(\theta)$ and $\bm{\Sigma}(\theta)$ still satisfy condition (3.5) so that using induction completes the proof. ∎
C.3.1 Proof of Eq. C.17 of Theorem 19: Contraction of the population level M-step of the mixing probabilities
Proof.
Fix any $\bar{\omega}=(\bar{\pi},\bar{M},\bar{J})\in\mathcal{B}_{d}(\omega^{*})$ given by (C.2) and pick any $\ell\in[L]$. Our goal is to bound from above
$\left|\bm{\pi}_{\ell}(\bar{\omega})-\bm{\pi}_{\ell}(\omega^{*})\right|=\left|\mathbb{E}[\gamma_{\ell}(X;\omega^{*})]-\mathbb{E}[\gamma_{\ell}(X;\bar{\omega})]\right|.$
For any $u\in(0,1)$, by writing
$\pi_{u}=(1-u)\pi^{*}+u\bar{\pi},\quad M_{u}=(1-u)M^{*}+u\bar{M},\quad J_{u}=(1-u)J^{*}+u\bar{J},$
we have
$\bm{\pi}_{\ell}(\bar{\omega})-\bm{\pi}_{\ell}(\omega^{*})=\nabla^{\pi}\bm{\pi}_{\ell}(\bar{\omega})+\nabla^{M}\bm{\pi}_{\ell}(\bar{\omega})+\nabla^{J}\bm{\pi}_{\ell}(\bar{\omega})$
where
$\begin{split}\nabla^{\pi}\bm{\pi}_{\ell}(\bar{\omega})&:=\sum_{k=1}^{L-1}\int_{0}^{1}\mathbb{E}\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial\pi_{k}}\Big|_{\omega=(\pi_{u},\bar{M},\bar{J})},\bar{\pi}_{k}-\pi_{k}^{*}\right\rangle{\rm d}u\\ \nabla^{M}\bm{\pi}_{\ell}(\bar{\omega})&:=\sum_{k=1}^{L}\int_{0}^{1}\mathbb{E}\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial\mu_{k}}\Big|_{\omega=(\pi^{*},M_{u},\bar{J})},\bar{\mu}_{k}-\mu_{k}^{*}\right\rangle{\rm d}u\\ \nabla^{J}\bm{\pi}_{\ell}(\bar{\omega})&:=\int_{0}^{1}\mathbb{E}\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial J}\Big|_{\omega=(\pi^{*},M^{*},J_{u})},\bar{J}-J^{*}\right\rangle{\rm d}u.\end{split}$ (C.21)
Below we bound each terms separately in terms of
$\begin{split}\mathcal{R}_{I}^{(0)}&:=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)\right],\\ \mathcal{R}_{I}^{(1)}&:=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}~\Bigl\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)NN^{\top}\right]\Bigr\|_{\rm op},\\ \mathcal{R}_{I}^{(2)}&:=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\left\|\mathbb{E}\left[I_{a\ell}(X,\omega)NN^{\top}\right]\right\|_{\rm op},\end{split}$ (C.22)
where, for any $a,\ell\in[L]$ and $x\in\mathbb{R}^{L}$, we write
$I_{a\ell}(x,\omega)$ $:=\frac{\exp\{(x-\frac{1}{2}(\mu_{\ell}+\mu_{a}))^{\top}J({\bm{e}}_{a}-{\bm{e}}_{\ell})\}}{(\pi_{\ell}+\sum_{k\neq\ell}\pi_{k}\exp\{(x-\frac{1}{2}(\mu_{\ell}+\mu_{k}))^{\top}J({\bm{e}}_{k}-{\bm{e}}_{\ell}\})^{2}}.$ (C.23)
Clearly, we have $\mathcal{R}_{I}^{(1)}\leq\mathcal{R}_{I}^{(2)}$.
Controlling the partial derivative w.r.t. $\pi_{k}$’s.
We bound from above
$|\nabla^{\pi}\bm{\pi}_{\ell}(\bar{\omega})|~\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left|\sum_{k=1}^{L-1}\mathbb{E}\left[\frac{\partial\gamma_{\ell}(X;\omega)}{\partial\pi_{k}}\bigg|_{\omega}\right]\left(\bar{\pi}_{k}-\pi_{k}^{*}\right)\right|.$
Pick any $\omega\in\mathcal{B}_{d}(\omega^{*})$. Recall that for any $x\in\mathbb{R}^{d}$ and $\ell\in[L]$,
$\gamma_{\ell}(x;\omega)=\frac{\pi_{\ell}}{\sum_{k=1}^{L}\pi_{k}\exp\left\{(x-{1\over 2}(\mu_{\ell}+\mu_{k}))^{\top}J({\bm{e}}_{k}-{\bm{e}}_{\ell})\right\}}.$
Since $\pi_{L}=1-\sum_{\ell=1}^{L-1}\pi_{\ell}$, basic calculation gives that: for all $1\leq k,\ell\leq L-1$,
$\frac{\partial\gamma_{\ell}(x;\omega)}{\partial\pi_{k}}$ $=\pi_{\ell}I_{L\ell}(x,\omega)-\pi_{\ell}I_{k\ell}(x,\omega),\qquad k\neq\ell,$ (C.24)
$\frac{\partial\gamma_{\ell}(x;\omega)}{\partial\pi_{\ell}}$ $=\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}I_{a\ell}(x,\omega)+\pi_{\ell}I_{L\ell}(x,\omega),$ (C.25)
$\frac{\partial\gamma_{L}(x;\omega)}{\partial\pi_{k}}$ $=\sum_{a=1}^{L-1}\pi_{a}I_{aL}(x,\omega)-\pi_{L}I_{kL}(x,\omega),$ (C.26)
so that for all $1\leq k,\ell<L$ with $k\neq\ell$,
$|\nabla^{\pi}\bm{\pi}_{\ell}(\bar{\omega})|$ $~\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left|\sum_{k<L,k\neq\ell}\mathbb{E}\left[\pi_{\ell}I_{L\ell}(X,\omega)-\pi_{\ell}I_{k\ell}(X,\omega)\right]\left(\bar{\pi}_{k}-\pi_{k}^{*}\right)\right$.
$\qquad\qquad+\left.\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)\right]\left(\bar{\pi}_{\ell}-\pi_{\ell}^{*}\right)+\pi_{\ell}\mathbb{E}\left[I_{L\ell}(X,\omega)\right]\left(\bar{\pi}_{\ell}-\pi_{\ell}^{*}\right)\right|$
$~\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\Bigl\{(\pi_{\ell}+\pi_{\ell}^{*})\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}[I_{a\ell}(X,\omega)]\Bigr\}\max_{k<L}\frac{|\bar{\pi}_{k}-\pi_{k}^{*}|}{\pi_{k}^{*}}$
$\qquad+\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\pi_{\ell}\mathbb{E}\left[I_{L\ell}(X,\omega)\right]\left|\sum_{k<L}\left(\bar{\pi}_{k}-\pi_{k}^{*}\right)\right|$
$~\lesssim~\pi_{\ell}^{*}~\mathcal{R}_{I}^{(0)}d(\bar{\pi},\pi^{*}).$ (C.27)
In the last step, we use $\sum_{k<L}(\bar{\pi}_{k}-\pi_{k}^{*})=\pi_{L}^{*}-\bar{\pi}_{L}$, the definition of $\mathcal{R}_{I}^{(0)}$ in (C.22) and
$1/2\leq\pi_{\ell}/\pi_{\ell}^{*}\leq 3/2,\qquad\forall~\ell\in[L],$ (C.28)
which is implied by $d(\pi,\pi^{*})\leq 1/2$ from (C.2). By (C.26), the bound in (C.3.1) also holds for $\ell=L$ as
$|\nabla^{\pi}\bm{\pi}_{L}(\bar{\omega})|$ $~\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left|\sum_{k=1}^{L-1}\left(\sum_{a=1}^{L-1}\pi_{a}\mathbb{E}\left[I_{aL}(X,\omega)\right]-\pi_{L}\mathbb{E}\left[I_{kL}(X,\omega)\right]\right)\left(\bar{\pi}_{k}-\pi_{k}^{*}\right)\right|$
$~\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left|{\pi_{L}^{*}-\bar{\pi}_{L}\over\pi_{L}^{*}}\pi_{L}^{*}\sum_{a=1}^{L-1}\pi_{a}\mathbb{E}\left[I_{aL}(X,\omega)\right]-\pi_{L}\sum_{k=1}^{L-1}\pi_{k}^{*}\mathbb{E}\left[I_{kL}(X,\omega)\right]{\bar{\pi}_{k}-\pi_{k}^{*}\over\pi_{k}^{*}}\right|$
$~\lesssim~\pi_{L}^{*}~\mathcal{R}_{I}^{(0)}~d(\bar{\pi},\pi^{*}).$
Controlling the partial derivative w.r.t. $\mu_{k}$’s
We bound from above
$|\nabla^{M}\bm{\pi}_{\ell}(\bar{\omega})|~\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left|\sum_{k=1}^{L}\mathbb{E}\left[\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial\mu_{k}}\bigg|_{\omega},~\bar{\mu}_{k}-\mu_{k}^{*}\right\rangle\right]\right|.$
For any $k\neq\ell$, note that
$\frac{\partial\gamma_{\ell}(x;\omega)}{\partial\mu_{k}}$ $=\frac{\pi_{\ell}\pi_{k}}{2}I_{k\ell}(x,\omega)J({\bm{e}}_{k}-{\bm{e}}_{\ell}),$ (C.29)
$\frac{\partial\gamma_{\ell}(x;\omega)}{\partial\mu_{\ell}}$ $=\sum_{a\in[L]\setminus\{\ell\}}\frac{\pi_{\ell}\pi_{a}}{2}I_{a\ell}(x,\omega)J({\bm{e}}_{a}-{\bm{e}}_{\ell})=\sum_{a\in[L]\setminus\{\ell\}}\frac{\partial\gamma_{\ell}(x;\omega)}{\partial\mu_{a}}.$ (C.30)
Fix any $a\in[L]$ with $a\neq\ell$. By using (C.29), we have
$\left|\mathbb{E}\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial\mu_{a}},\bar{\mu}_{a}-\mu_{a}^{*}\right\rangle\right|$ $={\pi_{\ell}\pi_{a}\over 2}\mathbb{E}\left[I_{a\ell}(X,\omega)\right]\left|(\bar{\mu}_{a}-\mu_{a}^{*})^{\top}J({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|$
$\leq{\pi_{\ell}\pi_{a}\over 2}\mathbb{E}\left[I_{a\ell}(X,\omega)\right]\|\Sigma^{*1/2}J({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}\|\bar{\mu}_{a}-\mu_{a}^{*}\|_{\Sigma^{*}},$
$\leq\pi_{\ell}\pi_{a}\sqrt{\Delta_{a\ell}}~\mathbb{E}\left[I_{a\ell}(X,\omega)\right]\|\bar{\mu}_{a}-\mu_{a}^{*}\|_{\Sigma^{*}}$ (C.31)
where in the last step we used
$\|\Sigma^{*1/2}J({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$
$\leq\|\Sigma^{*1/2}(J-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}+\|\Sigma^{*1/2}J^{*}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$
[…]
Similarly, we also have
$\left|\mathbb{E}\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial\mu_{\ell}},\bar{\mu}_{\ell}-\mu_{\ell}^{*}\right\rangle\right|$ $\leq\pi_{\ell}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}[I_{a\ell}(X,\omega)]\sqrt{\Delta_{a\ell}}~\|\bar{\mu}_{\ell}-\mu_{\ell}^{*}\|_{\Sigma^{*}}$
$\leq\pi_{\ell}\sqrt{\Delta_{\max}}~\mathcal{R}_{I}^{(0)}\|\bar{\mu}_{\ell}-\mu_{\ell}^{*})\|_{\Sigma^{*}}.$ (C.33)
Combining (C.31) and (C.33) together with (C.28) gives that for all $\ell\in[L]$,
$|\nabla^{M}\bm{\pi}_{\ell}(\bar{\omega})|~\lesssim~\pi_{\ell}^{*}\sqrt{\Delta_{\max}}~\mathcal{R}_{I}^{(0)}d(\bar{M},M^{*}).$ (C.34)
Controlling the partial derivative w.r.t. $J$
We proceed to bound
$|\nabla^{J}\bm{\pi}_{\ell}(\bar{\omega})|~\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left|\mathbb{E}\left[\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial J}\bigg|_{\omega},~\bar{J}-J^{*}\right\rangle\right]\right|.$
Note that the partial derivatives of $\gamma_{\ell}(x,\omega)$ with respect to $J$ is
$\frac{\partial\gamma_{\ell}(x;\omega)}{\partial J}=-\sum_{a\in[L]\setminus\{\ell\}}\pi_{\ell}\pi_{a}I_{a\ell}(x,\omega)\left(x-\frac{\mu_{\ell}+\mu_{a}}{2}\right)({\bm{e}}_{a}-{\bm{e}}_{\ell})^{\top}\in\mathbb{R}^{d\times L}$ (C.35)
with $I_{a\ell}(x,\omega)$ given in (C.23). Since
$\left\langle\left(X-\frac{\mu_{a}+\mu_{\ell}}{2}\right)({\bm{e}}_{a}-{\bm{e}}_{\ell})^{\top},\bar{J}-J^{*}\right\rangle=$~ $\left(X-\frac{\mu_{a}+\mu_{\ell}}{2}\right)^{\top}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell}),$
the main task is to analyze
$\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)\left(X-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right]$
$=~\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\sum_{b=1}^{L}\pi_{b}^{*}\mathbb{E}\left[I_{a\ell}(X,\omega)\mid Y=b\right]\left(\mu_{b}^{*}-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})$ (C.36)
$\qquad+\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)~N^{\top}\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right]$ (C.37)
where we use the fact that
$X~\stackrel{{\scriptstyle d}}{{=}}~\Sigma^{*1/2}N+\sum_{b=1}^{L}1\{Y=b\}\mu_{b}^{*}$ (C.38)
with $N\sim\mathcal{N}_{d}(0,\mathbf{I}_{d})$. Since condition (C.2) ensures that
$\|\mu_{\ell}^{*}-(\mu_{\ell}+\mu_{a})/2\|_{\Sigma^{*}}\leq{\sqrt{\Delta_{a\ell}}\over 2}+{1\over 2}\|\mu^{*}_{\ell}-\mu_{\ell}\|_{\Sigma^{*}}+{1\over 2}\|\mu^{*}_{a}-\mu_{a}\|_{\Sigma^{*}}\leq\sqrt{\Delta_{a\ell}},$ (C.39)
by the definition in (C.22), we find
(C.36) $\leq\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}[I_{a\ell}(X,\omega)]\max_{b\in[L]}\left(\sqrt{\Delta_{b\ell}}+\sqrt{\Delta_{a\ell}}\right)\left\|\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}$
$\leq 4\sqrt{\Delta_{\max}}~\mathcal{R}_{I}^{(0)}~d(\bar{J},J^{*}).$
To bound (C.37), applying Cauchy-Schwarz inequality twice and using $\|\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}\leq 2d(\bar{J},J^{*})$ yield
$\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)~N^{\top}\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right]$
$\leq 2d(\bar{J},J^{*})\sqrt{\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}[I_{a\ell}(X,\omega)]}\sqrt{\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\|\mathbb{E}[I_{a\ell}(X,\omega)NN^{\top}\|_{\rm op}}$
$\leq 2d(\bar{J},J^{*})\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}.$ (C.40)
The last step use the definitions of $\mathcal{R}_{I}^{(0)}$ and $\mathcal{R}_{I}^{(2)}$ in (C.22). Combining the bounds of (C.36) and (C.37) and using (C.28) give
$|\nabla^{J}\bm{\pi}_{\ell}(\bar{\omega})|~\lesssim\pi_{\ell}^{*}\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}+\sqrt{\Delta_{\max}}~\mathcal{R}_{I}^{(0)}\right)~d(\bar{J},J^{*}).$ (C.41)
Finally, by collecting the bounds in Sections C.3.1, C.41 and C.34, we conclude that
${|\bm{\pi}_{\ell}(\bar{\theta})-\pi_{\ell}^{*}|\over\pi_{\ell}^{*}}$ $\leq{1\over\pi_{\ell}^{*}}\left|\mathbb{E}[\gamma_{\ell}(X;\bar{\omega})]-\mathbb{E}[\gamma_{\ell}(X;\omega^{*})]\right|$
$\lesssim\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}d(\bar{J},J^{*})+\mathcal{R}_{I}^{(0)}d(\bar{\omega},\omega^{*})$
[…]
Tis completes the proof of (C.17). ∎
C.3.2 Proof of Eq. C.18 of Theorem 19: Contraction of the population level M-step of the mean vectors
Proof.
For any $\omega=(\pi,M,J)$ and $\ell\in[L]$, note that
$\bm{\mu}_{\ell}(\omega)-\mu_{\ell}^{*}=\frac{\mathbb{E}[\gamma_{\ell}(X;\omega)(X-\mu_{\ell}^{*})]}{\mathbb{E}[\gamma_{\ell}(X;\omega)]}={1\over\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[(\gamma_{\ell}(X;\omega)-\gamma_{\ell}(X;\omega^{*}))(X-\mu_{\ell}^{*})\right].$
Fix any $\bar{\omega}\in\mathcal{B}_{d}(\omega^{*})$ and any $\ell\in[L]$. It then follows that
$\bm{\mu}_{\ell}(\bar{\omega})-\bm{\mu}_{\ell}(\omega^{*})$ $=\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})+\nabla^{M}\bm{\mu}_{\ell}(\bar{\omega})+\nabla^{J}\bm{\mu}_{\ell}(\bar{\omega})$
where
$\begin{split}\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})&:=\sum_{k=1}^{L-1}\int_{0}^{1}\frac{1}{\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[\left(X-\mu_{\ell}^{*}\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial\pi_{k}}\Big|_{\omega=(\pi_{u},\bar{M},\bar{J})},~\left(\bar{\pi}_{k}-\pi_{k}^{*}\right)\right\rangle\right]{\rm d}u\\ \nabla^{M}\bm{\mu}_{\ell}(\bar{\omega})&:=\sum_{k=1}^{L}\int_{0}^{1}\frac{1}{\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[\left(X-\mu_{\ell}^{*}\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial\mu_{k}}\Big|_{\omega=(\pi^{*},M_{u},\bar{J})},~\bar{\mu}_{k}-\mu_{k}^{*}\right\rangle\right]{\rm d}u\\ \nabla^{J}\bm{\mu}_{\ell}(\bar{\omega})&:=\int_{0}^{1}\frac{1}{\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[\left(X-\mu_{\ell}^{*}\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial J}\Big|_{\omega=(\pi^{*},M^{*},J_{u})},~\bar{J}-J^{*}\right\rangle\right]{\rm d}u.\end{split}$ (C.42)
The main task below is to bound each terms in terms of $\mathcal{R}_{I}^{(j)}$ with $j\in\{0,1,2\}$. For future reference, we notice that
$1/2\leq\bm{\pi}_{\ell}(\omega)/\pi_{\ell}^{*}\leq 3/2,\qquad\forall~\omega\in\mathcal{B}_{d}(\omega^{*}),$ (C.43)
since
[…]
Controlling the partial derivative of $\bm{\mu}_{\ell}(\omega)$ w.r.t. $\pi_{k}$’s
We need to bound from above
$\left\|\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\frac{1}{\bm{\pi}_{\ell}(\omega)}\left\|\sum_{k=1}^{L-1}(\bar{\pi}_{k}-\pi_{k}^{*})~\mathbb{E}\left[\left(X-\mu_{\ell}^{*}\right){\partial\gamma_{\ell}(X;\omega)\over\partial\pi_{k}}\right]\right\|_{\Sigma^{*}}.$
From the expressions of $\partial\gamma_{\ell}(x;\omega)$ in Eqs. C.24, C.25 and C.26 as well as the definition of $I_{a\ell}(x,\omega)$ in (C.23), by similar arguments of proving (C.3.1), we obtain that for any $\ell<L$,
$\left\|\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim d(\bar{\pi},\pi^{*})\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\frac{\pi_{\ell}+\pi_{\ell}^{*}}{\bm{\pi}_{\ell}(\omega)}\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)(X-\mu_{\ell}^{*})\right]\right\|_{\Sigma^{*}}$
$\lesssim d(\bar{\pi},\pi^{*})\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)(X-\mu_{\ell}^{*})\right]\right\|_{\Sigma^{*}}.$ (C.45)
The last step uses (C.28) and (C.43). For any $\omega\in\mathcal{B}_{d}(\omega^{*})$, using (C.38) and the definitions of $\mathcal{R}_{I}^{(0)}$ and $\mathcal{R}_{I}^{(1)}$ in (C.22) gives
$\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)(X-\mu_{\ell}^{*})\right]\right\|_{\Sigma^{*}}$
$\leq\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)N\right]\right\|_{2}+\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\sum_{b=1}^{L}\pi_{b}^{*}\mathbb{E}\left[I_{a\ell}(X,\omega)\left(\mu_{b}^{*}-\mu_{\ell}^{*}\right)\mid Y=b\right]\right\|_{\Sigma^{*}}$
$\leq\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}+\mathcal{R}_{I}^{(0)}\max_{b\in[L]}\left\|\mu_{b}^{*}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}$
$\leq\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}+\mathcal{R}_{I}^{(0)}\sqrt{\Delta_{\max}}.$ (C.46)
We thus obtain that for any $\ell<L$,
$\left\|\nabla^{\pi}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim~\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}+\mathcal{R}_{I}^{(0)}\sqrt{\Delta_{\max}}\right)~d(\bar{\pi},\pi^{*}).$ (C.47)
From (C.26) and by analogous argument, it is easy to see that (C.47) also holds for $\ell=L$.
Controlling the partial derivative of $\bm{\mu}_{\ell}(\omega)$ w.r.t. $M$
By the expression of $\partial\gamma_{\ell}(x;\omega)/\partial\mu_{k}$ in (C.29) – (C.30) and the definition of $I_{a\ell}(x,\omega)$ in (C.23), $\|\nabla^{M}\bm{\mu}_{\ell}(\bar{\omega})\|_{\Sigma^{*}}$ is bounded from above by
$\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left\|\sum_{k=1}^{L}{1\over\bm{\pi}_{\ell}(\omega)}\mathbb{E}\left[(X-\mu_{\ell}^{*})\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial\mu_{k}},\bar{\mu}_{k}-\mu_{k}^{*}\right\rangle\right]\right\|_{\Sigma^{*}}$
$\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}{2\pi_{\ell}\over\bm{\pi}_{\ell}(\omega)}\left\|\sum_{k\in[L]\setminus\{\ell\}}\pi_{k}\mathbb{E}\left[I_{k\ell}(X,\omega)(X-\mu_{\ell}^{*})\right]({\bm{e}}_{k}-{\bm{e}}_{\ell})^{\top}J^{\top}(\bar{\mu}_{k}-\mu_{k}^{*})\right\|_{\Sigma^{*}}$
$\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}6\left\|\sum_{k\in[L]\setminus\{\ell\}}\pi_{k}\mathbb{E}\left[I_{k\ell}(X,\omega)N\right]({\bm{e}}_{k}-{\bm{e}}_{\ell})^{\top}J^{\top}(\bar{\mu}_{k}-\mu_{k}^{*})\right\|_{2}$
$~+\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}6\left\|\sum_{k\in[L]\setminus\{\ell\}}\pi_{k}\mathbb{E}\left[I_{k\ell}(X,\omega)\sum_{b=1}^{L}1\{Y=b\}\right](\mu_{b}^{*}-\mu_{\ell}^{*})({\bm{e}}_{k}-{\bm{e}}_{\ell})^{\top}J^{\top}(\bar{\mu}_{k}-\mu_{k}^{*})\right\|_{\Sigma^{*}}.$
The second step uses (C.28) and (C.43). By Cauchy-Schwarz inequality, the first term is no greater than
$\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}6\Bigl\|\sum_{k\in[L]\setminus\{\ell\}}\pi_{k}\mathbb{E}\left[I_{k\ell}(X,\omega)NN^{\top}\right]\Bigr\|_{\rm op}^{1/2}$
$\qquad\qquad\Bigl(\sum_{k\in[L]\setminus\{\ell\}}\pi_{k}\mathbb{E}\left[I_{k\ell}(X,\omega)\right]\Bigr)^{1/2}\max_{k\in[L]}\left|({\bm{e}}_{k}-{\bm{e}}_{\ell})^{\top}J^{\top}(\bar{\mu}_{k}-\mu_{k}^{*})\right|$
$\lesssim\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}~\|\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})\|_{2}\|\bar{\mu}_{k}-\mu_{k}^{*}\|_{\Sigma^{*}}$
$\lesssim\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}~\sqrt{\Delta_{\max}}~d(\bar{M},M^{*})$ (C.48)
with the last step due to (C.32). Furthermore, the second term is bounded by
$6\mathcal{R}_{I}^{(0)}\max_{b\in[L]}\left\|\mu_{b}^{*}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}\max_{k\in[L]}\|\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})\|_{2}\|\bar{\mu}_{k}-\mu_{k}^{*}\|_{\Sigma^{*}}\lesssim\Delta_{\max}~\mathcal{R}_{I}^{(0)}~d(\bar{M},M^{*})$
using (C.32) and (C.43). We thus have
$\left\|\nabla^{M}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}~\lesssim~\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}+\mathcal{R}_{I}^{(0)}\sqrt{\Delta_{\max}}\right)\sqrt{\Delta_{\max}}~d(\bar{M},M^{*}).$ (C.49)
Controlling the partial derivative of $\bm{\mu}_{\ell}(\omega)$ w.r.t. $J$
We proceed to bound from above
$\left\|\nabla^{J}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $~\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\frac{1}{\bm{\pi}_{\ell}(\omega)}\left\|\mathbb{E}\left[\left(X-\mu_{\ell}^{*}\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial J}\Big|_{\omega},\bar{J}-J^{*}\right\rangle\right]\right\|_{\Sigma^{*}}.$
By (C.35) and (C.43), we need to bound the following term
$\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[(X-\mu_{\ell}^{*})\left\langle I_{a\ell}(X,\omega)\left(X-\frac{\mu_{\ell}+\mu_{a}}{2}\right)({\bm{e}}_{a}-{\bm{e}}_{\ell})^{\top},\bar{J}-J^{*}\right\rangle\right]\right\|_{\Sigma^{*}}$
$=\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)\Sigma^{*-1/2}(X-\mu_{\ell}^{*})\left(X-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right]\right\|_{2}.$
Recalling (C.38), we have the decomposition
$\Sigma^{*-1/2}(X-\mu_{\ell}^{*})\left(X-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}\Sigma^{*-1/2}\stackrel{{\scriptstyle d}}{{=}}~NN^{\top}+T_{1}+T_{2}+T_{3}$ (C.50)
where the matrices $T_{1},T_{2},T_{3}$ are given by
$T_{1}$ $=\Sigma^{*-1/2}\sum_{b=1}^{L}1\{Y=b\}(\mu_{b}^{*}-\mu_{\ell}^{*})N^{\top}$
$T_{2}$ $=N\sum_{b=1}^{L}1\{Y=b\}\left(\mu_{b}^{*}-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}\Sigma^{*-1/2}$
$T_{3}$ $=\Sigma^{*-1/2}\sum_{b=1}^{L}1\{Y=b\}\left(\mu_{b}^{*}-\mu_{\ell}^{*}\right)\sum_{b=1}^{L}1\{Y=b\}\left(\mu_{b}^{*}-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}\Sigma^{*-1/2}.$
By using Section C.3.1, we find
$\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)T_{1}\right]\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}$
$\leq\left|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)N^{\top}\right]\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|\max_{b\in[L]}\left\|\mu_{b}^{*}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}$
$\leq 2\sqrt{\Delta_{\max}}\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}~d(\bar{J},J^{*}).$
Similarly, by (C.39) and the argument in (C.3.2), we also find
$\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)T_{2}\right]\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}$
$\leq\max_{b\in[L]}\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)N\right]\left(\mu_{b}^{*}-\frac{\mu_{\ell}+\mu_{a}}{2}\right)^{\top}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}$
$\leq\Bigl\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)NN^{\top}\right]\Bigr\|_{\rm op}^{1/2}$
$\qquad\Bigl(\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)\right]\Bigr)^{1/2}\max_{b\in[L]}\left\|\mu_{b}^{*}-\frac{\mu_{\ell}+\mu_{a}}{2}\right\|_{\Sigma^{*}}\|\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$
$\leq 2\sqrt{\Delta_{\max}}\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(1)}}~d(\bar{J},J^{*}).$
as well as
$\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)T_{3}\right]\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}$
$=\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(X,\omega)\right]\max_{b\in[L]}\left\|\mu_{b}^{*}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}\left\|\mu_{b}^{*}-\frac{\mu_{\ell}+\mu_{a}}{2}\right\|_{\Sigma^{*}}2d(\bar{J},J)$
$\leq 4\Delta_{\max}\mathcal{R}_{I}^{(0)}d(\bar{J},J).$
Finally, since one can very by using the Cauchy-Schwarz inequality that
$\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}[I_{a\ell}(X,\omega)NN^{\top}]\Sigma^{*1/2}(\bar{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}\leq 2\sqrt{\mathcal{R}_{I}^{(1)}\mathcal{R}_{I}^{(2)}}d(\bar{J},J^{*}),$
combining with the previous three bounds gives
$\left\|\nabla^{J}\bm{\mu}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim\left[\sqrt{\mathcal{R}_{I}^{(1)}\mathcal{R}_{I}^{(2)}}+\sqrt{\Delta_{\max}}\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}+\sqrt{\Delta_{\max}}\mathcal{R}_{I}^{(0)}\right)\right]d(\bar{J},J^{*}).$ (C.51)
In view of Eqs. C.47, C.49 and C.51 and using $\mathcal{R}_{I}^{(1)}\leq\mathcal{R}_{I}^{(2)}$, we conclude that
$\begin{split}\left\|\bm{\mu}_{\ell}(\bar{\omega})-\bm{\mu}_{\ell}(\omega^{*})\right\|_{\Sigma^{*}}&\lesssim\left(\sqrt{\mathcal{R}_{I}^{(0)}\mathcal{R}_{I}^{(2)}}+\mathcal{R}_{I}^{(0)}\sqrt{\Delta_{\max}}\right)~d(\bar{\omega},\omega^{*})+\sqrt{\mathcal{R}_{I}^{(1)}\mathcal{R}_{I}^{(2)}}d(\bar{J},J^{*})\\ &\lesssim\kappa_{w}\sqrt{\Delta_{\max}}~d(\bar{\omega},\omega^{*})\end{split}$ (C.52)
where the last step invokes Lemma 22. This completes the proof of (C.18). ∎
C.4 Key technical lemmas used in the proof of Theorem 19
For any $k,\ell\in[L]$ and $x\in\mathbb{R}^{L}$, define
$W_{k\ell}(x,\omega)$ $:=\left(x-\frac{\mu_{\ell}+\mu_{k}}{2}\right)^{\top}J({\bm{e}}_{k}-{\bm{e}}_{\ell}).$ (C.53)
so that for any $j\in[L]$,
$\left(W_{k\ell}(X,\omega)\mid Y=j\right)~\stackrel{{\scriptstyle d}}{{=}}~\delta^{j}_{k\ell}(\omega)+N^{\top}\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})$ (C.54)
where $N\sim\mathcal{N}_{d}(0,\mathbf{I}_{d})$ and
$\delta_{k\ell}^{j}(\omega)=\bigl(\mu_{j}^{*}-\frac{\mu_{\ell}+\mu_{k}}{2}\bigr)^{\top}J({\bm{e}}_{k}-{\bm{e}}_{\ell}).$ (C.55)
For future reference, we note that
$\|\Sigma^{*1/2}J^{*}({\bm{e}}_{k}-{\bm{e}}_{\ell})\|_{2}^{2}=\Delta_{k\ell}=\delta_{k\ell}^{k}(\omega^{*})=-\delta_{k\ell}^{\ell}(\omega^{*}).$
Consider the event
$\mathcal{E}_{N}=\left\{|N^{\top}\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})|<\frac{c_{2}}{2c_{1}}\sqrt{\Delta_{k\ell}}~\|\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})\|_{2}\right\}$ (C.56)
with $c_{2}=(1-4c_{0}-4c_{0}^{2})/2$ and $c_{1}=1+2c_{0}$. We know that
$\mathbb{P}(\mathcal{E}_{N}^{c})\leq 2\exp\left(-\frac{c_{2}^{2}}{8c_{1}^{2}}\Delta_{k\ell}\right).$ (C.57)
Lemma 20 further ensures that for any fixed $\omega\in\mathcal{B}_{d}(\omega^{*})$ and any $k,\ell\in[L]$, on the event $\mathcal{E}_{N}$,
$(W_{k\ell}(X,\omega)\mid Y=\ell)$ $\leq-c_{2}\Delta_{k\ell}+{c_{2}\over 2}\sqrt{\Delta_{k\ell}}\sqrt{\Delta_{k\ell}}\leq-{c_{2}\over 2}\Delta_{k\ell},$ (C.58)
$(W_{k\ell}(X,\omega)\mid Y=k)$ $\geq~c_{2}\Delta_{k\ell}-{c_{2}\over 2}\sqrt{\Delta_{k\ell}}\sqrt{\Delta_{k\ell}}\geq~{c_{2}\over 2}\Delta_{k\ell}.$ (C.59)
Lemma 20.
For any $\omega=(\pi,M,J)\in\mathcal{B}_{d}(\omega^{*},c_{0})$ given in (C.2) and any $k,\ell\in[L]$, one has
$\|\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})\|_{2}\leq(1+2c_{0})\sqrt{\Delta_{k\ell}}$
and
$\delta_{k\ell}^{k}(\omega)\geq{1\over 2}\left(1-4c_{0}-4c_{0}^{2}\right)\Delta_{k\ell},\qquad\delta_{k\ell}^{\ell}(\omega)\leq-{1\over 2}\left(1-4c_{0}-4c_{0}^{2}\right)\Delta_{k\ell}.$
Proof.
The first result follows from
$\|\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})\|_{2}$ $\leq\|\Sigma^{*1/2}J^{*}({\bm{e}}_{k}-{\bm{e}}_{\ell})\|_{2}+\|\Sigma^{*1/2}(J-J^{*})({\bm{e}}_{k}-{\bm{e}}_{\ell})\|_{2}\hskip-56.9055pt$
[…]
For the other result, notice that
$\delta_{k\ell}^{\ell}(\omega)$ $=({\bm{e}}_{k}-{\bm{e}}_{\ell})^{\top}J^{\top}\left(\mu_{\ell}^{*}-\frac{\mu_{\ell}+\mu_{k}}{2})\right)$
$=({\bm{e}}_{k}-{\bm{e}}_{\ell})^{\top}J^{*\top}\left(\mu_{\ell}^{*}-\frac{\mu_{\ell}^{*}+\mu_{k}^{*}}{2}\right)+({\bm{e}}_{k}-{\bm{e}}_{\ell})^{\top}(J-J^{*})^{\top}\left(\mu_{\ell}^{*}-\frac{\mu_{\ell}+\mu_{k}}{2}\right)$
$\qquad+{1\over 2}({\bm{e}}_{k}-{\bm{e}}_{\ell})^{\top}J^{*\top}(\mu_{\ell}^{*}-\mu_{\ell}+\mu_{k}^{*}-\mu_{k}).$
The first term equals to
[…]
and the second and third terms are no greater than
${1\over 2}\|\Sigma^{*1/2}(J-J^{*})({\bm{e}}_{k}-{\bm{e}}_{\ell})\|_{2}\left(\|\mu_{\ell}-\mu_{\ell}^{*}-\mu_{k}+\mu_{k}^{*}\|_{\Sigma^{*}}+\sqrt{\Delta_{k\ell}}\right)$
$\quad+{1\over 2}\|\mu_{\ell}-\mu_{\ell}^{*}-\mu_{k}+\mu_{k}^{*}\|_{\Sigma^{*}}\sqrt{\Delta_{k\ell}}.$
By (C.2), we conclude
$\left({1\over 2}-2c_{0}-2c_{0}^{2}\right)\Delta_{k\ell}\geq\delta_{k\ell}^{\ell}(\omega)\geq-\left({1\over 2}+2c_{0}+2c_{0}^{2}\right)\Delta_{k\ell}.$
By the symmetric argument, we can also prove the result for $\delta_{k\ell}^{k}(\omega)$. The proof is complete. ∎
The following lemma controls the (conditional) expectation of various quantities related with $\gamma_{\ell}(X;\omega)$ and the Gaussian noise $N$. Note that due to independence between $N$ and $Y$, these conditional expectations are deterministic.
Lemma 21.
For any $\omega\in\mathcal{B}_{d}(\omega^{*},c_{0})$ given by (C.2), there exists some absolute positive constants $c=c(c_{0})$ and $C>0$ such that for any $b\neq\ell\in[L]$ and any $a\geq 1$,
$\mathbb{E}[\gamma_{\ell}(X;\omega)^{a}\mid Y=b]$ $~\leq~C\left({\pi_{\ell}^{*}\over\pi_{b}^{*}}\exp(-c\Delta_{b\ell})\right)^{a}+\exp(-c\Delta_{b\ell}),$ (C.60)
$\left\|\mathbb{E}\left[\gamma_{\ell}(X;\omega)N\mid Y=b\right]\right\|_{2}$ $~\leq~C{\pi_{\ell}^{*}+\pi_{b}^{*}\over\pi_{b}^{*}}\exp(-c\Delta_{b\ell}),$ (C.61)
$\left\|\mathbb{E}\left[\gamma_{\ell}(X;\omega)NN^{\top}\mid Y=b\right]\right\|_{{\rm op}}$ $~\leq~C{\pi_{\ell}^{*}+\pi_{b}^{*}\over\pi_{b}^{*}}\exp(-c\Delta_{b\ell}),$ (C.62)
$\mathbb{E}\left[\gamma_{\ell}(X;\omega)\|N\|_{2}^{2}\mid Y=b\right]$ $~\leq~C{\pi_{\ell}^{*}+\pi_{b}^{*}\over\pi_{b}^{*}}d\exp(-c\Delta_{b\ell}),$ (C.63)
$\left\|\mathbb{E}\left[\gamma_{\ell}(X;\omega)\|N\|_{2}^{2}NN^{\top}\mid Y=b\right]\right\|_{\rm op}$ $~\leq~C{\pi_{\ell}^{*}+\pi_{b}^{*}\over\pi_{b}^{*}}d\exp(-c\Delta_{b\ell}).$ (C.64)
In particular, the above holds for $\omega=\omega^{*}$ with $c_{0}=0$.
Proof.
Pick any $b\neq\ell$ and $\omega\in\mathcal{B}_{d}(\omega^{*})$. Recall $\mathcal{E}_{N}$ from (C.56). Start with
$\mathbb{E}[\gamma_{\ell}(X;\omega)^{a}\mid Y=b]$ $\leq\mathbb{E}\left[\gamma_{\ell}(X;\omega)^{a}1\{\mathcal{E}_{N}\}\mid Y=b\right]+\mathbb{P}(\mathcal{E}_{N}^{c}).$
Note that on $\mathcal{E}_{N}$ and conditioning on $Y=b$,
[…]
By using (C.28) and (C.57), we obtain
$\mathbb{E}[\gamma_{\ell}(X;\omega)^{a}\mid Y=b]$ $\leq 3^{a}\left(\frac{\pi^{*}_{\ell}}{\pi^{*}_{b}}\,\exp\Bigl(-{c_{2}\over 2}\Delta_{b\ell}\Bigr)\right)^{a}+\exp\Bigl(-\frac{c_{2}^{2}}{8c_{1}^{2}}\Delta_{b\ell}\Bigr).$
Setting $c=\min\{c_{2}/2,c_{2}^{2}/(16c_{1}^{2})\}$ gives (C.60).
Eq. C.61 follows by Cauchy-Schwarz inequality and (C.60) with $a=2$ as
$\left\|\mathbb{E}\left[\gamma_{\ell}(X;\omega)N\mid Y=b\right]\right\|_{2}$ $=\sup_{v\in\mathbb{S}^{d}}\mathbb{E}\left[\gamma_{\ell}(X;\omega^{*})v^{\top}N\mid Y=b\right]$
$\leq\sqrt{\mathbb{E}\left[\gamma_{\ell}(X;\omega^{*})^{2}\mid Y=b\right]}\sup_{v\in\mathbb{S}^{d}}\sqrt{\mathbb{E}[(v^{\top}N)^{2}]}.$
We use similar arguments to prove the other results. For (C.62), we find that
$\left\|\mathbb{E}\left[\gamma_{\ell}(X;\omega)NN^{\top}\mid Y=b\right]\right\|_{{\rm op}}$ $\leq\left\|\mathbb{E}\left[\gamma_{\ell}(X;\omega)NN^{\top}1\{\mathcal{E}_{N}\}\mid Y=b\right]\right\|_{{\rm op}}$
$\quad+\left\|\mathbb{E}\left[NN^{\top}1\{\mathcal{E}_{N}^{c}\}\right]\right\|_{{\rm op}}.$
The second term is bounded by
$\sup_{v\in\mathbb{S}^{d}}\mathbb{E}\left[v^{\top}NN^{\top}v1\{\mathcal{E}_{N}^{c}\}\right]$ $\leq\sup_{v\in\mathbb{S}^{d}}\sqrt{\mathbb{E}\left[(N^{\top}v)^{4}\right]}\sqrt{\mathbb{P}(\mathcal{E}_{N}^{c}\}}\leq 2\exp\Bigl(-\frac{c_{2}^{2}}{16c_{1}^{2}}\Delta_{b\ell}\Bigr).$ (C.66)
Since the first term can be bounded by using (C.65) with $a=1$:
$\left\|\mathbb{E}\left[\gamma_{\ell}(X;\omega)NN^{\top}1\{\mathcal{E}_{N}\}\mid Y=b\right]\right\|_{{\rm op}}$ $\leq{\pi_{\ell}\over\pi_{b}}\exp\left(-{c_{2}\over 2}\Delta_{b\ell}\right)\sup_{v\in\mathbb{S}^{d}}\mathbb{E}\left[(N^{\top}v)^{2}\right]$
$\leq 3{\pi_{\ell}^{*}\over\pi_{b}^{*}}\exp\left(-{c_{2}\over 2}\Delta_{b\ell}\right),$
we obtain (C.62). Eq. C.63 follows by noting that
$\mathbb{E}\left[\gamma_{\ell}(X;\omega)\|N\|_{2}^{2}\mid Y=b\right]$ $\leq\sqrt{\mathbb{E}[\gamma_{\ell}(X;\omega^{*})^{2}\mid Y=b]}\sqrt{\mathbb{E}[\|N\|_{2}^{4}]}\leq 6{\pi_{\ell}^{*}+\pi_{b}^{*}\over\pi_{b}^{*}}d\exp(-c\Delta_{b\ell}).$
Finally, by (C.65), we have
$\left\|\mathbb{E}\left[\gamma_{\ell}(X;\omega)\|N\|_{2}^{2}NN^{\top}\mid Y=b\right]\right\|_{\rm op}$
$\leq\sup_{v\in\mathbb{S}^{d}}{\pi_{\ell}\over\pi_{b}}\exp(-c\Delta_{b\ell})\mathbb{E}\left[\|N\|_{2}^{2}(v^{\top}N)^{2}\right]+\sup_{v\in\mathbb{S}^{d}}\sqrt{\mathbb{E}[(N^{\top}u)^{2}\|N\|_{2}^{2}]}\sqrt{\mathbb{P}(\mathcal{E}_{N}^{c})}$
$\lesssim{\pi_{\ell}^{*}+\pi_{b}^{*}\over\pi_{b}^{*}}d\exp(-c\Delta_{b\ell}),$
proving (C.64). This completes the proof. ∎
The following lemma establishes upper bounds of $\mathcal{R}_{I}^{(j)}$, for $j\in\{0,1,2\}$, given by (C.22).
Lemma 22.
For some constant $c>0$ depending only on $c_{1}$ and $c_{2}$ given in (C.56), and some absolute constant $C>0$, one has
$\max_{j\in\{0,1\}}\mathcal{R}_{I}^{(j)}~\leq~{C\over\pi_{\min}^{*}}\exp\left(-c\Delta_{\min}\right),\qquad\mathcal{R}_{I}^{(2)}~\leq~{CL\over\pi_{\min}^{*}}\exp\left(-c\Delta_{\min}\right).$ (C.67)
Proof.
We first prove $j=0$. For any $\ell\in[L]$ and $\omega$, define
$I_{\ell}(X,\omega):=\pi_{\ell}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}I_{a\ell}(X,\omega)=\gamma_{\ell}(X;\omega)(1-\gamma_{\ell}(X;\omega)).$
so that $\mathcal{R}_{I}^{(0)}=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}\pi_{\ell}^{-1}\mathbb{E}[I_{\ell}(X,\omega)]$. Note that $\mathbb{E}[I_{\ell}(X,\omega)]$ equals to
$\sum_{b=1}^{L}\pi_{b}^{*}\mathbb{E}\left[I_{\ell}(X,\omega)\mid Y=b\right]$
$\leq\pi_{\ell}^{*}~\mathbb{E}\left[(1-\gamma_{\ell}(X;\omega))\mid Y=\ell\right]+\sum_{b\in[L]\setminus\{\ell\}}\pi_{b}^{*}~\mathbb{E}\left[\gamma_{\ell}(X;\omega)\mid Y=b\right]$
$=\pi_{\ell}^{*}\sum_{a\in[L]\setminus\{\ell\}}\mathbb{E}\left[\gamma_{a}(X;\omega))\mid Y=\ell\right]+\sum_{b\in[L]\setminus\{\ell\}}\pi_{b}^{*}~\mathbb{E}\left[\gamma_{\ell}(X;\omega)\mid Y=b\right]$
$\lesssim\pi_{\ell}^{*}\sum_{a\in[L]\setminus\{\ell\}}{\pi_{a}^{*}+\pi_{\ell}^{*}\over\pi_{\ell}^{*}}\exp(-c\Delta_{\min})+\sum_{b\in[L]\setminus\{\ell\}}\pi_{b}^{*}{\pi_{\ell}^{*}+\pi_{b}^{*}\over\pi_{b}^{*}}\exp(-c\Delta_{\min})$ by (C.60)
$\lesssim(1+L\pi_{\ell}^{*})\exp(-c\Delta_{\min}).$
This together with (C.28) gives
$\mathcal{R}_{I}^{(0)}\lesssim\max_{\ell\in[L]}\left({1\over\pi_{\ell}^{*}}+L\right)\exp\left(-c\Delta_{\min}\right)\lesssim{1\over\pi_{\min}^{*}}\exp\left(-c\Delta_{\min}\right)$
proving the result for $j=0$.
To prove the result of $j=1$, we similarly bound $\|\mathbb{E}[I_{\ell}(X,\omega)NN^{\top}]$ by
$\pi_{\ell}^{*}\sum_{a\in[L]\setminus\{\ell\}}\left\|\mathbb{E}\left[\gamma_{a}(X;\omega))NN^{\top}\mid Y=\ell\right]\right\|_{\rm op}+\sum_{b\in[L]\setminus\{\ell\}}\pi_{b}^{*}\left\|\mathbb{E}\left[\gamma_{\ell}(X;\omega)NN^{\top}\mid Y=b\right]\right\|_{\rm op}.$
Invoking (C.62) and (C.28) gives the result of $j=1$.
Finally, for $j=2$, note that
$\mathcal{R}_{I}^{(2)}$ $\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}\sup_{\max_{a}\|v_{a}\|_{2}=1}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(N^{\top}v_{a})^{2}\right]$
$\leq\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}\sup_{\max_{b}\|v_{b}\|_{2}=1}\sum_{b\in[L]\setminus\{\ell\}}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}(N^{\top}v_{b})^{2}\right]$
$\leq(L-1)\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\left\|\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\mathbb{E}\left[I_{a\ell}NN^{\top}\right]\right\|_{\rm op}$
$=(L-1)\mathcal{R}_{I}^{(1)}.$
The proof is complete. ∎
Appendix D Proofs of the sample level results in Section 4.1
D.1 Proof of Theorem 5: concentration inequalities of the M-steps at the true parameter
[…]
Proof.
For any $\ell\in[L]$, we apply the Bernstein inequality to bound
$\widehat{\bm{\pi}}_{\ell}(\omega^{*})-\bm{\pi}_{\ell}(\omega^{*})$ $=\mathbb{E}_{n}[\gamma_{\ell}(X;\omega^{*})]-\mathbb{E}[\gamma_{\ell}(X;\omega^{*})]={1\over n}\sum_{i=1}^{n}\left\{\gamma_{\ell}(X_{i};\omega^{*})-\mathbb{E}[\gamma_{\ell}(X_{i};\omega^{*})]\right\}.$
Since the variance of each $\gamma_{\ell}(X_{i};\omega^{*})$ is no greater than
$\mathbb{E}[\gamma_{\ell}^{2}(X_{i};\omega^{*})]$ $\leq\pi_{\ell}^{*}+\sum_{b\in[L]\setminus\{\ell\}}\pi_{b}^{*}~\mathbb{E}[\gamma_{\ell}(X_{i};\omega^{*})^{2}\mid Y=b]$
$\leq\pi_{\ell}^{*}+(L\pi_{\ell}^{*}+1)\exp{(-c\Delta_{\min})}$ by Lemma 21
[…]
applying the Bernstein’s inequality gives that for all $t>0$,
$\mathbb{P}\left\{|\widehat{\bm{\pi}}_{\ell}(\omega^{*})-\bm{\pi}_{\ell}(\omega^{*})|\lesssim\sqrt{\pi_{\ell}^{*}t\over n}+{t\over n}\right\}\geq 1-e^{-t}.$
Taking the union bounds over $\ell\in[L]$, choosing $t=\log(n)$ and using $n\pi_{\min}^{*}\geq C\log n$ yield (4.2). ∎
[…]
Proof.
We work on the event
$\mathcal{E}_{\pi}:=\bigcap_{\ell\in[L]}\left\{{|\widehat{\bm{\pi}}_{\ell}(\omega^{*})-\pi_{\ell}^{*}|\over\pi_{\ell}^{*}}+{|\widetilde{\pi}_{\ell}-\pi_{\ell}^{*}|\over\pi_{\ell}^{*}}\leq C\sqrt{\log n\over n\pi_{\ell}^{*}}\right\}.$ (D.2)
which, by (4.2) and Lemma 29, holds with probability at least $1-n^{-1}$ under $n\pi_{\min}^{*}\geq C\log n$. On the event $\mathcal{E}_{\pi}$, we have
$1/2\leq\widehat{\bm{\pi}}_{\ell}(\omega^{*})/\pi_{\ell}^{*}\leq 3/2,\qquad c\leq\widetilde{\pi}_{\ell}/\pi_{\ell}^{*}\leq C,\qquad\forall~\ell\in[L].$ (D.3)
Pick any $\ell\in[L]$. Notice that
$\widehat{\bm{\mu}}_{\ell}(\omega^{*})-\mu_{\ell}^{*}={\mathbb{E}_{n}[\gamma_{\ell}(X;\omega^{*})(X-\mu_{\ell}^{*})]\over\mathbb{E}_{n}[\gamma_{\ell}(X;\omega^{*})]}$ $={1\over\widehat{\bm{\pi}}_{\ell}(\omega^{*})}{1\over n}\sum_{i=1}^{n}\underbrace{\gamma_{\ell}(X_{i};\omega^{*})(X_{i}-\mu_{\ell}^{*})}_{V_{\ell,i}}.$ (D.4)
By $\mathbb{E}[V_{\ell,i}]=\mathbb{E}[\gamma_{\ell}(X_{i};\omega^{*})](\bm{\mu}_{\ell}(\omega^{*})-\mu_{\ell}^{*})=0$, we have
$\|\widehat{\bm{\mu}}_{\ell}(\omega^{*})-\bm{\mu}_{\ell}(\omega^{*})\|_{\Sigma^{*}}={1\over n\widehat{\bm{\pi}}_{\ell}(\omega^{*})}\left\|\sum_{i=1}^{n}\left(V_{\ell,i}-\mathbb{E}[V_{\ell,i}]\right)\right\|_{\Sigma^{*}}.$
to which we use a truncation argument together with the vector-valued Bernstein inequality to analyze. From segol2021improved we know $\Sigma^{*-1/2}V_{\ell,i}$ is a sub-Gaussian random vector with some absolute sub-Gaussian constant $C_{v}>0$. Consider the event
$\mathcal{E}_{2}=\bigcap_{i=1}^{n}\mathcal{E}_{2,i},\qquad\mathcal{E}_{2,i}:=\left\{\|V_{\ell,i}\|_{\Sigma^{*}}\leq C_{v}(\sqrt{d}+2\sqrt{\log n})\right\}$ (D.5)
and by Lemma 30, one has $\mathbb{P}(\mathcal{E}_{2,i})\geq 1-n^{-2}$ and $\mathbb{P}(\mathcal{E}_{2})\geq 1-n^{-1}$. We first find that
$\mathbb{E}[\|V_{\ell,i}\|_{\Sigma^{*}}^{2}]$ $\leq\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})\|X_{i}-\mu_{\ell}^{*}\|_{\Sigma^{*}}^{2}\right]$
$\leq\pi_{\ell}^{*}\mathbb{E}\left[\|N_{i}\|_{2}^{2}\right]+2\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}^{*}\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})(\|N_{i}\|_{2}^{2}+\Delta_{a\ell})\mid Y_{i}=a\right]$
$\lesssim d\pi_{\ell}^{*}+\sum_{a\in[L]\setminus\{\ell\}}(\pi_{\ell}^{*}+\pi_{a}^{*})(d+\Delta_{a\ell})\exp(-c\Delta_{a\ell})$ by Lemma 21
$\lesssim d\pi_{\ell}^{*}+(L\pi_{\ell}^{*}+1)d\exp{(-c^{\prime}\Delta_{\min})}$ $\text{by }\mathcal{E}_{\pi}$
[…]
Then
${1\over n}\left\|\sum_{i=1}^{n}\mathbb{E}[V_{\ell,i}1\{\mathcal{E}_{2,i}^{c}\}]\right\|_{\Sigma^{*}}\leq~\max_{i\in[n]}\mathbb{E}\left[\|V_{\ell,i}\|_{\Sigma^{*}}^{2}\right]^{1/2}\mathbb{P}(\mathcal{E}_{2,i}^{c})^{1/2}\lesssim\sqrt{d\pi_{\ell}^{*}\over n^{2}}.$ (D.6)
On the event $\mathcal{E}_{2}$, it remains to bound from above
${1\over n}\left\|\sum_{i=1}^{n}\Bigl\{V_{\ell,i}1\{\mathcal{E}_{2,i}\}-\mathbb{E}\left[V_{\ell,i}1\{\mathcal{E}_{2,i}\}\right]\Bigr\}\right\|_{\Sigma^{*}}.$ (D.7)
Clearly, the following holds almost surely,
$\max_{i\in[n]}\|V_{\ell,i}1\{\mathcal{E}_{2,i}\}\|_{\Sigma^{*}}\leq C_{v}(\sqrt{d}+2\sqrt{\log n}).$
In conjunction with the above variance bound, applying the vector-valued Bernstein’s inequality in Lemma 33 gives that for all $t\geq 0$,
[…]
with probability at least $1-e^{-t}-2n^{-1}$. Together with (D.6), by choosing $t=\log n$ and using $n\pi_{\min}^{*}\geq C\log(n)(1\vee\log(n)/d)$, we obtain that for any $t\geq 1$, with probability $1-3n^{-1}$,
$\|\widehat{\bm{\mu}}_{\ell}(\omega^{*})-\bm{\mu}_{\ell}(\omega^{*})\|_{\Sigma^{*}}$ $\lesssim\sqrt{d\log n\over n\pi_{\ell}^{*}}+{\log n\over n\pi_{\ell}^{*}}\sqrt{d+\log n}\lesssim\sqrt{d\log n\over n\pi_{\ell}^{*}}.$ (D.8)
This completes the proof of (4.3). ∎
[…]
Proof.
We analyze the term
$d(\bm{\widehat{\Sigma}}(\omega^{*}),\bm{\Sigma}(\omega^{*}))=\left\|\Sigma^{*-1/2}(\bm{\widehat{\Sigma}}(\omega^{*})-\bm{\Sigma}(\omega^{*}))\Sigma^{*-1/2}\right\|_{{\rm op}}.$
It is easy to verify the decomposition
$\widehat{\Sigma}(\omega^{*})$ $=\mathbb{E}_{n}\left[\sum_{\ell=1}^{L}\gamma_{\ell}(X,\omega^{*})(X-\widehat{\bm{\mu}}_{\ell}(\omega^{*}))(X-\widehat{\bm{\mu}}_{\ell}(\omega^{*}))^{\top}\right]$
$={1\over n}\sum_{i=1}^{n}\sum_{\ell=1}^{L}\gamma_{\ell}(X_{i};\omega^{*})(X_{i}-\mu_{\ell}^{*})(X_{i}-\mu_{\ell}^{*})^{\top}-\sum_{\ell=1}^{L}\widehat{\bm{\pi}}_{\ell}(\omega^{*})(\mu_{\ell}^{*}-\widehat{\bm{\mu}}_{\ell}(\omega^{*}))(\mu_{\ell}^{*}-\widehat{\bm{\mu}}_{\ell}(\omega^{*}))^{\top}.$ (D.9)
For any $i\in[n]$, define
$H_{i}:=\sum_{\ell=1}^{L}\Sigma^{*-1/2}\gamma_{\ell}(X_{i};\omega^{*})(X_{i}-\mu_{\ell}^{*})(X_{i}-\mu_{\ell}^{*})^{\top}\Sigma^{*-1/2}$
and note that $\mathbb{E}[H_{i}]=\mathbf{I}_{d}.$ Adding and subtracting terms gives $d(\bm{\widehat{\Sigma}}(\omega^{*}),\bm{\Sigma}(\omega^{*}))\leq\textrm{I}+\textrm{II}$ where
$\textrm{I}:=$~ ${1\over n}\left\|\sum_{i=1}^{n}\left(H_{i}-\mathbb{E}[H_{i}]\right)\right\|_{\rm op}$
$\textrm{II}:=$~ $\left\|\sum_{\ell=1}^{L}\widehat{\bm{\pi}}_{\ell}(\omega^{*})\Sigma^{*-1/2}(\mu_{\ell}^{*}-\widehat{\bm{\mu}}_{\ell}(\omega^{*}))(\mu_{\ell}^{*}-\widehat{\bm{\mu}}_{\ell}(\omega^{*}))^{\top}\Sigma^{*-1/2}\right\|_{\rm op}.$ (D.10)
By (D.8) and $\mathcal{E}_{\pi}$ in (D.2), with probability at least $1-3n^{-1}$,
II $\lesssim\sum_{\ell=1}^{L}\pi_{\ell}^{*}\left\|\mu_{\ell}^{*}-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right\|_{\Sigma^{*}}^{2}\lesssim{dL\log n\over n}.$ (D.11)
To bound I, note that $\gamma_{\ell}(X_{i};\omega^{*})]^{1/2}(X_{i}-\mu_{\ell}^{*})$ is also sub-Gaussian with some constant $C>0$. In the following we work on the event
$\mathcal{E}_{2}=\bigcap_{i=1}^{n}\mathcal{E}_{2,i},\qquad\mathcal{E}_{2,i}:=\left\{\|[\gamma_{\ell}(X_{i};\omega^{*})]^{1/2}(X_{i}-\mu_{\ell}^{*})\|_{\Sigma^{*}}\leq C_{v}(\sqrt{d}+2\sqrt{\log n})\right\}$
We have $\textrm{I}\leq\textrm{I}_{1}+\textrm{I}_{2}$ with
$\textrm{I}_{1}$ $:={1\over n}\left\|\sum_{i=1}^{n}\left(H_{i}1\{\mathcal{E}_{2,i}\}-\mathbb{E}[H_{i}1\{\mathcal{E}_{2,i}\}]\right)\right\|_{\rm op},$
$\textrm{I}_{2}$ $:={1\over n}\left\|\sum_{i=1}^{n}\mathbb{E}\left[H_{i}1\{\mathcal{E}_{2,i}^{c}\}\right]\right\|_{\rm op}.$
We first bound $\textrm{I}_{1}$ by the matrix-valued Bernstein inequality of Lemma 34. For any $i\in[n]$, we have that almost surely,
$\|H_{i}1\{\mathcal{E}_{2,i}\}\|_{\rm op}$ $\leq\sum_{\ell=1}^{L}\gamma_{\ell}(X_{i};\omega^{*})\|X_{i}-\mu_{\ell}^{*}\|_{\Sigma^{*}}^{2}1\{\mathcal{E}_{2,i}\}\lesssim L(d+\log n).$
To bound the variance term in the Bernstein inequality, we decompose
$\left\|\mathbb{E}[H_{i}^{2}1\{\mathcal{E}_{2,i}\}]\right\|_{\rm op}$
$\leq\left\|\sum_{a=1}^{L}\pi_{a}^{*}\mathbb{E}\left[\gamma_{a}^{2}(X_{i};\omega^{*})\Sigma^{*-1/2}(X_{i}-\mu_{a}^{*})(X_{i}-\mu_{a}^{*})^{\top}\Sigma^{*-1/2}\|X_{i}-\mu_{a}^{*}\|_{\Sigma^{*}}^{2}\mid Y_{i}=a\right]\right\|_{\rm op}$
$\quad+\left\|\sum_{a=1}^{L}\pi_{a}^{*}\mathbb{E}\left[\left(\sum_{\ell\in[L]\setminus\{a\}}\gamma_{\ell}(X_{i};\omega^{*})\Sigma^{*-1/2}(X_{i}-\mu_{\ell}^{*})(X_{i}-\mu_{\ell}^{*})^{\top}\Sigma^{*-1/2}\right)^{2}\mid Y_{i}=a\right]\right\|_{\rm op}.$
The first term is bounded by
$\left\|\sum_{a=1}^{L}\pi_{a}^{*}\mathbb{E}\left[\gamma_{a}(X_{i};\omega^{*})N_{i}N_{i}^{\top}\|N_{i}\|_{2}^{2}\mid Y_{i}=a\right]\right\|_{\rm op}\leq\sup_{v\in\mathbb{S}^{d}}\mathbb{E}[(N_{i}^{\top}v)^{2}\|N_{i}\|_{2}^{2}]\leq 3d.$ (D.12)
while the second term is no greater than
$\sup_{v\in\mathbb{S}^{d}}\sum_{a=1}^{L}\pi_{a}^{*}\mathbb{E}\left[\sum_{\ell\in[L]\setminus\{a\}}\gamma_{\ell}(X_{i};\omega^{*})\langle v,X_{i}-\mu_{\ell}^{*}\rangle_{\Sigma^{*}}^{2}\sum_{\ell\in[L]\setminus\{a\}}\gamma_{\ell}(X_{i};\omega^{*})\left\|X_{i}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}^{2}\mid Y_{i}=a\right]$
$\leq 4\sup_{v\in\mathbb{S}^{d}}\sum_{a=1}^{L}\pi_{a}^{*}\sum_{\ell\in[L]\setminus\{a\}}\mathbb{E}\left[\gamma_{\ell}(X_{i};\omega^{*})\left((v^{\top}N_{i})^{2}+\Delta_{a\ell}\right)\left(\|N_{i}\|_{2}^{2}+\Delta_{a\ell}\right)\mid Y_{i}=a\right]$
$\lesssim\sum_{a=1}^{L}\sum_{\ell\in[L]\setminus\{a\}}(\pi_{\ell}^{*}+\pi_{a}^{*})(d+\Delta_{a\ell})\Delta_{a\ell}\exp(-c\Delta_{a\ell})$
$\lesssim dL\exp(-c\Delta_{\min})$ (D.13)
where the penultimate step uses Lemma 21. Combining (D.12) and (D.1.3) together with (3.4) yields that
$\left\|\sum_{i=1}^{n}\mathbb{E}[H_{i}^{2}1\{\mathcal{E}_{2,i}\}]\right\|_{\rm op}\lesssim nd+ndL\exp(-c\Delta_{\min})\lesssim nd.$
An application of the matrix-valued Bernstein inequality in Lemma 34 gives: for all $t\geq 0$,
$\textrm{I}_{1}\lesssim\sqrt{dt\over n}+{tL(d+\log n)\over n}$ (D.14)
with probability at least $1-de^{-t}-2n^{-1}$. Regarding $\textrm{I}_{2}$, by using the arguments in (D.12) and (D.1.3), one can deduce that
$\textrm{I}_{2}\leq\max_{i\in[n]}\left\|\mathbb{E}\left[H_{i}1\{\mathcal{E}_{2,i}^{c}\}\right]\right\|_{{\rm op}}\leq\max_{i\in[n]}\sqrt{\|\mathbb{E}[H_{i}\|H_{i}\|_{\rm op}]\|_{\rm op}}\sqrt{\mathbb{P}(\mathcal{E}_{2,i}^{c})}\lesssim{\sqrt{d}\over n}.$
Thus the bound in (D.14) holds for I for all $t\geq 1$. In conjunction with (D.11), by choosing $t=2\log n$ and using $d\leq n$ and $\log n\leq n\pi_{\min}^{*}\leq n/L$, we conclude that with probability at least 14/n1-4/n,
$d(\bm{\widehat{\Sigma}}(\omega^{*}),\bm{\Sigma}(\omega^{*}))$ $\lesssim\sqrt{d\log n\over n}+{(d+\log n)L\log n\over n}$
completing the proof of (4.4). ∎
D.2 Contraction of the sample level M-steps
For any $\theta$ satisfying (3.5), by recalling that $\omega=\omega(\theta)=(\pi,M,J(\theta))$, we establish in this subsection the sample level contraction. We introduce a few additional quantities:
[…]
which are the sample level counterparts of those in (C.22). In Lemma 24 we show that with probability at least $1-n^{-1}$,
$\widehat{\mathcal{R}}_{I}^{(0)}+\widehat{\mathcal{R}}_{I}^{(1)}=\mathcal{O}\left(\xi_{n}\right),\qquad\widehat{\mathcal{R}}_{I}^{(2)}=\mathcal{O}\left(L\xi_{n}\right).$
with
$\xi_{n}={1\over\pi_{\min}^{*2}}\exp(-c\Delta_{\min})+{dL\log(q)\log(n)\over n\pi_{\min}^{*}}.$ (D.16)
and $q:=n(\Delta_{\max}+1/\pi_{\min}^{*})$.
The following theorem gives the one-step contraction of the sample level M-steps. Recall $d(\omega,\omega^{*})$ from (C.9).
Theorem 23.
Grant condition (3.4). On the event
$\Delta_{\max}(\Delta_{\max}+L)\xi_{n}^{2}+d(\widehat{\bm{\pi}}(\omega^{*}),\bm{\pi}(\omega^{*}))\leq c,$ (D.17)
for any $\omega\in\mathcal{B}_{d}(\omega^{*})$, we have
$d(\widehat{\bm{\pi}}(\omega),\widehat{\bm{\pi}}(\omega^{*}))$ $~\leq~\xi_{n}(1+\sqrt{L/\Delta_{\max}})~d(\omega,\omega^{*}),$ (D.18)
$d(\widehat{\mathbf{M}}(\omega),\widehat{\mathbf{M}}(\omega^{*}))$ $~\leq~\xi_{n}\sqrt{L+\Delta_{\max}}~d(\omega,\omega^{*}).$ (D.19)
Furthermore, if the event
$d(\widehat{\mathbf{M}}(\omega^{*}),\mathbf{M}(\omega^{*}))\leq\sqrt{\Delta_{\max}}$ (D.20)
holds, then
$d(\bm{\widehat{\Sigma}}(\omega),\bm{\widehat{\Sigma}}(\omega^{*}))$ $~\leq~\xi_{n}\sqrt{(L+\Delta_{\max})\Delta_{\max}}~d(\omega,\omega^{*}).$ (D.21)
Proof.
We work on the events in (D.17) and (D.20).
Contraction of $\widehat{\pi}_{\ell}$.
Fix any $\ell\in[L]$ and $\bar{\omega}$. We have
$\widehat{\bm{\pi}}_{\ell}(\bar{\omega})-\widehat{\bm{\pi}}_{\ell}(\omega)$ $=\mathbb{E}_{n}[\gamma_{\ell}(X,\bar{\omega})-\gamma_{\ell}(X,\omega^{*})]=\nabla^{\pi}\widehat{\bm{\pi}}_{\ell}(\bar{\omega})+\nabla^{M}\widehat{\bm{\pi}}_{\ell}(\bar{\omega})+\nabla^{J}\widehat{\bm{\pi}}_{\ell}(\bar{\omega})$
where, similar as (C.21),
$\nabla^{\pi}\widehat{\bm{\pi}}_{\ell}(\bar{\omega})$ $:=\sum_{k=1}^{L-1}\int_{0}^{1}\mathbb{E}_{n}\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial\pi_{k}}\Big|_{\omega=(\pi_{u},\bar{M},\bar{J})},\bar{\pi}_{k}-\pi_{k}^{*}\right\rangle{\rm d}u$
$\nabla^{M}\widehat{\bm{\pi}}_{\ell}(\bar{\omega})$ $:=\sum_{k=1}^{L}\int_{0}^{1}\mathbb{E}_{n}\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial\mu_{k}}\Big|_{\omega=(\pi^{*},M_{u},\bar{J})},\bar{\mu}_{k}-\mu_{k}^{*}\right\rangle{\rm d}u$
$\nabla^{J}\widehat{\bm{\pi}}_{\ell}(\bar{\omega})$ $:=\int_{0}^{1}\mathbb{E}_{n}\left\langle\frac{\partial\gamma_{\ell}(X;\omega)}{\partial J}\Big|_{\omega=(\pi^{*},M^{*},J_{u})},\bar{J}-J^{*}\right\rangle{\rm d}u.$
By repeating the same argument of proving Sections C.3.1, C.34 and C.41, we have
$|\nabla^{\pi}\widehat{\bm{\pi}}_{\ell}(\bar{\omega})|$ $~\lesssim~\pi_{\ell}^{*}~\widehat{\mathcal{R}}_{I}^{(0)}~d(\bar{\pi},\pi^{*}),$
$|\nabla^{M}\widehat{\bm{\pi}}_{\ell}(\bar{\omega})|$ $~\lesssim~\pi_{\ell}^{*}~\widehat{\mathcal{R}}_{I}^{(0)}\sqrt{\Delta_{\max}}~d(\bar{M},M^{*}),$
$|\nabla^{J}\widehat{\bm{\pi}}_{\ell}(\bar{\omega})|$ $~\lesssim~\pi_{\ell}^{*}\left(\widehat{\mathcal{R}}_{I}^{(0)}\sqrt{\Delta_{\max}}+\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(2)}}\right)d(\bar{J},J^{*})$
so that by recalling (C.9),
${1\over\pi_{\ell}^{*}}|\widehat{\bm{\pi}}_{\ell}(\bar{\omega})-\widehat{\bm{\pi}}_{\ell}(\omega^{*})|~\lesssim~\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(2)}}d(\bar{J},J^{*})+\widehat{\mathcal{R}}_{I}^{(0)}d(\bar{\omega},\omega^{*}).$
Contraction of $\widehat{\mu}_{\ell}$.
Similar to (C.42), we have
$\widehat{\bm{\mu}}_{\ell}(\bar{\omega})-\widehat{\bm{\mu}}_{\ell}(\omega^{*})=\nabla^{\pi}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})+\nabla^{M}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})+\nabla^{J}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})$
where
$\nabla^{\pi}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})$ $:=\sum_{k=1}^{L-1}\int_{0}^{1}\frac{1}{\widehat{\bm{\pi}}_{\ell}(\omega)}\mathbb{E}_{n}\left[\left(X-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial\pi_{k}}\Big|_{\omega=(\pi_{u},\bar{M},\bar{J})},~\left(\bar{\pi}_{k}-\pi_{k}^{*}\right)\right\rangle\right]{\rm d}u$
$\nabla^{M}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})$ $:=\sum_{k=1}^{L}\int_{0}^{1}\frac{1}{\widehat{\bm{\pi}}_{\ell}(\omega)}\mathbb{E}_{n}\left[\left(X-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial\mu_{k}}\Big|_{\omega=(\pi^{*},M_{u},\bar{J})},~\bar{\mu}_{k}-\mu_{k}^{*}\right\rangle\right]{\rm d}u$
$\nabla^{J}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})$ $:=\int_{0}^{1}\frac{1}{\widehat{\bm{\pi}}_{\ell}(\omega)}\mathbb{E}_{n}\left[\left(X-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right)\left\langle{\partial\gamma_{\ell}(X;\omega)\over\partial J}\Big|_{\omega=(\pi^{*},M^{*},J_{u})},~\bar{J}-J^{*}\right\rangle\right]{\rm d}u.$
Note that on the event (D.17), for any $\omega\in\mathcal{B}_{d}(\omega^{*})$,
$d(\widehat{\bm{\pi}}(\omega),\bm{\pi}(\omega^{*}))$ $\leq d(\widehat{\bm{\pi}}(\omega),\widehat{\bm{\pi}}(\omega^{*}))+d(\widehat{\bm{\pi}}(\omega^{*}),\bm{\pi}(\omega^{*}))$
$\leq\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(2)}}d(\bar{J},J^{*})+\widehat{\mathcal{R}}_{I}^{(0)}d(\omega,\omega^{*})+d(\widehat{\bm{\pi}}(\omega^{*}),\bm{\pi}(\omega^{*}))$
$\leq 1/2$ by (C.9) & (C.44)
which further ensures
$1/2\leq\widehat{\bm{\pi}}_{\ell}(\omega)/\pi_{\ell}^{*}\leq 3/2,\qquad\forall~\ell\in[L].$
Meanwhile the event (D.20) implies
$\|\widehat{\bm{\mu}}_{\ell}(\omega^{*})-\mu_{b}^{*}\|_{\Sigma^{*}}\leq\sqrt{\Delta_{\max}}+d(\widehat{\mathbf{M}}(\omega^{*}),\mathbf{M}(\omega^{*}))\leq 2\sqrt{\Delta_{\max}}.$ (D.22)
By similar arguments of proving Eqs. C.47, C.49 and C.51, we find that
$\left\|\nabla^{\pi}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim~\left[\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(1)}}+\widehat{\mathcal{R}}_{I}^{(0)}\sqrt{\Delta_{\max}}\right]d(\bar{\pi},\pi^{*})$
$\left\|\nabla^{M}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim~\left[\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(1)}}+\widehat{\mathcal{R}}_{I}^{(0)}\Delta_{\max}\right]d(\bar{M},M^{*})$
$\left\|\nabla^{J}\widehat{\bm{\mu}}_{\ell}(\bar{\omega})\right\|_{\Sigma^{*}}$ $\lesssim~\left[\sqrt{\widehat{\mathcal{R}}_{I}^{(1)}\widehat{\mathcal{R}}_{I}^{(2)}}+\left(\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(2)}}+\widehat{\mathcal{R}}_{I}^{(0)}\sqrt{\Delta_{\max}}\right)\sqrt{\Delta_{\max}}\right]d(\bar{J},J^{*}).$
Therefore, similar as (C.52),
$\left\|\widehat{\bm{\mu}}_{\ell}(\bar{\omega})-\widehat{\bm{\mu}}_{\ell}(\omega^{*})\right\|_{\Sigma^{*}}$~ $\leq~\sqrt{\widehat{\mathcal{R}}_{I}^{(1)}\widehat{\mathcal{R}}_{I}^{(2)}}d(\bar{J},J^{*})+\left(\sqrt{\widehat{\mathcal{R}}_{I}^{(0)}\widehat{\mathcal{R}}_{I}^{(2)}}+\widehat{\mathcal{R}}^{(0)}_{I}\sqrt{\Delta_{\max}}\right)d(\bar{\omega},\omega^{*}),$
as desired.
Contraction of $\bm{\widehat{\Sigma}}$.
Since
$\bm{\widehat{\Sigma}}(\bar{\omega})=\widehat{\Sigma}_{T}-\sum_{\ell=1}^{L}\widehat{\bm{\pi}}_{\ell}(\bar{\omega})\widehat{\bm{\mu}}_{\ell}(\bar{\omega})\widehat{\bm{\mu}}_{\ell}(\bar{\omega})^{\top}$
and using (D.19) gives
[…]
invoking part (a) of Lemma 18 with $\theta=(\widehat{\bm{\pi}}(\bar{\omega}),\widehat{\mathbf{M}}(\bar{\omega}),\bm{\widehat{\Sigma}}(\bar{\omega}))$ and $\theta^{\prime}=(\widehat{\bm{\pi}}(\omega^{*}),\widehat{\mathbf{M}}(\omega^{*}),\bm{\widehat{\Sigma}}(\omega^{*}))$ gives
$d(\bm{\widehat{\Sigma}}(\bar{\omega}),\bm{\widehat{\Sigma}}(\omega^{*}))~\leq 3\max_{\ell\in[L]}\|\widehat{\bm{\mu}}_{\ell}(\omega^{*})\|_{\Sigma^{*}}d(\widehat{\mathbf{M}}(\bar{\omega}),\widehat{\mathbf{M}}(\omega^{*}))+\max_{\ell\in[L]}2\|\widehat{\bm{\mu}}_{\ell}(\omega^{*})\|_{\Sigma^{*}}^{2}d(\widehat{\bm{\pi}}(\bar{\omega}),\widehat{\bm{\pi}}(\omega^{*})).$
Since the event (D.20) ensures
$\|\widehat{\bm{\mu}}_{\ell}(\omega^{*})\|_{\Sigma^{*}}\leq\|\bm{\mu}_{\ell}(\omega^{*})\|_{\Sigma^{*}}+d(\widehat{\mathbf{M}}(\omega^{*}),\mathbf{M}(\omega^{*}))\leq 2\sqrt{\Delta_{\max}}$
invoking the contraction results of $\widehat{\pi}$ and $\widehat{M}$ completes the proof. ∎
D.2.1 Deviation inequalities of $\widehat{\mathcal{R}}_{I}^{(j)}$ for $j\in\{0,1,2\}$
The following lemma states stochastic bounds for $\widehat{\mathcal{R}}_{I}^{(j)}$, $j\in\{0,1,2\}$, given by (D.15).
Lemma 24.
With probability at least $1-n^{-dL}$,
$\widehat{\mathcal{R}}_{I}^{(0)}$~ $\lesssim~{L^{2}\over\pi_{\min}^{*}}\exp(-c\Delta_{\min})+{dL\log(q)\over n\pi_{\min}^{*}},$ (D.23)
$\widehat{\mathcal{R}}_{I}^{(1)}$~ $\lesssim~{L\over\pi_{\min}^{*2}}\exp(-c\Delta_{\min})+{dL\log(q)\log(n)\over n\pi_{\min}^{*}},$ (D.24)
$\widehat{\mathcal{R}}_{I}^{(2)}$~ $\lesssim~{L^{2}\over\pi_{\min}^{*2}}\exp(-c\Delta_{\min})+{dL^{2}\log(q)\log(n)\over n\pi_{\min}^{*}}$ (D.25)
We prove Lemma 24 in the following a few subsections. Since it requires to establish uniform convergence over $\omega\in\mathcal{B}_{d}(\omega^{*})$, our proof is based on a discretization argument of this ball. We start by constructing an $\epsilon$-net of $\mathcal{B}_{d}(\omega^{*})$ in the following subsection.
D.2.2 An $\epsilon$-net construction of $\mathcal{B}_{d}(\omega^{*})$
Let $\mathcal{N}_{\epsilon_{1}}$ be an $\epsilon_{1}$-net of $\{v\in\mathbb{R}^{d}:\|v\|_{2}\leq c_{0}\sqrt{\Delta_{\min}}\}$. Define
$\mathcal{N}_{\epsilon_{1}}(\mu_{1}^{*})=\left\{\mu_{1}^{*}+\Sigma^{*1/2}v:v\in\mathcal{N}_{\epsilon_{1}}\right\}.$
Then for any $\mu$ such that $\|\mu-\mu_{1}^{*}\|_{\Sigma^{*}}\leq c_{0}\sqrt{\Delta_{\min}}$, there exists $v_{\mu}\in\mathcal{N}_{\epsilon_{1}}$ such that $\|\Sigma^{*-1/2}(\mu-\mu_{1}^{*})-v_{\mu}\|_{2}\leq\epsilon_{1}$ which further implies that $\mu^{\prime}:=\mu_{1}^{*}+\Sigma^{*1/2}v_{\mu}\in\mathcal{N}_{\epsilon_{1}}(\mu_{1}^{*})$ satisfies $\|\mu-\mu^{\prime}\|_{\Sigma^{*}}=\|\Sigma^{*-1/2}(\mu-\mu_{1}^{*})-v_{\mu}\|_{2}\leq\epsilon_{1}.$ Therefore, $\mathcal{N}_{\epsilon_{1}}(\mu_{1}^{*})$ is an $\epsilon_{1}$-net of $\{\mu:\|\mu-\mu_{1}^{*}\|_{\Sigma^{*}}\leq c_{0}\sqrt{\Delta_{\min}}\}$ with respect to $\|\cdot\|_{\Sigma^{*}}$. Moreover
$\mathcal{N}_{\epsilon_{1}}(M^{*})=\mathcal{N}_{\epsilon_{1}}(\mu_{1}^{*})\otimes\cdots\otimes\mathcal{N}_{\epsilon_{1}}(\mu_{L}^{*})$
is an $\epsilon_{1}$-net of $\{M\in\mathbb{R}^{d\times L}:d(M,M^{*})\leq c_{0}\sqrt{\Delta_{\min}}\}$ with respect to the distance $d(M,M^{\prime})$ in (3.2). By construction, we have $d(M,M^{*})\leq c_{0}\sqrt{\Delta_{\min}}$ for any $M\in\mathcal{N}_{\epsilon_{1}}(M^{*})$.
By analogous argument, we can construct an $\epsilon_{2}$-net of $\{J\in\mathbb{R}^{d\times L}:d(J,J^{*})\leq c_{0}\sqrt{\Delta_{\min}}\}$ with respect to $d(J,J^{\prime})$ in (C.8),
$\mathcal{N}_{\epsilon_{2}}(J^{*})=\mathcal{N}_{\epsilon_{2}}(J_{1}^{*})\otimes\cdots\otimes\mathcal{N}_{\epsilon_{2}}(J_{L}^{*})$
with
$\mathcal{N}_{\epsilon_{2}}(J_{a}^{*})=\left\{J_{a}^{*}+\Sigma^{*-1/2}v:v\in\mathcal{N}_{\epsilon_{2}}\right\},\quad\forall~a\in[L].$
Then for any $J_{a}$ such that $\|\Sigma^{*1/2}(J_{a}-J^{*}_{a})\|_{2}\leq c_{0}\sqrt{\Delta_{\min}}$, there exists $v\in\mathcal{N}_{\epsilon_{2}}$ such that $\|\Sigma^{*1/2}(J_{a}-J_{a}^{*})-v\|_{2}\leq\epsilon_{2}$ which further implies that $J_{a}^{\prime}:=J_{a}^{*}+\Sigma^{*-1/2}v\in\mathcal{N}_{\epsilon_{2}}(J_{a}^{*})$ satisfies $\|\Sigma^{*1/2}(J_{a}-J_{a}^{\prime})\|_{2}=\|\Sigma^{*1/2}(J_{a}-J_{a}^{*})-v\|_{2}\leq\epsilon_{2}.$
Finally, regarding $\pi$, there exists an $\epsilon_{3}$-net $\mathcal{N}_{\epsilon_{3}}(\mathcal{S}^{L})$ of the simplex in $\mathbb{R}^{L}$ with respect to the sup-norm [ghosal2001entropies][Lemma A.4]. As a result, for any $\pi$ such that $d(\pi,\pi^{*})\leq c_{0}$, there exists $\pi^{\prime}\in\mathcal{N}_{\epsilon_{3}}(\mathcal{S}^{L})$ such that
$d(\pi,\pi^{\prime})\leq{\epsilon_{3}\over\pi_{\min}^{*}},\qquad d_{\pi}(\pi^{\prime},\pi^{*})\leq c_{0}+{\epsilon_{3}\over\pi_{\min}^{*}}\leq 2c_{0}$
provided that $\epsilon_{3}\leq c_{0}\pi_{\min}^{*}$.
Combining all three nets constructed above, we have
$|\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}|:=\left|\mathcal{N}_{\epsilon_{1}}(M^{*})\otimes\mathcal{N}_{\epsilon_{2}}(J^{*})\otimes\mathcal{N}_{\epsilon_{3}}(\mathcal{S}^{L})\right|\leq\left(5\over\epsilon_{3}\right)^{L-1}\left(3c_{0}\sqrt{\Delta_{\min}}\over\epsilon_{1}\epsilon_{2}\right)^{dL}.$ (D.26)
D.2.3 Proof of Lemma 24 for $j=0$
Proof.
The entire proof is based on the event $\mathcal{E}_{2}$ given in (D.5). By the definitions in (D.15) and (C.23), we have
$\widehat{\mathcal{R}}_{I}^{(0)}$ $=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}{1\over\pi_{\ell}}\mathbb{E}_{n}\left[I_{\ell}(X,\omega)\right]$
where we write
$I_{\ell}(X,\omega)=\pi_{\ell}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}I_{a\ell}(X,\omega).$ (D.27)
We use a discretization argument to handle the supremum. Recall $\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}$ from (D.26). For any $\omega\in\mathcal{B}_{d}(\omega^{*})$, there exists some $\omega^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}$ such that for all $\ell\in[L]$,
$\mathbb{E}_{n}\left[I_{\ell}(X,\omega)\right]~\leq~\mathbb{E}_{n}\left[I_{\ell}(X,\omega^{\prime})\right]+\left|\mathbb{E}_{n}\left[I_{\ell}(X,\omega)-I_{\ell}(X,\omega^{\prime})\right]\right|.$
By invoking Lemma 25 and choosing
$\epsilon_{1}=\frac{dL}{n\sqrt{\Delta_{\max}}},\qquad\epsilon_{2}=\frac{L}{n}\sqrt{d\log n\over\Delta_{\max}},\qquad\epsilon_{3}=\pi_{\min}^{*}{dL\over n},$
one has that for all $\omega\in\mathcal{B}_{d}(\omega^{*})$,
$\max_{\ell\in[L]}\max_{i\in[n]}\left|I_{\ell}(X_{i},\omega)-I_{\ell}(X_{i},\omega^{\prime})\right|$ $\lesssim\epsilon_{2}\sqrt{d+\log n+\Delta_{\max}}+\epsilon_{1}\sqrt{\Delta_{\max}}+\frac{\epsilon_{3}}{\pi_{\min}^{*}}$
$\lesssim{dL\log n\over n}.$ (D.28)
Thus, for all $\ell\in[L]$
$\left|\mathbb{E}_{n}\left[I_{\ell}(X,\omega)-I_{\ell}(X,\omega^{\prime})\right]\right|$ $\lesssim{dL\log n\over n}.$ (D.29)
We proceed to bound $\max_{\ell\in[L]}\mathbb{E}_{n}[I_{\ell}(X,\omega)]$ over $\omega\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}$. For any $\omega\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}$, define the set
$S(\omega)=\left\{i\in[n]:\max_{k\neq\ell\in[L]}\left(|N_{i}^{\top}\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})|-\nu\Delta_{k\ell}\right)>0\right\}$ (D.30)
for some constant $0<\nu<(1/2-2c_{0}-2c_{0}^{2})$. Further let $S(\omega)]^{c}:=[n]\setminus S(\omega)$ be its complement set.
Recall the model (1.1), we partition $n]=\bigcup_{a=1}^{L}\widehat{G}_{a}$ with
$\widehat{G}_{a}=\left\{i\in[n]:Y_{i}=a\right\},\qquad n_{a}:=|\widehat{G}_{a}|,\qquad\forall\ a\in[L].$ (D.31)
Further let
$\widetilde{\pi}_{\ell}:={n_{\ell}\over n},\qquad\forall~\ell\in[L].$ (D.32)
We first observe that for any $i\in\widehat{G}_{\ell}\cap[S(\omega)]^{c}$, using (C.54), Lemma 20 and the choice of $\nu$ gives
$I_{\ell}(X_{i},\omega)\leq{1\over\pi_{\ell}}\sum_{a\in[L]\setminus\{\ell\}}\pi_{a}\exp(N_{i}^{\top}\sigma_{a\ell}(\omega)+\delta_{a\ell}^{\ell}(\omega))\leq{1\over\pi_{\ell}}\exp(-c^{\prime}\Delta_{\min}).$ (D.33)
Similarly, for any $b\in[L]\setminus\{\ell\}$ and $i\in\widehat{G}_{b}\cap[S(\omega)]^{c}$, we have
[…]
Combining these two gives
${1\over n}\sum_{i\in[S(\omega)]^{c}}I_{\ell}(X_{i},\omega)$ $\leq{1\over n}\sum_{i\in[S(\omega)]^{c}}\left(1\{Y_{i}=\ell\}+\sum_{b\in[L]\setminus\{\ell\}}1\{Y_{i}=b\}\right)I_{\ell}(X_{i},\omega)$
$\leq{n_{\ell}\over n\pi_{\ell}}\exp(-c^{\prime}\Delta_{\min})+\sum_{b\in[L]\setminus\{\ell\}}{n_{b}\over n}{\pi_{\ell}\over\pi_{b}}\exp(-c^{\prime}\Delta_{\min})$
$\lesssim(1+L\pi_{\ell})\exp(-c^{\prime}\Delta_{\min}).$ (D.35)
On the other hand, by letting
$Z_{i}(\omega):=1\left\{\max_{k\neq\ell\in[L]\setminus\{\ell\}}\left(|N_{i}^{\top}\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})|-\nu\Delta_{k\ell}\right)>0\right\}$
we know that $Z_{1}(\omega),\ldots,Z_{n}(\omega)$ are i.i.d. with
$\mathbb{E}[Z_{i}(\omega)]$ $\leq 2\sum_{k\neq\ell\in[L]}\mathbb{P}\left(N_{i}^{\top}\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})>\nu\Delta_{k\ell}\right)$
$\leq~2\sum_{k\neq\ell\in[L]}\exp\left(-\nu^{2}\Delta_{k\ell}^{2}/\|\Sigma^{*1/2}J({\bm{e}}_{k}-{\bm{e}}_{\ell})\|_{2}^{2}\right)$
[…]
An application of the Bernstein inequality together with taking the union bound over $\omega\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}$ gives that with probability at least $1-n^{-dL}$,
${1\over n}\left|\sum_{i\in[n]}\left(Z_{i}(\omega)-\mathbb{E}[Z_{i}(\omega)]\right)\right|\lesssim\sqrt{dL^{3}\log q\over n}\exp(-c\Delta_{\min})+{dL\log q\over n}$ (D.37)
holds for all $\omega\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}$. Here we let $q:=n(\pi_{\min}^{*-1}\vee\Delta_{\max})$, With the same probability, we therefore have
$\max_{\omega\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}}~|S(\omega)|$ $\leq\sqrt{ndL^{3}\log(q)}\exp(-c\Delta_{\min})+dL\log(q)+\max_{\omega\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}}\ \mathbb{E}[Z_{i}(\omega)]$
$\lesssim~dL\log(q)+nL^{2}\exp(-c\Delta_{\min}).$ (D.38)
The last step uses $n\geq d$, (D.2.3) and (3.4). Armed with (D.2.3) and (D.38), we find that
$\max_{\omega\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}}\max_{\ell\in[L]}{1\over\pi_{\ell}}\mathbb{E}_{n}\left[I_{\ell}(X,\omega)\right]$ $\leq\max_{\omega\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}}\max_{\ell\in[L]}{1\over n\pi_{\ell}}\left\{\sum_{i\in S(\omega)}+\sum_{i\in[S(\omega)]^{c}}\right\}I_{\ell}(X_{i},\omega)$
$\leq\max_{\omega\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}}\max_{\ell\in[L]}\left\{{|S(\omega)|\over n\pi_{\ell}}+\left({1\over\pi_{\ell}}+L\right)\exp(-c^{\prime}\Delta_{\min})\right\}$
$\lesssim~{L^{2}\over\pi_{\min}^{*}}\exp(-c\Delta_{\min})+{dL\log(q)\over n\pi_{\min}^{*}}.$
The last step also uses (C.28). This completes the proof of (D.23). ∎
The following lemma controls the discretization error in $I_{\ell}(X_{i},\omega)$.
Lemma 25.
For any $\omega\in\mathcal{B}_{d}(\omega^{*})$ and $\omega^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}$, on the event $\mathcal{E}_{2}$, we have
$\max_{\ell\in[L]}\max_{i\in[n]}~\left|I_{\ell}(X_{i},\omega)-I_{\ell}(X_{i},\omega^{\prime})\right|\lesssim\epsilon_{2}\sqrt{d+\log n+\Delta_{\max}}+\epsilon_{1}\sqrt{\Delta_{\max}}+\frac{\epsilon_{3}}{\pi_{\min}^{*}}.$
Proof.
Recall from (D.27) and (C.23) that
[…]
It is easy to see that
$\left|I_{\ell}(x,\omega)-I_{\ell}(x,\omega^{\prime})\right|$ $\leq 3\left|\gamma_{\ell}(x;\omega)-\gamma_{\ell}(x;\omega^{\prime})\right|.$
Define
$f_{h}^{(\ell)}(x,\omega)=(x-\frac{1}{2}(\mu_{h}+\mu_{\ell}))^{\top}J({\bm{e}}_{h}-{\bm{e}}_{\ell})+\log{\frac{\pi_{h}}{\pi_{\ell}}},\quad\forall~h\in[L].$
For any $a\in\mathbb{R}^{L}$ with $a_{\ell}=0$, define $\psi_{\ell}(a)=1/(\sum_{h=1}^{L}e^{a_{h}}).$ Since
$\psi_{\ell}(a)-\psi_{\ell}(b)$ $=\frac{\sum_{h}e^{a_{h}}(1-e^{b_{h}-a_{h}})}{\sum_{h}e^{a_{h}}\sum_{h}e^{b_{h}}}$
$\leq\frac{\sum_{h}e^{a_{h}}|b_{h}-a_{h}|}{\sum_{h}e^{a_{h}}\sum_{h}e^{b_{h}}}$ $\text{by }1-e^{x}\leq|x|\text{ and }e^{b_{\ell}}=1$
$\leq\max_{h\in[L]}|a_{h}-b_{h}|,$
applying it with $a=f^{(\ell)}(X,\omega)$ and $b=f^{(\ell)}(X,\omega^{\prime})$ gives
$\left|\gamma_{\ell}(X,\omega)-\gamma_{\ell}(X,\omega^{\prime})\right|$ $=\left|\psi(f^{(\ell)}(X,\omega))-\psi(f^{(\ell)}(X,\omega^{\prime}))\right|\leq\max_{h\in[L]}|f^{(\ell)}_{h}(X,\omega)-f^{(\ell)}_{h}(X,\omega^{\prime})|.$
By using (C.38), we have
$|f_{h}^{(\ell)}(X,\omega)-f^{(\ell)}_{h}(X,\omega^{\prime})|$
$\leq\left|N^{\top}\Sigma^{*1/2}(J-J^{\prime})({\bm{e}}_{h}-{\bm{e}}_{\ell})\right|+\max_{b\in[L]}\left|\left(\mu_{b}^{*}-\frac{1}{2}(\mu_{h}+\mu_{\ell})\right)^{\top}(J-J^{\prime})({\bm{e}}_{h}-{\bm{e}}_{\ell})\right|$
$\quad+\frac{1}{2}\left|(\mu_{h}-\mu_{h}^{\prime}+\mu_{\ell}-\mu_{\ell}^{\prime})^{\top}J({\bm{e}}_{h}-{\bm{e}}_{\ell}^{\prime})\right|+\left|\log(\pi_{h}/\pi_{h}^{\prime})-\log(\pi_{\ell}/\pi_{\ell}^{\prime})\right|$
$\lesssim\epsilon_{2}\|N\|_{2}+\sqrt{\Delta_{\max}}(\epsilon_{1}+\epsilon_{2})+\frac{\epsilon_{3}}{\pi_{\min}}$
where the last step uses (C.39), (C.32) and
$\left|\log\frac{\pi_{h}}{\pi_{h}^{\prime}}\right|\leq\frac{|\pi_{h}-\pi^{\prime}_{h}|}{\pi_{h}\wedge\pi^{\prime}_{h}}\lesssim{\epsilon_{3}\over\pi_{\min}}$
due to the concavity of $x\mapsto\log(x)$. Invoking $\mathcal{E}_{2}$ and (C.28) completes the proof. ∎
D.2.4 Proof of Lemma 24 for $j=1$
Proof.
The proof is based on $\mathcal{E}_{2}$ and
$\mathcal{E}_{{\rm op}}:=\left\{{1\over n}\left\|\sum_{i=1}^{n}N_{i}N_{i}^{\top}\right\|_{\rm op}\leq{2d\over n}+12\right\}$
which according to Lemma 32 holds with probability $1-e^{-n}$. By definition in (D.15) and (D.27), we have
$\widehat{\mathcal{R}}_{I}^{(1)}$ $=\sup_{\omega\in\mathcal{B}_{d}(\omega^{*})}\max_{\ell\in[L]}{1\over\pi_{\ell}}\left\|\mathbb{E}_{n}\left[I_{\ell}(X,\omega)NN^{\top}\right]\right\|_{\rm op}.$
Since for any $\omega\in\mathcal{B}_{d}(\omega^{*})$, there exists some $\omega^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}$ such that
$\left\|\mathbb{E}_{n}\left[I_{\ell}(X,\omega)NN^{\top}\right]\right\|_{\rm op}-\left\|\mathbb{E}_{n}\left[I_{\ell}(X,\omega^{\prime})NN^{\top}\right]\right\|_{\rm op}$
$\leq{1\over n}\left\|\sum_{i=1}^{n}N_{i}N_{i}^{\top}\right\|_{\rm op}\max_{i\in[n]}\left|I_{\ell}(X_{i},\omega)-I_{\ell}(X_{i},\omega^{\prime})\right|$
[…]
it suffices to bound, for all $\omega\in\mathcal{N}_{\epsilon_{1},\epsilon_{2},\epsilon_{3}}$ and $\ell\in[L]$,
$\left\|\mathbb{E}_{n}\left[I_{\ell}(X,\omega)NN^{\top}\right]\right\|_{{\rm op}}$ $\leq{1\over n}\left\|\sum_{i\in S(\omega)}I_{\ell}(X_{i},\omega)N_{i}N_{i}^{\top}\right\|_{\rm op}+{1\over n}\left\|\sum_{i\in[S(\omega)]^{c}}I_{\ell}(X_{i},\omega)N_{i}N_{i}^{\top}\right\|_{\rm op}$
$\leq{1\over n}\left\|\sum_{i\in S(\omega)}N_{i}N_{i}^{\top}\right\|_{\rm op}+{1\over n}\left\|\sum_{i=1}^{n}N_{i}N_{i}^{\top}\right\|_{\rm op}\max_{i\in[S(\omega)]^{c}}I_{\ell}(X_{i},\omega)$
$\lesssim{1\over n}\left\|\sum_{i\in S(\omega)}N_{i}N_{i}^{\top}\right\|_{\rm op}+\left({1\over\pi_{\ell}}+{\pi_{\ell}\over\pi_{\min}}\right)\exp(-c^{\prime}\Delta_{\min}).$
The last step uses $\mathcal{E}_{{\rm op}}$, (D.33) and (D.34). Let $m:=C(dL\log(q)+nL\exp(-c\Delta_{\min}))$ for some large constant $C>0$. It suffices to consider $m\leq n$ as we can set $m=n$ otherwise. Invoking Lemma 32 for $S=S(\omega)\leq m$ together with (D.38) gives that with probability at least $1-e^{-m}\geq 1-n^{-dL}$,
${1\over n}\left\|\sum_{i\in S(\omega)}N_{i}N_{i}^{\top}\right\|_{\rm op}$ $\leq{2d+12m\log(en/m)\over n}\lesssim{dL\log(q)\log(n)\over n}+L^{2}\exp(-c\Delta_{\min}),$
thereby completing the proof. ∎
D.2.5 Proof of Lemma 24 for $j=2$
Proof.
By repeating the same argument of bounding $\mathcal{R}_{I}^{(2)}$ in (22) with $\mathbb{E}$ replaced by $\mathbb{E}_{n}$, we have $\widehat{\mathcal{R}}_{I}^{(2)}\leq(L-1)\widehat{\mathcal{R}}_{I}^{(1)}$. ∎
D.3 Proof of Theorem 7 & Theorem 8
Proof.
For any $t\geq 1$ and each $h\in\{\pi,M,\Sigma\}$, triangle inequality gives
$d(\widehat{h}^{(t)},\bm{h}(\theta^{*}))$ $\leq d\left(\widehat{\bm{h}}(\widehat{\theta}^{(t-1)}),\widehat{\bm{h}}(\theta^{*})\right)+d\left(\widehat{\bm{h}}(\theta^{*}),\bm{h}(\theta^{*})\right)$
$=d\left(\widehat{\bm{h}}(\widehat{\omega}^{(t-1)}),\widehat{\bm{h}}(\omega^{*})\right)+d\left(\widehat{\bm{h}}(\omega^{*}),\bm{h}(\omega^{*})\right)$
where in the last step we also re-parametrize by using $\widehat{\omega}^{(t-1)}:=\omega(\widehat{\theta}^{(t-1)})$. Provided that $\widehat{\theta}^{(t-1)}$ satisfies (3.5), its reparametrization $\widehat{\omega}^{(t-1)}\in\mathcal{B}_{d}(\omega^{*})$ given by (C.2). We work on the event under which Theorem 5 and Lemma 24 holds and for notational simplicity, define
$\epsilon_{\pi}=C\sqrt{\log n\over n\pi_{\min}^{*}},\quad\epsilon_{M}=C\sqrt{d\log n\over n\pi_{\min}^{*}},\quad\epsilon_{\Sigma}=C\sqrt{d\log n\over n}.$
For known $\Sigma^{*}$, invoking Theorem 23 and Lemma 17 thus gives
$d(\widehat{\bm{\pi}}(\widehat{\omega}^{(t-1)}),\widehat{\bm{\pi}}(\omega^{*}))$ $\leq~(\kappa/3)\left\{d(\widehat{\pi}^{(t-1)},\pi^{*})+d(\widehat{M}^{(t-1)},M^{*})+d(J(\widehat{\theta}^{(t-1)}),J(\theta^{*}))\right\}$
$\leq~\kappa\left\{d(\widehat{\pi}^{(t-1)},\pi^{*})+d(\widehat{M}^{(t-1)},M^{*})\right\}$ (D.39)
where
[…]
We note $\log q\asymp\log n$ under (4.5). Similarly, we have
$d(\widehat{\mathbf{M}}(\widehat{\omega}^{(t-1)}),\widehat{\mathbf{M}}(\omega^{*}))$ $\leq\kappa\sqrt{\Delta_{\max}}\left\{d(\widehat{\pi}^{(t-1)},\pi^{*})+d(\widehat{M}^{(t-1)},M^{*})\right\}.$ (D.40)
Together with Theorem 5, we find that
$d(\widehat{\pi}^{(t)},\pi^{*})$ $\leq\epsilon_{\pi}+\kappa\left\{d(\widehat{\pi}^{(t-1)},\pi^{*})+d(\widehat{M}^{(t-1)},M^{*})\right\}$
$\leq\epsilon_{\pi}+(1/2+c_{0}\sqrt{\Delta_{\min}})\kappa$
$\leq 1/2$ (D.41)
where the last step uses (3.5) with $c_{\Sigma}=0$ and conditions (3.4) & (4.5). Similarly,
$d(\widehat{M}^{(t)},M^{*})$ $\leq\epsilon_{M}+\kappa\sqrt{\Delta_{\max}}\left\{d(\widehat{\pi}^{(t-1)},\pi^{*})+d(\widehat{M}^{(t-1)},M^{*})\right\}$
$\leq\epsilon_{M}+(1/2+c_{0}\sqrt{\Delta_{\min}})\kappa\sqrt{\Delta_{\max}}$
$\leq c_{0}\sqrt{\Delta_{\min}}.$ (D.42)
The last two displays ensure that $\widehat{\theta}^{(t)}$ satisfies (3.5) hence $\widehat{\omega}^{(t)}\in\mathcal{B}_{d}(\omega^{*})$. By Sections D.3 and D.40 and induction, we obtain
$d(\widehat{\pi}^{(t)},\pi^{*})$ $\leq\epsilon_{\pi}+\kappa\left\{d(\widehat{\pi}^{(t-1)},\pi^{*})+d(\widehat{M}^{(t-1)},M^{*})\right\}$
$\leq\epsilon_{\pi}+\kappa(\epsilon_{\pi}+\epsilon_{M})+\kappa^{2}(1+\sqrt{\Delta_{\max}})\left\{d(\widehat{\pi}^{(t-2)},\pi^{*})+d(\widehat{M}^{(t-2)},M^{*})\right\}$
$\leq\epsilon_{\pi}+(\epsilon_{\pi}+\epsilon_{M})\kappa\left[1+\kappa(1+\sqrt{\Delta_{\max}})+\cdots+\kappa^{t-2}(1+\sqrt{\Delta_{\max}})^{t-2}\right]$
$\qquad+\kappa^{t}(1+\sqrt{\Delta_{\max}})^{t-1}\left\{d(\widehat{\pi}^{(0)},\pi^{*})+d(\widehat{M}^{(0)},M^{*})\right\}$
$\leq\epsilon_{\pi}+{\kappa\over 1-\kappa-\kappa\sqrt{\Delta_{\max}}}(\epsilon_{\pi}+\epsilon_{M})+(2\kappa\sqrt{\Delta_{\max}})^{t}\left\{d(\widehat{\pi}^{(0)},\pi^{*})+d(\widehat{M}^{(0)},M^{*})\right\}$
as well as, by using (D.40),
$d(\widehat{M}^{(t)},M^{*})$ $\leq\epsilon_{M}+\kappa\sqrt{\Delta_{\max}}\left\{d(\widehat{\pi}^{(t-1)},\pi^{*})+d(\widehat{M}^{(t-1)},M^{*})\right\}$
$\leq\epsilon_{M}+{\kappa\sqrt{\Delta_{\max}}\over 1-2\kappa\sqrt{\Delta_{\max}}}(\epsilon_{\pi}+\epsilon_{M})+(2\kappa\sqrt{\Delta_{\max}})^{t}\left\{d(\widehat{\pi}^{(0)},\pi^{*})+d(\widehat{M}^{(0)},M^{*})\right\}.$
Therefore, with $\kappa_{n}=2\kappa\sqrt{\Delta_{\max}}$, one has that for all $t\geq 1$,
$d(\widehat{\pi}^{(t)},\pi^{*})+d(\widehat{M}^{(t)},M^{*})$ $~\lesssim~\sqrt{d\log n\over n\pi_{\min}^{*}}+\kappa_{n}^{t}\left\{d(\widehat{\pi}^{(0)},\pi^{*})+d(\widehat{M}^{(0)},M^{*})\right\},$
completing the proof of Theorem 7.
For unknown $\Sigma^{*}$, denote the distance between the initialization $\widehat{\theta}^{(0)}$ and $\theta^{*}$ by
$d(\widehat{\theta}^{(0)},\theta^{*}):=d(\widehat{\pi}^{(0)},\pi^{*})+d(\widehat{M}^{(0)},M^{*})+\sqrt{\Delta_{\max}}~d(\widehat{\Sigma}^{(0)},\Sigma^{*}).$ (D.43)
Note that condition (3.5) ensures $d(\widehat{\theta}^{(0)},\theta^{*})\leq 1/2+c_{\pi}\sqrt{\Delta_{\min}}+c_{\Sigma}\sqrt{\Delta_{\max}}$. By similar arguments, we note from Theorem 23 and Lemma 17 that
$d(\widehat{\bm{\pi}}(\widehat{\omega}^{(t-1)}),\widehat{\bm{\pi}}(\omega^{*}))$ $~\leq~\kappa~d(\widehat{\theta}^{(t-1)},\theta^{*})$
$d(\widehat{\mathbf{M}}(\widehat{\omega}^{(t-1)}),\widehat{\mathbf{M}}(\omega^{*}))$ $~\leq~\kappa\sqrt{\Delta_{\max}}~d(\widehat{\theta}^{(t-1)},\theta^{*})$
$d(\bm{\widehat{\Sigma}}(\widehat{\omega}^{(t-1)}),\bm{\widehat{\Sigma}}(\omega^{*}))$ $~\leq~\kappa{\Delta_{\max}}~d(\widehat{\theta}^{(t-1)},\theta^{*}).$
Invoking Theorem 5 and using the fact that $\widehat{\theta}^{(t-1)}$ satisfies (3.5) give
$d(\widehat{\Sigma}^{(t)},\Sigma^{*})$ $\leq\epsilon_{\Sigma}+\kappa{\Delta_{\max}}~d(\widehat{\theta}^{(t-1)},\theta^{*})$
$\leq\epsilon_{\Sigma}+\kappa\Delta_{\max}\left(1/2+(c_{\mu}+c_{\Sigma})\sqrt{\Delta_{\max}}\right)$
[…]
By similar arguments, it is easy to see that (D.41) and (D.3) still hold so that $\widehat{\theta}^{(t)}$ satisfies (3.5) hence $\widehat{\omega}^{(t)}\in\mathcal{B}_{d}(\omega^{*})$. By the analogous induction argument, letting $\kappa_{n}^{\prime}=\kappa+\kappa\sqrt{\Delta_{\max}}+\kappa\Delta_{\max}^{3/2}$ and using (D.43) give
$d(\widehat{\pi}^{(t)},\pi^{*})$ $\leq\epsilon_{\pi}+{\kappa\over 1-\kappa_{n}^{\prime}}\left(\epsilon_{\pi}+\epsilon_{M}+\sqrt{\Delta_{\max}}~\epsilon_{\Sigma}\right)+(\kappa_{n}^{\prime})^{t}~d(\widehat{\theta}^{(0)},\theta^{*}),$
$d(\widehat{M}^{(t)},M^{*})$ $\leq\epsilon_{M}+{\kappa\sqrt{\Delta_{\max}}\over 1-\kappa_{n}^{\prime}}\left(\epsilon_{\pi}+\epsilon_{M}+\sqrt{\Delta_{\max}}~\epsilon_{\Sigma}\right)+(\kappa_{n}^{\prime})^{t}~d(\widehat{\theta}^{(0)},\theta^{*})$
$d(\widehat{\Sigma}^{(t)},\Sigma^{*})$ $\leq\epsilon_{\Sigma}+{\kappa\Delta_{\max}\over 1-\kappa_{n}^{\prime}}(\epsilon_{\pi}+\epsilon_{M}+\sqrt{\Delta_{\max}}~\epsilon_{\Sigma})+(\kappa_{n}^{\prime})^{t}~d(\widehat{\theta}^{(0)},\theta^{*}).$
This in conjunction with (4.7) yields
$d(\widehat{\pi}^{(t)},\pi^{*})+d(\widehat{M}^{(t)},M^{*})$ $~\lesssim~\sqrt{d\log n\over n\pi_{\min}^{*}}+(\kappa_{n}^{\prime})^{t}~d(\widehat{\theta}^{(0)},\theta^{*}),$
$d(\widehat{\Sigma}^{(t)},\Sigma^{*})$ $~\lesssim~\sqrt{d\log n\over n}+(\kappa_{n}^{\prime})^{t}~d(\widehat{\theta}^{(0)},\theta^{*}),$
thereby completing the proof of Theorem 8. ∎
D.4 Proof of Remark 4
Proof.
By (C.14), we have
$d(\widehat{\Sigma}^{(0)},\Sigma^{*})\leq 3\sqrt{\Delta_{\max}}d(\widehat{M}^{(0)},M^{*})+2\Delta_{\max}d(\widehat{\pi}^{(0)},\pi^{*})+d(\widehat{\Sigma}_{T},\Sigma_{T}^{*}).$
The claim follows by invoking Lemma 26 stated and proved below together with $\Delta_{\max}\sqrt{d\log n/n}\leq c$ under condition (4.7). ∎
Recall $\widehat{\Sigma}_{T}=n^{-1}\sum_{i=1}^{n}X_{i}X_{i}^{\top}$. The following bounds from above $d(\widehat{\Sigma}_{T},\Sigma_{T}^{*})$.
Lemma 26.
Under model (1.1), assume $d\log n\leq n$. With probability at least $1-n^{-d}$, one has
$d(\widehat{\Sigma}_{T},\Sigma_{T}^{*})~\lesssim~(\Delta_{\max}+1)\sqrt{d\log n\over n}.$
Proof.
Let $\mathcal{N}_{1/3}$ be the $(1/3)$-net of $\mathbb{S}^{d}$ and $|\mathcal{N}_{1/3}|\leq 7^{d}$. Standard discretization argument gives
$d(\widehat{\Sigma}_{T},\Sigma_{T}^{*})=\sup_{v\in\mathbb{S}^{d}}v^{\top}\Sigma^{*-1/2}(\widehat{\Sigma}_{T}-\Sigma_{T}^{*})\Sigma^{*-1/2}v\leq 3\max_{v\in\mathcal{N}_{1/3}}v^{\top}\Sigma^{*-1/2}(\widehat{\Sigma}_{T}-\Sigma_{T}^{*})\Sigma^{*-1/2}v.$
Fix any $v\in\mathcal{N}_{1/3}$. By definition, we have
$\widehat{\Sigma}_{T}$ $={1\over n}{\mathbf{X}}^{\top}{\mathbf{X}}={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}1\{Y_{i}=k\}X_{i}X_{i}^{\top}$
$={1\over n}\sum_{k=1}^{L}\left[\sum_{i=1}^{n}1\{Y_{i}=k\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}+n_{k}(\widetilde{\mu}_{k}\mu_{k}^{*\top}+\mu_{k}^{*}\widetilde{\mu}_{k}^{\top})-n_{k}\mu_{k}\mu_{k}^{\top}\right]$
$=\bar{\Sigma}+\sum_{k=1}^{L}{n_{k}\over n}\mu_{k}^{*}\mu_{k}^{*\top}+\sum_{k=1}^{L}{n_{k}\over n}\left[(\widetilde{\mu}_{k}-\mu_{k}^{*})\mu_{k}^{*\top}+\mu_{k}(\widetilde{\mu}_{k}-\mu_{k}^{*})^{\top}\right]$
where $\widetilde{\mu}_{k}$ is given in (4.10) and
$\bar{\Sigma}:={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}1\{Y_{i}=k\}(X_{i}-\mu_{k}^{*})(X_{i}-\mu_{k}^{*})^{\top},$
By further using the decomposition of $\Sigma_{T}$ in (C.7) and $\widetilde{\pi}_{k}=n_{k}/n$ for $k\in[L]$, we find that
$v^{\top}\Sigma^{*-1/2}(\widehat{\Sigma}_{T}-\Sigma_{T}^{*})\Sigma^{*-1/2}v\leq v^{\top}\Sigma^{*-1/2}(\bar{\Sigma}-\Sigma^{*})\Sigma^{*-1/2}v$
$\qquad+\sum_{k=1}^{L}(\widetilde{\pi}_{k}-\pi_{k})v^{\top}\Sigma^{*-1/2}\mu_{k}\mu_{k}^{\top}\Sigma^{*-1/2}v+2\left|\sum_{k=1}^{L}\widetilde{\pi}_{k}v^{\top}\Sigma^{*-1/2}(\widetilde{\mu}_{k}-\mu_{k}^{*})\mu_{k}^{*\top}\Sigma^{*-1/2}v\right|.$
bing2024lineardiscriminantregularizedregression give that for any $t\geq 0$, with probability at least $1-4e^{-t}$,
$v^{\top}\Sigma^{*-1/2}(\bar{\Sigma}-\Sigma^{*})\Sigma^{*-1/2}v\lesssim\sqrt{t\over n}+{t\over n}$
as well as
$\sum_{k=1}^{L}(\widetilde{\pi}_{k}-\pi_{k})v^{\top}\Sigma^{*-1/2}\mu_{k}\mu_{k}^{\top}\Sigma^{*-1/2}v$ $\lesssim\Delta_{\max}\left(\sqrt{t\over n}+{t\over n}\right).$ (D.44)
By inspecting the proof of bing2024lineardiscriminantregularizedregression, one can also deduce that
$\left|\sum_{k=1}^{L}\widetilde{\pi}_{k}v^{\top}\Sigma^{*-1/2}(\widetilde{\mu}_{k}-\mu_{k}^{*})\mu_{k}^{*\top}\Sigma^{*-1/2}v\right|\lesssim\|M^{\top}\Sigma^{*-1/2}v\|_{\infty}\sqrt{t\over n}\leq\sqrt{\Delta_{\max}}\sqrt{t\over n}$
holds with probability at least $1-2e^{-t}$. The proof is completed by collecting all three bounds, choosing $t=Cd\log(n)$ for some large constant $C>0$ and using $d\log n\leq n$. ∎
Appendix E Proofs of the sample level results in Section 4.2
E.1 Proof of Theorem 10
Proof.
We first prove Theorem 10 when $\widehat{\theta}^{(t)}$ satisfies (3.5) and
$L\left(1+{1\over\pi_{\min}^{*2}\Delta_{\min}^{2}}\right){\phi(\widehat{\theta}^{(t)})\over n}\leq c.$ (E.1)
Later we will argue that these conditions are met under either of the initialization requirements in Theorem 10.
Pick any $\theta$ satisfying (3.5) and rewrite $\omega=(\pi,M,J(\theta))\in\mathcal{B}_{d}(\omega^{*})$. For any $a\in[L]$ and $i\in\widehat{G}_{a}$, we have
$\gamma_{\ell}(X_{i};\omega)$ $\leq{1\over 1+\exp(N_{i}^{\top}\Sigma^{*1/2}J({\bm{e}}_{a}-{\bm{e}}_{\ell})+\delta_{a\ell}^{a}(\omega)+\log(\pi_{a}/\pi_{\ell}))}$ by (C.54)
$\leq{1\over 1+\exp(N_{i}^{\top}\Sigma^{*1/2}J({\bm{e}}_{a}-{\bm{e}}_{\ell})+2c\Delta_{a\ell}+\log(\pi_{a}^{*}/\pi_{\ell}^{*}))}$ by (C.2) & Lemma 20
$\leq{1\over 1+\exp(N_{i}^{\top}\Sigma^{*1/2}J({\bm{e}}_{a}-{\bm{e}}_{\ell})+c\Delta_{a\ell})}$ $\text{by }\Delta_{\min}\geq C\log(1/\pi_{\min}^{*})$
$\leq{1\over 1+\exp(N_{i}^{\top}\Sigma^{*1/2}J^{*}({\bm{e}}_{a}-{\bm{e}}_{\ell})+(c/2)\Delta_{a\ell})}$
$\quad+1\left\{\left|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|>(c/4)\Delta_{a\ell}\right\}$
$\quad+1\left\{\left|N_{i}^{\top}\Sigma^{*1/2}(J-\widetilde{J})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|>(c/4)\Delta_{a\ell}\right\}.$
Here we write $\widetilde{J}:=\widetilde{\Sigma}^{-1}\widetilde{M}$ with $\widetilde{M}=(\widetilde{\mu}_{1},\ldots,\widetilde{\mu}_{L})$ and $\widetilde{\Sigma}$ being the estimators that use the true labels $Y_{1},\ldots,Y_{n}$. Specifically, recalling $\widetilde{\pi}_{\ell}=n_{\ell}/n$ from (D.32), we have
[…]
By applying the above bound of $\gamma_{\ell}(X_{i};\theta)$ to $\widehat{\theta}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\widehat{\Sigma}^{(t)})$ with the corresponding $\widehat{\omega}^{(t)}=(\widehat{\pi}^{(t)},\widehat{M}^{(t)},J(\widehat{\theta}^{(t)}))\in\mathcal{B}_{d}(\omega^{*})$, we obtain
$\phi(\widehat{\theta}^{(t)})=\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}\gamma_{\ell}(X_{i};\widehat{\theta}^{(t)})\|\mu_{\ell}^{*}-\mu_{a}^{*}\|_{\Sigma^{*}}^{2}\leq\textrm{I}+\textrm{II}+\textrm{III}$ (E.2)
with
I $=\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}{1\over 1+\exp(N_{i}^{\top}\Sigma^{*1/2}J^{*}({\bm{e}}_{a}-{\bm{e}}_{\ell})+(c/2)\Delta_{a\ell})}\Delta_{a\ell}$
II $=\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}1\left\{\left|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|>(c/4)\Delta_{a\ell}\right\}\Delta_{a\ell}$
III $=\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}1\left\{\left|N_{i}^{\top}\Sigma^{*1/2}(J(\widehat{\theta}^{(t)})-\widetilde{J})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|>(c/4)\Delta_{a\ell}\right\}\Delta_{a\ell}.$
We proceed to bound each term in the sequel on the event $\mathcal{E}_{\pi}$, given in (D.2), intersecting with
$\mathcal{E}_{{\rm op}}$ $:=\bigcap_{\ell\in[L]}\left\{\left\|\sum_{i\in\widehat{G}_{\ell}}N_{i}N_{i}^{\top}\right\|_{\rm op}\lesssim d+n_{\ell}\right\}\bigcap\left\{\left\|\sum_{i=1}^{n}N_{i}N_{i}^{\top}\right\|_{\rm op}\lesssim n+d\right\},$ (E.3)
$\mathcal{E}_{{\rm op}}^{\prime}$ $:=\left\{\left\|{1\over n}\sum_{i=1}^{n}N_{i}N_{i}^{\top}-\mathbf{I}_{d}\right\|_{\rm op}\lesssim\sqrt{d+\log n\over n}\right\},$ (E.4)
$\mathcal{E}_{\ell_{2}}$ $:=\bigcap_{\ell\in[L]}\left\{\left\|{1\over n_{\ell}}\sum_{i\in\widehat{G}_{\ell}}N_{i}\right\|_{2}\lesssim\sqrt{d+\log n\over n_{\ell}}\right\}.$ (E.5)
According to Lemmas 32, 31 and 35, the intersection event holds with probability $1-e^{-n\pi_{\min}^{*}}-n^{-2}-n^{-1}\geq 1-2n^{-1}$ by $n\pi_{\min}^{*}\geq C\log n$. The event $\mathcal{E}_{\ell_{2}}$ ensures that
$\left\|\widetilde{\mu}_{\ell}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}={1\over n_{\ell}}\left\|\sum_{i\in\widehat{G}_{\ell}}N_{i}\right\|_{2}\lesssim\sqrt{d+\log n\over n_{\ell}},\qquad\forall~\ell\in[L],$ (E.6)
which, together with (C.2) and $Cd\log n\leq n\pi_{\min}^{*}$, further implies
$\|\widetilde{\mu}_{\ell}\|_{\Sigma^{*}}\leq\left\|\widetilde{\mu}_{\ell}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}+\|\mu_{\ell}^{*}\|_{\Sigma^{*}}\leq 2\sqrt{\Delta_{\max}}.$ (E.7)
Moreover, using the decomposition in (D.1.3) and (E.6), we find that on the event $\mathcal{E}_{\ell_{2}}\cap\mathcal{E}_{{\rm op}}^{\prime}$,
$d(\widetilde{\Sigma},\Sigma^{*})$ $\leq\left\|{1\over n}\sum_{i=1}^{n}N_{i}N_{i}^{\top}-\mathbf{I}_{d}\right\|_{\rm op}+\left\|\sum_{\ell=1}^{L}{n_{\ell}\over n}\Bigl({1\over n_{\ell}}\sum_{i\in\widehat{G}_{\ell}}N_{i}\Bigr)\Bigl({1\over n_{\ell}}\sum_{i\in\widehat{G}_{\ell}}N_{i}\Bigr)^{\top}\right\|_{\rm op}$
$\lesssim\sqrt{d+\log n\over n}+{L(d+\log n)\over n},$ (E.8)
which, by Weyl’s inequality and $Cd\log n\leq n\pi_{\min}^{*}\leq n/L$, implies that
$1/2\leq\lambda_{j}\bigl(\Sigma^{*-1/2}\widetilde{\Sigma}\Sigma^{*-1/2}\bigr)\leq 2,\qquad\forall~1\leq j\leq d.$ (E.9)
Bounding I.
By Eqs. C.56 and C.57, we have
$\mathbb{E}\left[{1\over 1+\exp(N_{i}^{\top}\Sigma^{*1/2}J^{*}({\bm{e}}_{a}-{\bm{e}}_{\ell})+(c/2)\Delta_{a\ell})}\right]\leq\exp(-c^{\prime}\Delta_{a\ell})$
which in conjunction with $\Delta_{\min}\geq C\log(L)$ implies
$\mathbb{E}[\textrm{I}]$ $\leq\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}\exp(-c^{\prime}\Delta_{a\ell}+\log(\Delta_{a\ell}))\leq n\exp(-c^{\prime\prime}\Delta_{\min}).$
By Markov’s inequality, we conclude with probability at least $1-\exp(-(c^{\prime\prime}/2)\Delta_{\min})$,
$\textrm{I}\leq\mathbb{E}[\textrm{I}]\exp((c^{\prime\prime}/2)\Delta_{\min})\leq n\exp(-(c^{\prime\prime}/2)\Delta_{\min}).$ (E.10)
Bounding II.
Lemma 27 states that with probability at least $1-4n^{-1}$,
$\max_{\ell\in[L]}\max_{a\in[L]\setminus\{\ell\}}\max_{i\in\widehat{G}_{a}}{1\over\Delta_{a\ell}}\left|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|\leq c/4,$
whence $\textrm{II}=0$.
Bounding III.
Note that
III $\leq\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}{16\over c^{2}\Delta_{a\ell}}\left[N_{i}^{\top}\Sigma^{*1/2}(J(\widehat{\theta}^{(t)})-\widetilde{J})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right]^{2}$
$\leq{16\over c^{2}}\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\left\|\sum_{i\in\widehat{G}_{a}}N_{i}N_{i}^{\top}\right\|_{\rm op}{1\over\Delta_{a\ell}}\left\|\Sigma^{*1/2}(J(\widehat{\theta}^{(t)})-\widetilde{J})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}^{2}$
$\leq{16\over c^{2}}\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}(d+n_{a}){1\over\Delta_{a\ell}}\left\|\Sigma^{*1/2}(J(\widehat{\theta}^{(t)})-\widetilde{J})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}^{2}$ $\text{by }\mathcal{E}_{{\rm op}}.$
Using $d(\widehat{\Sigma}^{(t)},\Sigma^{*})\leq 1/2$ implied by $\widehat{\theta}^{(t)}$ satisfying (3.5), reasoning as (C.2) gives
$\|\Sigma^{*1/2}(J(\widehat{\theta}^{(t)})-\widetilde{J})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$
$\leq\|\Sigma^{*1/2}\bm{\widehat{\Sigma}}^{(t)-1}\Sigma^{*1/2}\|_{\rm op}\left(\|\widehat{\mu}_{\ell}^{(t)}-\widetilde{\mu}_{\ell}\|_{\Sigma^{*}}+\|\widehat{\mu}_{a}^{(t)}-\widetilde{\mu}_{a}\|_{\Sigma^{*}}\right)$
$\quad+\|\Sigma^{*1/2}\bm{\widehat{\Sigma}}^{(t)-1}\Sigma^{*1/2}\|_{\rm op}\|(\bm{\widehat{\Sigma}}^{(t)}-\widetilde{\Sigma})\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{\Sigma^{*}}$
$\leq 2\|\widehat{\mu}_{\ell}^{(t)}-\widetilde{\mu}_{\ell}\|_{\Sigma^{*}}+2\|\widehat{\mu}_{a}^{(t)}-\widetilde{\mu}_{a}\|_{\Sigma^{*}}+2\|(\bm{\widehat{\Sigma}}^{(t)}-\widetilde{\Sigma})\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{\Sigma^{*}}.$
Note that $(\widehat{\pi}^{(t)},\widehat{M}^{(t)},\widehat{\Sigma}^{(t)})=(\widehat{\bm{\pi}}(\widehat{\theta}^{(t-1)}),\widehat{\mathbf{M}}(\widehat{\theta}^{(t-1)}),\bm{\widehat{\Sigma}}(\widehat{\theta}^{(t-1)}))$. Invoking Lemma 28 with $\theta=\widehat{\theta}^{(t-1)})$ gives
III $\lesssim\left({dL^{2}\log(q)\over n\pi_{\min}^{*2}\Delta_{\min}^{2}}+{L\phi(\widehat{\theta}^{(t-1)})\over n\pi_{\min}^{*2}\Delta_{\min}^{2}}+\exp(-2c^{\prime}\Delta_{\min})\right)\phi(\widehat{\theta}^{(t-1)})\hskip-56.9055pt$
$\quad+\left\{\left({\log n\over\Delta_{\min}}+1\right)\left({dL^{2}\log(q)\over n}+L\exp(-2c\Delta_{\min})\right)+{L\phi(\widehat{\theta}^{(t-1)})\over n}\right\}\phi(\widehat{\theta}^{(t-1)})$
$\leq{1\over 2}\phi(\widehat{\theta}^{(t-1)})$
where the last step is due to $\Delta_{\min}\geq C\log(1/\pi_{\min}^{*}+\log n)$, condition (E.1) and the initialization requirement in (4.11). Collecting the bounds for I, II and III yields that: with probability at least $1-4n^{-1}-\exp(-c\Delta_{\min})$,
$\phi(\widehat{\theta}^{(t)})~\leq~n\exp(-c\Delta_{\min})+{1\over 2}\phi(\widehat{\theta}^{(t-1)}).$ (E.11)
Since this together with $\Delta_{\min}\geq C\log(1/\pi_{\min}^{*})$ further ensures $\phi(\widehat{\theta}^{(t)})$ satisfying (E.1), by induction and noting that the high probability bound in (E.11) holds uniformly for all $\widehat{\theta}^{(t-1)}$, the claim is proved.
Finally, we note that when $\widehat{\theta}^{(t)}$ meets the requirement in (a), the above proof implies that with probability $1-4n^{-1}-\exp(-c\Delta_{\min})$,
$\phi(\widehat{\theta}^{(t)})\leq n\exp(-c\Delta_{\min})+{64nL\over c^{2}}\left(2c_{\mu}^{2}+c_{\Sigma}^{2}\right)$
so that (E.1) holds as
$L\left(1+{1\over\pi_{\min}^{*2}\Delta_{\min}^{2}}\right){\phi(\widehat{\theta}^{(t)})\over n}\leq\left(1+{1\over\pi_{\min}^{*2}\Delta_{\min}^{2}}\right)\left(L\exp(-c\Delta_{\min})+{64L^{2}\over c^{2}}\left(2c_{\mu}^{2}+c_{\Sigma}^{2}\right)\right)\leq c^{\prime}$
by $\Delta_{\min}\geq C\log(1/\pi_{\min}^{*})$ and (a). On the other hand, when $\widehat{\theta}^{(t)}$ meets the requirement in (b), then Eqs. E.27, E.28 and E.29 and (b) ensure that $\widehat{\theta}^{(t)}$ also satisfies (3.5). The proof is complete. ∎
E.2 Proof of Theorem 9
Proof.
The proof follows from that of Theorem 10 and the only difference is in bounding III where the term $\|(\bm{\widehat{\Sigma}}^{(t)}-\widetilde{\Sigma})\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{\Sigma^{*}}=0$ in this case.
Regarding the initialization requirement, for $\widehat{\theta}^{(0)}=(\widehat{\pi}^{(0)},\widehat{M}^{(0)},\Sigma^{*})$, if $\phi(\widehat{\theta}^{(0)})/n\leq c\pi_{\min}^{*}\Delta_{\min}(1+\pi_{\min}^{*}\Delta_{\min})$, then (E.27) and (E.28) ensure that $\widehat{\theta}^{(0)}$ satisfies (3.5) with $c_{\Sigma}=0$. On the other hand, if $\widehat{\theta}^{(0)}=(\widehat{\pi}^{(0)},\widehat{M}^{(0)},\Sigma^{*})$ satisfies (3.5), then the proof reveals
${\phi(\widehat{\theta}^{(0)})\over n}\leq{CL\over\Delta_{\min}}d(\widehat{M}^{(0)},M^{*})^{2}\leq CLc_{\mu}^{2}$
so that $c_{\mu}\leq\min\{(\sqrt{2}-1)/2,c\pi_{\min}^{*}\Delta_{\min}/L\}$ suffices. ∎
E.3 Technical lemmas used in the proof of Theorem 10
We first state two useful facts.
Fact 1.
The following equality holds for any $\theta$:
$\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}}\left(1-\gamma_{\ell}(X_{i};\theta)\right)$ $=\sum_{\ell=1}^{L}\sum_{i\notin\widehat{G}_{\ell}}\gamma_{\ell}(X_{i};\theta),$
$\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}}\left(1-\gamma_{\ell}(X_{i};\theta)\right)N_{i}$ $=\sum_{\ell=1}^{L}\sum_{i\notin\widehat{G}_{\ell}}\gamma_{\ell}(X_{i};\theta)N_{i}.$
Proof.
By definition,
$\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}}\left(1-\gamma_{\ell}(X_{i};\theta)\right)$ $=\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}}\sum_{k\in[L]\setminus\{\ell\}}\gamma_{k}(X_{i};\theta)$
$=\sum_{\ell=1}^{L}\left(\sum_{k=1}^{L}\sum_{i\in\widehat{G}_{\ell}}\gamma_{k}(X_{i};\theta)-\sum_{i\in\widehat{G}_{\ell}}\gamma_{\ell}(X_{i};\theta)\right)$
$=\sum_{k=1}^{L}\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}}\gamma_{k}(X_{i};\theta)-\sum_{k=1}^{L}\sum_{i\in\widehat{G}_{k}}\gamma_{k}(X_{i};\theta)$
$=\sum_{k=1}^{L}\sum_{\ell\in[L]\setminus\{k\}}\sum_{i\in\widehat{G}_{\ell}}\gamma_{k}(X_{i};\theta)$
$=\sum_{k=1}^{L}\sum_{i\in\widehat{G}_{k}}\gamma_{k}(X_{i};\theta).$
Similar argument can be used to prove the second statement. ∎
There are a few useful inequalities about $\phi(\theta)$ which follow directly by definitions of $\phi(\theta)$ in (4.9) and $\Delta_{\ell k}$ in (C.1) as well as 1.
Fact 2.
For any $\theta$, we has the following inequalities.
$\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\theta))=\sum_{\ell=1}^{L}\sum_{i\notin\widehat{G}_{a}}\gamma_{\ell}(X_{i};\theta)$ $~\leq~{\phi(\theta)\over\Delta_{\min}},$ (E.12)
$\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\sum_{i\in\widehat{G}_{a}}\gamma_{\ell}(X_{i};\theta)\sqrt{\Delta_{a\ell}}$ $~\leq~{\phi(\theta)\over\sqrt{\Delta_{\min}}}.$ (E.13)
Lemma 27.
Grant $\Delta_{\min}\gtrsim 1$, $n\geq dL^{2}$ and $Cd\log n\leq n\pi_{\min}^{*}$ for some large constant $C>0$. The following holds with probability at least $1-4n^{-1}$,
$\max_{\ell\in[L]}\max_{a\in[L]\setminus\{\ell\}}\max_{i\in\widehat{G}_{a}}~{1\over\Delta_{a\ell}}\left|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|\leq c/4.$
Proof.
We work on the event $\mathcal{E}_{\ell_{2}}\cap\mathcal{E}_{2}\cap\mathcal{E}_{\pi}\cap\mathcal{E}_{{\rm op}}^{\prime}$ and first condition on $Y_{1},\ldots,Y_{n}$. Fix any $a\neq\ell$ and $i\in\widehat{G}_{a}$. We use the leave-one-out technique by decomposing
$\left|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|\leq\left|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}-\widetilde{J}^{(i)})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|+\left|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}^{(i)}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|$
where $\widetilde{J}^{(i)}:=\widetilde{\Sigma}^{(i)-1}\widetilde{M}^{(i)}$ with $\widetilde{\Sigma}^{(i)}$ and $\widetilde{M}^{(i)}$ being the same as $\widetilde{\Sigma}$ and $\widetilde{M}$ except leaving out the sample $X_{i}$. Specifically,
$\widetilde{\mu}_{a}^{(i)}={1\over n_{a}-1}\sum_{j\in\widehat{G}_{a}\setminus\{i\}}X_{j},\qquad\widetilde{\mu}_{b}^{(i)}=\widetilde{\mu}_{b},\quad\forall~b\in[L]\setminus\{a\},$ (E.14)
as well as
$\widetilde{\Sigma}^{(i)}={1\over n}\left[\sum_{b\in[L]\setminus\{a\}}\sum_{i\in\widehat{G}_{b}}(X_{i}-\widetilde{\mu}_{b})(X_{i}-\widetilde{\mu}_{b})^{\top}+\sum_{j\in\widehat{G}_{a}\setminus\{i\}}(X_{j}-\widetilde{\mu}_{a}^{(i)})(X_{j}-\widetilde{\mu}_{a}^{(i)})^{\top}\right].$ (E.15)
Note that $\mathcal{E}_{\pi}$ implies $n_{a}\geq n\pi_{a}^{*}/2\geq C\log n$. By construction, the independence between $N_{i}$ and $\widetilde{J}^{(i)}$ yields that for all $t\geq 0$,
$\mathbb{P}\left\{\left|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}^{(i)}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|\geq t\left\|\Sigma^{*1/2}(\widetilde{J}^{(i)}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}\right\}\leq 2\exp\left(-t^{2}/2\right).$ (E.16)
On the other hand, by $\mathcal{E}_{2}$ in (D.5), we have
$\left|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}-\widetilde{J}^{(i)})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|$ $\leq\left\|\Sigma^{*1/2}(\widetilde{J}-\widetilde{J}^{(i)})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}\max_{i\in[n]}\|N_{i}\|_{2}$
$\leq 4\left\|\Sigma^{*1/2}(\widetilde{J}-\widetilde{J}^{(i)})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}\sqrt{d+\log n}.$ (E.17)
We proceed to control $\|\Sigma^{*1/2}(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$ and $\|\Sigma^{*1/2}(\widetilde{J}-\widetilde{J}^{(i)})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$, which by triangle inequality also gives a bound for $\|\Sigma^{*1/2}(\widetilde{J}^{(i)}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$.
Recall $d(\widetilde{\Sigma},\Sigma^{*})\leq 1/2$ from (E.9). Reasoning as (C.2) and using Eqs. E.6 and E.1 together with $n\geq dL^{2}$ give
$\|\Sigma^{*1/2}(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$ $\leq 2\|\widetilde{\mu}_{a}-\mu_{a}^{*}\|_{\Sigma^{*}}+2\|\widetilde{\mu}_{\ell}-\mu_{\ell}^{*}\|_{\Sigma^{*}}+2d(\widetilde{\Sigma},\Sigma^{*})\sqrt{\Delta_{a\ell}}$
$\lesssim\sqrt{d+\log n\over n_{a}\wedge n_{\ell}}+\sqrt{d+\log n\over n}\sqrt{\Delta_{a\ell}}.$ (E.18)
In conjunction with $n\pi_{\min}^{*}\Delta_{\min}\geq Cd\log n$ and $\mathcal{E}_{\pi}$, one also has
$\|\Sigma^{*1/2}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}\leq\|\Sigma^{*1/2}(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}+\|\Sigma^{*1/2}J^{*}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}\leq 2\sqrt{\Delta_{a\ell}}.$ (E.19)
Regarding $\|\Sigma^{*1/2}(\widetilde{J}-\widetilde{J}^{(i)})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$, by the definition in (E.15), one has
$d(\widetilde{\Sigma},\widetilde{\Sigma}^{(i)})$ $={n_{a}\over n(n_{a}-1)}\left\|\Sigma^{*-1/2}(X_{i}-\widetilde{\mu}_{a})(X_{i}-\widetilde{\mu}_{a})^{\top}\Sigma^{*-1/2}\right\|_{\rm op}\hskip-28.45274pt$
$\leq{2n_{a}\over n(n_{a}-1)}\left(\|N_{i}\|_{2}^{2}+\|\widetilde{\mu}_{a}-\mu_{a}^{*}\|_{\Sigma^{*}}^{2}\right)$
$\lesssim{d\log n\over n}$ by (E.6) and $\mathcal{E}_{2}$
whence
$1/2\leq\lambda_{j}\Bigl(\Sigma^{*-1/2}\widetilde{\Sigma}^{(i)}\Sigma^{*-1/2}\Bigr)\leq 2,\qquad\forall~1\leq j\leq d.$ (E.20)
Meanwhile, by (E.14), we obtain
$\|\widetilde{\mu}_{a}-\widetilde{\mu}_{a}^{(i)}-\widetilde{\mu}_{\ell}+\widetilde{\mu}_{\ell}^{(i)}\|_{\Sigma^{*}}$ $={1\over n_{a}-1}\|X_{i}-\widetilde{\mu}_{a}\|_{\Sigma^{*}}$
$\leq{1\over n_{a}-1}\left(\left\|N_{i}\right\|_{2}+\|\widetilde{\mu}_{a}-\mu_{a}^{*}\|_{\Sigma^{*}}\right)$
$\leq{2\over n_{a}-1}\sqrt{d+\log n}$ (E.21)
where the last step uses $\mathcal{E}_{2}$ and (E.6). We thus have, by (E.20) and reasoning as (C.2),
$\|\Sigma^{*1/2}(\widetilde{J}-\widetilde{J}^{(i)})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$ $\leq 2\|\widetilde{\mu}_{a}-\widetilde{\mu}_{a}^{(i)}\|_{\Sigma^{*}}+2\left\|(\widetilde{\Sigma}^{(i)}-\widetilde{\Sigma})\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{\Sigma^{*}}$
$\leq{8\over n_{a}}\sqrt{d+\log n}+2\left\|(\widetilde{\Sigma}^{(i)}-\widetilde{\Sigma})\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{\Sigma^{*}}$ (E.22)
where the last step uses Eq. E.21. Also observe that
$\left\|(\widetilde{\Sigma}^{(i)}-\widetilde{\Sigma})\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{\Sigma^{*}}$
$={n_{a}\over n(n_{a}-1)}\left\|(X_{i}-\widetilde{\mu}_{a})(X_{i}-\widetilde{\mu}_{a})^{\top}\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{\Sigma^{*}}$
$\leq{2\over n}\left\|N_{i}N_{i}^{\top}\Sigma^{*1/2}\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}+{2\over n}\left\|N_{i}(\mu_{a}^{*}-\widetilde{\mu}_{a})^{\top}\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}$
$\quad+{2\over n}\left\|(\mu_{a}^{*}-\widetilde{\mu}_{a})N_{i}^{\top}\Sigma^{*1/2}\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{\Sigma^{*}}$
$\quad+{2\over n}\left\|(\mu_{a}^{*}-\widetilde{\mu}_{a})(\mu_{a}^{*}-\widetilde{\mu}_{a})^{\top}\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{\Sigma^{*}}\hskip-28.45274pt$
$\lesssim{\sqrt{d+\log n}\over n}|N_{i}^{\top}\Sigma^{*1/2}\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})|+{d\log n\over n\sqrt{n_{a}}}\|\Sigma^{*1/2}\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}.$ (E.23)
The last step is due to (E.6) and $\mathcal{E}_{2}$. To bound the remaining term $N_{i}^{\top}\Sigma^{*1/2}\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})$, note that $\widetilde{J}^{(i)}$ is independent of $X_{i}$ hence of $N_{i}$. We thus have
$\mathbb{P}\left(\left|N_{i}^{\top}\Sigma^{*1/2}\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|\geq 2\sqrt{2\log n}\left\|\widetilde{\Sigma}^{*1/2}\widetilde{J}^{(i)}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}\right)\leq{2/n^{4}},$
which, by taking the union bounds, holds for all $i\in[N]$ and $a\neq\ell$, with probability $1-n^{-2}$. Combining with (E.22) and (E.23) yields
$\|\Sigma^{*1/2}(\widetilde{J}-\widetilde{J}^{(i)})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$ $\lesssim{\sqrt{d+\log n}\over n_{a}}+\left({\sqrt{d}\log n\over n}+{d\log n\over n\sqrt{n_{a}}}\right)\|\Sigma^{*1/2}(\widetilde{J}-\widetilde{J}^{(i)})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$
$\quad+\left({\sqrt{d}\log n\over n}+{d\log n\over n\sqrt{n_{a}}}\right)\|\Sigma^{*1/2}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}.$
By $n\geq d\log n$ and (E.19), we conclude
$\|\Sigma^{*1/2}(\widetilde{J}-\widetilde{J}^{(i)})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}\lesssim{\sqrt{d+\log n}\over n_{a}}+\left({\sqrt{d}\log n\over n}+{d\log n\over n\sqrt{n_{a}}}\right)\sqrt{\Delta_{a\ell}}.$ (E.24)
In conjunction with (E.3), we also have
$\|\Sigma^{*1/2}(\widetilde{J}^{(i)}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}\lesssim\sqrt{d+\log n\over n_{a}\wedge n_{\ell}}+\sqrt{d+\log n\over n}\sqrt{\Delta_{a\ell}}.$ (E.25)
Finally, plugging (E.24) and (E.25) into Eqs. E.16 and E.3, choosing $t=C\sqrt{\log n}$, taking the union bounds over $i\in[n]$, $a,\ell\in[L]$ and invoking $\mathcal{E}_{\pi}$ give
${1\over\Delta_{a\ell}}\left|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|$
$~\lesssim~{d+\log n\over n\pi_{a}^{*}\Delta_{a\ell}}+{\sqrt{d(d+\log n)}\log n\over n\sqrt{\Delta_{a\ell}}}+{d\log n\over n}\sqrt{d+\log n\over n\pi_{a}^{*}\Delta_{a\ell}}$
$\quad+{\sqrt{\log n}\over\Delta_{a\ell}}\left({\sqrt{d+\log n}\over n\pi_{a}^{*}}+\sqrt{d+\log n\over n(\pi_{a}^{*}\wedge\pi_{\ell}^{*})}+\sqrt{d+\log n\over n}\sqrt{\Delta_{a\ell}}\right)$ (E.26)
with probability at least $1-4n^{-1}$. The proof is complete. ∎
Lemma 28.
Assume $n\pi_{\min}^{*}\geq Cd\log n$, $n\geq dL^{2}$ and $\Delta_{\min}\geq C\log(1/\pi_{\min}^{*})$. On the event $\mathcal{E}_{\pi}\cap\mathcal{E}_{{\rm op}}\cap\mathcal{E}_{\ell_{2}}\cap\mathcal{E}_{{\rm op}}^{\prime}$, the following holds for any $\theta$ satisfying $\phi(\theta)/n\leq\Delta_{\min}\pi_{\min}^{*}(1+\Delta_{\min}\pi_{\min}^{*})$. For all $\ell\in[L]$:
[…]
$\|\widehat{\bm{\mu}}_{\ell}(\theta)-\widetilde{\mu}_{\ell}\|_{\Sigma^{*}}$ $~\lesssim~{1\over\pi_{\ell}^{*}\sqrt{\Delta_{\min}}}\left(\sqrt{\phi(\theta)\over n}+{\phi(\theta)\over n}\right)$ (E.28)
and
$d(\bm{\widehat{\Sigma}}(\theta),\widetilde{\Sigma})~\lesssim~\sqrt{\Delta_{\max}\over\Delta_{\min}}\sqrt{\phi(\theta)\over n}+{\Delta_{\max}\over\Delta_{\min}}{\phi(\theta)\over n}.$ (E.29)
Furthermore, if additionally $\theta$ satisfies (3.5) and $dL\log(q)\leq n\pi_{\min}^{*}\Delta_{\min}$, then (E.28) can be improved to: with probability at least $1-n^{-1}$, for all $\ell\in[L]$,
$\|\widehat{\bm{\mu}}_{\ell}(\theta)-\widetilde{\mu}_{\ell}\|_{\Sigma^{*}}$ $~\lesssim~{1\over\pi_{\ell}^{*}\sqrt{\Delta_{\min}}}\left(\sqrt{\phi(\theta)\over n}\sqrt{dL\log(q)\over n}+{\phi(\theta)\over n}\right)+\sqrt{\phi(\theta)\over n}\exp(-c^{\prime}\Delta_{\min}).$ (E.30)
In this case, with the same probability, we also have
$d(\bm{\widehat{\Sigma}}(\theta),\widetilde{\Sigma})~\lesssim~{\phi(\theta)\over n}+{dL\log(q)\over n}+\exp(-c^{\prime}\Delta_{\min})$ (E.31)
and
${1\over\sqrt{\Delta_{a\ell}}}\|(\bm{\widehat{\Sigma}}(\theta)-\widetilde{\Sigma})\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{\Sigma^{*}}$
$\qquad\qquad\lesssim~\sqrt{{\log n\over\Delta_{\min}}+1}\left(\sqrt{dL\log(q)\over n}+\exp(-c^{\prime}\Delta_{\min})\right)\sqrt{\phi(\theta)\over n}+{\phi(\theta)\over n}.$ (E.32)
Proof.
To prove (E.27), it suffices to note
[…]
$\leq{1\over n}\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\theta))+{1\over n}\sum_{\ell=1}^{L}\sum_{i\notin G_{\ell}}\gamma_{\ell}(X_{i};\theta)$
$\leq{2\phi(\theta)\over n\Delta_{\min}}$
where the last inequality uses Eq. E.12. By $8\phi(\theta)/n\leq\pi_{\min}^{*}\Delta_{\min}$ and $\mathcal{E}_{\pi}$, we thus have
$1/4\leq\widehat{\bm{\pi}}_{\ell}(\theta)/\pi_{\ell}^{*}\leq 5/4$ (E.33)
To show (E.28), start with the decomposition
$\widehat{\bm{\mu}}_{\ell}(\theta)-\widetilde{\mu}_{\ell}$ $={1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}\sum_{i=1}^{n}\gamma_{\ell}(X_{i};\theta)(X_{i}-\mu_{\ell}^{*})-{1\over n\widetilde{\pi}_{\ell}}\sum_{i\in\widehat{G}_{\ell}}(X_{i}-\mu_{\ell}^{*})$
$={\widetilde{\pi}_{\ell}-\widehat{\bm{\pi}}_{\ell}(\theta)\over\widehat{\bm{\pi}}_{\ell}(\theta)}{1\over n_{\ell}}\sum_{i\in\widehat{G}_{\ell}}N_{i}-{1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\theta))N_{i}$
$\qquad+{1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}\sum_{j\notin\widehat{G}_{\ell}}\gamma_{\ell}(X_{j};\theta)(X_{j}-\mu_{\ell}^{*}).$ (E.34)
First, using (E.27), $\mathcal{E}_{\ell_{2}}\cap\mathcal{E}_{\pi}$ and (E.33) gives
${|\widetilde{\pi}_{\ell}-\widehat{\bm{\pi}}_{\ell}(\theta)|\over\widehat{\bm{\pi}}_{\ell}(\theta)}\left\|{1\over n_{\ell}}\sum_{i\in\widehat{G}_{\ell}}N_{i}\right\|_{\Sigma^{*}}\lesssim{\phi(\theta)\over n\Delta_{\min}\widehat{\bm{\pi}}_{\ell}(\theta)}\sqrt{d+\log n\over n_{\ell}}\lesssim{\phi(\theta)\over n\Delta_{\min}\pi_{\ell}^{*}}\sqrt{d+\log n\over n\pi_{\ell}^{*}}.$
Second, an application of the Cauchy-Schwarz inequality yields
${1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}\left\|\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\theta))N_{i}\right\|_{\Sigma^{*}}$
$\leq{1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}\sqrt{\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\theta))}\sqrt{\|\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\theta))N_{i}N_{i}^{\top}\|_{\rm op}}\hskip-71.13188pt$
$\leq{1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}\sqrt{\phi(\theta)\over\Delta_{\min}}\sqrt{\|\sum_{i\in\widehat{G}_{\ell}}N_{i}N_{i}^{\top}\|_{\rm op}}$ by (E.12)
$\lesssim\sqrt{\phi(\theta)\over n\pi_{\ell}^{*}\Delta_{\min}}\sqrt{{d\over n\pi_{\ell}^{*}}+1}$ $\text{by }\mathcal{E}_{{\rm op}}\cap\mathcal{E}_{\pi}.$ (E.35)
Regarding the remaining term,
${1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}\left\|\sum_{j\notin\widehat{G}_{\ell}}\gamma_{\ell}(X_{j};\theta)(X_{j}-\mu_{\ell}^{*})\right\|_{\Sigma^{*}}$
$\leq{1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}\left\|\sum_{j\notin\widehat{G}_{\ell}}\gamma_{\ell}(X_{j};\theta)N_{j}^{*}\right\|_{2}+{1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}\sum_{a\in[L]\setminus\{\ell\}}\sum_{j\in\widehat{G}_{a}}\gamma_{\ell}(X_{j};\theta)\left\|\mu_{a}^{*}-\mu_{\ell}^{*}\right\|_{\Sigma^{*}}$
$\leq{1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}\sqrt{\sum_{j\notin\widehat{G}_{\ell}}\gamma_{\ell}(X_{j};\theta)}\sqrt{\|\sum_{j\notin\widehat{G}_{\ell}}\gamma_{\ell}(X_{j};\theta)N_{j}N_{j}^{\top}\|_{\rm op}}+{1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}{\phi(\theta)\over\sqrt{\Delta_{\min}}}$ by (E.13)
$\lesssim{1\over\pi_{\ell}^{*}}\sqrt{\phi(\theta)\over n\Delta_{\min}}\sqrt{n+d\over n}+{\phi(\theta)\over n\pi_{\ell}^{*}\sqrt{\Delta_{\min}}}.$ (E.36)
Combining the last three displays together with $n\pi_{\min}^{*}\Delta_{\min}\geq Cd\log n$ gives (E.28), which, in conjunction with $\phi(\theta)/n\leq c\pi_{\min}^{*2}\Delta_{\min}^{2}$, (E.7) and (C.2), further ensures
$\|\widehat{\bm{\mu}}_{\ell}(\theta)\|_{\Sigma^{*}}\leq\|\widehat{\bm{\mu}}_{\ell}(\theta)-\widetilde{\mu}_{\ell}\|_{\Sigma^{*}}+\|\widetilde{\mu}_{\ell}-\mu_{\ell}^{*}\|_{\Sigma^{*}}+\|\mu_{\ell}^{*}\|_{\Sigma^{*}}\leq 3\sqrt{\Delta_{\max}}.$ (E.37)
To prove (E.29), observe from (2.11) that
$\bm{\widehat{\Sigma}}(\theta)-\widetilde{\Sigma}=\widehat{\mathbf{M}}(\theta){\rm diag}(\widehat{\bm{\pi}}(\theta))\widehat{\mathbf{M}}(\theta)^{\top}-\widetilde{M}{\rm diag}(\widetilde{\pi})\widetilde{M}^{\top}$.
Adding and subtracting terms gives
[…]
$\quad+\left\|\sum_{\ell=1}^{L}(\widehat{\bm{\pi}}_{\ell}(\theta)-\widetilde{\pi}_{\ell})\Sigma^{*-1/2}\widetilde{\mu}_{\ell}\widetilde{\mu}_{\ell}^{\top}\Sigma^{*-1/2}\right\|_{\rm op}.$
By $\mathcal{E}_{\pi}$ and Eqs. E.28, E.37 and E.7, the first term on the right-hand-side is bounded by
$\sum_{\ell=1}^{L}\widehat{\bm{\pi}}_{\ell}(\theta)\left\|\widehat{\bm{\mu}}_{\ell}(\theta)-\widetilde{\mu}_{\ell}\right\|_{\Sigma^{*}}\max_{\ell\in[L]}\|\widehat{\bm{\mu}}_{\ell}(\theta)+\widetilde{\mu}_{\ell}\|_{\Sigma^{*}}$ $\lesssim\sqrt{\Delta_{\max}\over\Delta_{\min}}\left(\sqrt{\phi(\theta)\over n}+{\phi(\theta)\over n}\right).$ (E.38)
On the other hand, since (E.27) and (E.37) imply
$\left\|\sum_{\ell=1}^{L}(\widehat{\bm{\pi}}_{\ell}(\theta)-\widetilde{\pi}_{\ell})\Sigma^{*-1/2}\widetilde{\mu}_{\ell}\widetilde{\mu}_{\ell}^{\top}\Sigma^{*-1/2}\right\|_{\rm op}\lesssim{\Delta_{\max}\over\Delta_{\min}}{\phi(\theta)\over n},$
the proof of (E.29) is complete.
To prove the improved bounds, we start with (E.30) and refine the bounds of $\|\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\theta))N_{i}N_{i}^{\top}\|_{\rm op}$ and $\|\sum_{j\notin\widehat{G}_{\ell}}\gamma_{\ell}(X_{j};\theta)N_{j}N_{j}^{\top}\|_{\rm op}$ in (E.35) and (E.36), respectively. We use the arguments similar as the proof of Lemma 24 and recall $S(\omega)$ in (D.30) for $\omega=\omega(\theta)\in\mathcal{B}_{d}(\omega^{*})$. The latter is ensured as $\theta$ satisfies (3.5). By repeating the similar arguments therein,
$\Bigl\|\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\theta))N_{i}N_{i}^{\top}\Bigr\|_{\rm op}$
$\leq~\Bigl\|\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}\cap S(\omega)}N_{i}N_{i}^{\top}\Bigr\|_{\rm op}+\Bigl\|\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}\cap[S(\omega)]^{c}}(1-\gamma_{\ell}(X_{i};\theta))N_{i}N_{i}^{\top}\Bigr\|_{\rm op}$
[…]
Invoking Lemma 32, $d\leq n$, (D.38) and $\mathcal{E}_{\pi}$ yields that: with probability $1-n^{-dL}-e^{-n}$,
$\left\|\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\theta))N_{i}N_{i}^{\top}\right\|_{\rm op}$ $\lesssim dL\log(q)+{nL\over\pi_{\min}^{*}}\exp(-c\Delta_{\min}).$ (E.39)
By 1, the same bound also holds for $\|\sum_{\ell=1}^{L}\sum_{j\notin\widehat{G}_{\ell}}\gamma_{\ell}(X_{j};\theta)N_{j}N_{j}^{\top}\|_{\rm op}$. Plugging the above two bounds into (E.35) and (E.36) and using $\Delta_{\min}\geq C\log(1/\pi_{\min}^{*})$ gives
${1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}\left\|\sum_{i\in\widehat{G}_{\ell}}(1-\gamma_{\ell}(X_{i};\theta))N_{i}\right\|_{\Sigma^{*}}+{1\over n\widehat{\bm{\pi}}_{\ell}(\theta)}\left\|\sum_{j\notin\widehat{G}_{\ell}}\gamma_{\ell}(X_{j};\theta)(X_{j}-\mu_{\ell}^{*})\right\|_{\Sigma^{*}}$
$\lesssim\sqrt{\phi(\theta)\over n\pi_{\ell}^{*}\Delta_{\min}}\sqrt{dL\log(q)\over n\pi_{\ell}^{*}}+\sqrt{\phi(\theta)\over n}\exp(-c^{\prime}\Delta_{\min})+{\phi(\theta)\over n\pi_{\ell}^{*}\sqrt{\Delta_{\min}}},$
which implies the improved rate in (E.30).
To prove (E.31), we have the decomposition
$\bm{\widehat{\Sigma}}(\theta)-\widetilde{\Sigma}$ $=\textrm{I}+\textrm{II}+\textrm{III}$ (E.40)
where
I $={1\over n}\sum_{\ell=1}^{L}\sum_{j\notin\widehat{G}_{\ell}}\gamma_{\ell}(X_{j};\theta)(X_{j}-\mu_{\ell}^{*})(X_{j}-\mu_{\ell}^{*})^{\top}$
II $=-{1\over n}\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}}\left(1-\gamma_{\ell}(X_{i};\theta)\right)\Sigma^{*1/2}N_{i}N_{i}^{\top}\Sigma^{*1/2}$
III $=\sum_{\ell=1}^{L}\widehat{\bm{\pi}}_{\ell}(\theta)(\widehat{\bm{\mu}}_{\ell}(\theta)-\mu_{\ell}^{*})(\widehat{\bm{\mu}}_{\ell}(\theta)-\mu_{\ell}^{*})^{\top}-\sum_{\ell=1}^{L}\widetilde{\pi}_{\ell}(\widetilde{\mu}_{\ell}-\mu_{\ell}^{*})(\widetilde{\mu}_{\ell}-\mu_{\ell}^{*})^{\top}.$
The term $\|\Sigma^{*-1/2}\textrm{II}\Sigma^{*-1/2}\|_{\rm op}$ is bounded in (E.39) while, by Cauchy-Schwarz inequality,
$\|\Sigma^{*-1/2}~\textrm{I}~\Sigma^{*-1/2}\|_{\rm op}$ $\leq\left\|{2\over n}\sum_{\ell=1}^{L}\sum_{j\notin\widehat{G}_{\ell}}\gamma_{\ell}(X_{j};\theta)N_{j}N_{j}^{\top}\right\|_{\rm op}+{2\over n}\sum_{\ell=1}^{L}\sum_{a\in[L]\setminus\{\ell\}}\sum_{j\in\widehat{G}_{a}}\gamma_{\ell}(X_{j};\theta)\Delta_{a\ell}$
$=2\|\Sigma^{*-1/2}\textrm{II}\Sigma^{*-1/2}\|_{\rm op}+{2\phi(\theta)\over n}$
using 1 and (4.9). Regarding III, by adding and subtracting terms,
$\left\|\Sigma^{*-1/2}~\textrm{III}~\Sigma^{*-1/2}\right\|_{\rm op}$ $\leq\left\|\sum_{\ell=1}^{L}\widehat{\bm{\pi}}_{\ell}(\theta)\Sigma^{*-1/2}(\widehat{\bm{\mu}}_{\ell}(\theta)-\widetilde{\mu}_{\ell})(\widehat{\bm{\mu}}_{\ell}(\theta)-\widetilde{\mu}_{\ell})^{\top}\Sigma^{*-1/2}\right\|_{\rm op}$
$\quad+2\left\|\sum_{\ell=1}^{L}\widehat{\bm{\pi}}_{\ell}(\theta)\Sigma^{*-1/2}(\widehat{\bm{\mu}}_{\ell}(\theta)-\widetilde{\mu}_{\ell})(\widetilde{\mu}_{\ell}-\mu_{\ell}^{*})^{\top}\Sigma^{*-1/2}\right\|_{\rm op}$
$\quad+\left\|\sum_{\ell=1}^{L}(\widehat{\bm{\pi}}_{\ell}(\theta)-\widetilde{\pi}_{\ell})\Sigma^{*-1/2}(\widetilde{\mu}_{\ell}-\mu_{\ell}^{*})(\widetilde{\mu}_{\ell}-\mu_{\ell}^{*})^{\top}\Sigma^{*-1/2}\right\|_{\rm op}$
$\leq\sum_{\ell=1}^{L}\widehat{\bm{\pi}}_{\ell}(\theta)\left\|\widehat{\bm{\mu}}_{\ell}(\theta)-\widetilde{\mu}_{\ell}\right\|_{\Sigma^{*}}\max_{\ell\in[L]}\left(\|\widehat{\bm{\mu}}_{\ell}(\theta)-\widetilde{\mu}_{\ell}\|_{\Sigma^{*}}+2\|\widetilde{\mu}_{\ell}-\mu_{\ell}^{*}\|_{\Sigma^{*}}\right)$
$\quad+\sum_{\ell=1}^{L}\left|\widehat{\bm{\pi}}_{\ell}(\theta)-\widetilde{\pi}_{\ell}\right|\max_{\ell\in[L]}\|\widetilde{\mu}_{\ell}-\mu_{\ell}^{*}\|_{\Sigma^{*}}^{2}.$
By invoking the improved bound in Eq. E.30 and using Eqs. E.6 and E.27 together with the event $\mathcal{E}_{\ell_{2}}\cap\mathcal{E}_{\pi}$, after some algebra, we obtain that
$\left\|\Sigma^{*-1/2}~\textrm{III}~\Sigma^{*-1/2}\right\|_{\rm op}$
$~\lesssim\left({dL\log(q)\over n\pi_{\min}^{*}\Delta_{\min}}+\sqrt{d+\log n\over n\pi_{\min}^{*}\Delta_{\min}}+\exp(-2c^{\prime}\Delta_{\min})+{\phi(\theta)\over n\pi_{\min}^{*}\Delta_{\min}}\right){\phi(\theta)\over n}$
$\quad+\left(\sqrt{dL\log(q)\over n}+\exp(-c^{\prime}\Delta_{\min})\right)\sqrt{d+\log n\over n\pi_{\min}^{*}}\sqrt{\phi(\theta)\over n}.$ (E.41)
Combining with (E.39) gives
[…]
$\qquad+\sqrt{dL\log(q)\over n}\sqrt{d+\log n\over n\pi_{\min}^{*}}\sqrt{\phi(\theta)\over n}+{dL\log(q)\over n}+\exp(-c^{\prime}\Delta_{\min})$
which, under $\phi(\theta)+dL\log(q)\leq n\pi_{\min}^{*}\Delta_{\min}$ and $d\log n\leq n\pi_{\min}^{*}$, simplifies to (E.31).
Finally, we prove (28) by using the decomposition (E.40) and 1
$\|(\bm{\widehat{\Sigma}}(\theta)-\widetilde{\Sigma})\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{\Sigma^{*}}$ $\leq{1\over n}\|(\textrm{I}+\textrm{II}+\textrm{III})\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{\Sigma^{*}}$
$\leq{1\over n}\left\|\sum_{k=1}^{L}\sum_{i\notin\widehat{G}_{b}}\gamma_{k}(X_{i};\theta)(X_{i}-\mu_{k}^{*})(X_{i}-\mu_{k}^{*})^{\top}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{\Sigma^{*}}$
$\quad+{1\over n}\left\|\sum_{k=1}^{L}\sum_{i\in\widehat{G}_{k}}(1-\gamma_{k}(X_{i};\theta))N_{i}N_{i}^{\top}\Sigma^{*1/2}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2}$
$\quad+{1\over n}\left\|\Sigma^{*-1/2}~\textrm{III}~\Sigma^{*-1/2}\right\|_{\rm op}\|\Sigma^{*1/2}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}.$ (E.42)
Note that the first term is bounded by $R_{1}+R_{2}+R_{3}+R_{4}$ with
$R_{1}$ $={1\over n}\left\|\sum_{k=1}^{L}\sum_{i\notin\widehat{G}_{k}}\gamma_{k}(X_{i};\theta)N_{i}N_{i}^{\top}\Sigma^{*1/2}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2},$
$R_{2}$ $={1\over n}\left\|\sum_{k=1}^{L}\sum_{b\in[L]\setminus\{k\}}\sum_{i\in\widehat{G}_{b}}\gamma_{k}(X_{i};\theta)N_{i}(\mu_{b}^{*}-\mu_{k}^{*})^{\top}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{2},$
$R_{3}$ $={1\over n}\left\|\sum_{k=1}^{L}\sum_{b\in[L]\setminus\{k\}}\sum_{i\in\widehat{G}_{b}}\gamma_{k}(X_{i};\theta)(\mu_{b}^{*}-\mu_{k}^{*})N_{i}^{\top}\Sigma^{*1/2}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right\|_{\Sigma^{*}},$
$R_{4}$ $={1\over n}\sum_{k=1}^{L}\sum_{b\in[L]\setminus\{k\}}\sum_{i\in\widehat{G}_{b}}\gamma_{k}(X_{i};\theta)\Delta_{bk}\|\Sigma^{*1/2}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}.$
By Cauchy-Schwarz inequality,
$R_{1}$ $\leq{1\over n}\left\|\sum_{k=1}^{L}\sum_{i\notin\widehat{G}_{k}}\gamma_{k}(X_{i};\theta)N_{i}N_{i}^{\top}\right\|_{\rm op}^{1/2}\left(\sum_{k=1}^{L}\sum_{i\notin\widehat{G}_{k}}\gamma_{k}(X_{i};\theta)(N_{i}^{\top}\Sigma^{*1/2}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell}))^{2}\right)^{1/2}$
$\lesssim\left(\sqrt{dL\log(q)\over n}+\exp(-c\Delta_{\min})\right)\sqrt{\phi(\theta)\over n\Delta_{\min}}\max_{i\in[n]}\left|N_{i}^{\top}\Sigma^{*1/2}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|$
with probability at least $1-n^{-dL}-e^{-n}-n^{-2}$. The second step is due to (E.39) & (E.12). Furthermore, using $N_{i}^{\top}\Sigma^{*1/2}J^{*}({\bm{e}}_{a}-{\bm{e}}_{\ell})\sim\mathcal{N}(0,\Delta_{a\ell})$ and (E.26) gives
$\max_{i\in[n]}\left|N_{i}^{\top}\Sigma^{*1/2}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\right|$ $\leq\max_{i\in[n]}\left(|N_{i}^{\top}\Sigma^{*1/2}J^{*}({\bm{e}}_{a}-{\bm{e}}_{\ell})|+|N_{i}^{\top}\Sigma^{*1/2}(\widetilde{J}-J^{*})({\bm{e}}_{a}-{\bm{e}}_{\ell})|\right)$
$\lesssim\sqrt{\Delta_{a\ell}+\log n}$
with probability at least $1-4n^{-1}$. By 1, the same bound of $R_{1}$ also holds for the second term in (E.3).
For $R_{2}$, by similar arguments as well as (E.19), one has
$R_{2}$ $\leq{1\over n}\Bigl\|\sum_{k=1}^{L}\sum_{i\notin\widehat{G}_{k}}\gamma_{k}(X_{i};\theta)N_{i}N_{i}^{\top}\Bigr\|_{\rm op}^{1/2}\Bigl(\sum_{k=1}^{L}\sum_{b\in[L]\setminus\{k\}}\sum_{i\in\widehat{G}_{b}}\gamma_{k}(X_{i};\theta)\Delta_{bk}\Bigr)^{1/2}\|\Sigma^{*1/2}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$
$\lesssim\left(\sqrt{dL\log(q)\over n}+\exp(-c\Delta_{\min})\right)\sqrt{\phi(\theta)\over n}\sqrt{\Delta_{a\ell}}$
with probability at least $1-n^{-dL}-e^{-n}$. It is easy to see that the same bound also holds for $R_{3}$. Regarding $R_{4}$, observe that
[…]
Finally, invoking (E.3) bounds the term $n^{-1}\|\Sigma^{*-1/2}~\textrm{III}~\Sigma^{*-1/2}\|_{\rm op}\|\Sigma^{*1/2}\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{2}$ so that collecting the previous bounds yields
${1\over\sqrt{\Delta_{a\ell}}}\|(\bm{\widehat{\Sigma}}(\theta)-\widetilde{\Sigma})\widetilde{J}({\bm{e}}_{a}-{\bm{e}}_{\ell})\|_{\Sigma^{*}}$ $\lesssim\sqrt{{\log n\over\Delta_{\min}}+1}\left(\sqrt{dL\log(q)\over n}+\exp(-c\Delta_{\min})\right)\sqrt{\phi(\theta)\over n}$
$\quad+\left({dL\log(q)\over n\pi_{\min}^{*}\Delta_{\min}}+{\phi(\theta)\over n\pi_{\min}^{*}\Delta_{\min}}+1\right){\phi(\theta)\over n}$
thereby completing the proof of (28). ∎
E.4 Proof of Theorem 11
Proof.
From Theorem 10, by induction and (E.1), we know that for all $t\geq 1$,
$\phi(\widehat{\theta}^{(t)})$ $\leq n\exp(-c\Delta_{\min})\left(1+1/2+\cdots+1/2^{t-1}\right)+2^{-t}\phi(\widehat{\theta}^{(0)})$
$\leq 2n\exp(-c\Delta_{\min})+2^{-t}.$ (E.43)
Since $\widehat{\theta}$ satisfies (3.5), invoking (E.30) gives that with probability $1-n^{-1}-\exp(-c\Delta_{\min})$, for all $t\geq 1$,
$\|\widehat{\mu}_{\ell}^{(t+1)}-\widetilde{\mu}_{\ell}\|_{\Sigma^{*}}$ $=\|\widehat{\bm{\mu}}_{\ell}(\widehat{\theta}^{(t)})-\widetilde{\mu}_{\ell}\|_{\Sigma^{*}}$
$\lesssim{1\over\pi_{\ell}^{*}\sqrt{\Delta_{\min}}}\left(\sqrt{\phi(\widehat{\theta}^{(t)})\over n}\sqrt{dL\log(q)\over n}+{\phi(\widehat{\theta}^{(t)})\over n}\right)+\sqrt{\phi(\widehat{\theta}^{(t)})\over n}\exp(-c^{\prime}\Delta_{\min})$
$\lesssim\exp(-c\Delta_{\min})+{2^{-t}\over\pi_{\min}^{*}}.$
Together with (E.6), we conclude that with the same probability,
$\|\widehat{\mu}_{\ell}^{(t+1)}-\mu_{\ell}^{*}\|_{\Sigma^{*}}$ $\leq\|\widehat{\mu}_{\ell}^{(t+1)}-\mu_{\ell}^{*}\|_{\Sigma^{*}}+\|\widetilde{\mu}_{\ell}-\mu_{\ell}^{*}\|_{\Sigma^{*}}$
$\lesssim\sqrt{d+\log n\over n\pi_{\ell}^{*}}+\exp(-c\Delta_{\min}),\qquad\text{ for $t\geq C\log n$.}$
Similarly, using (E.27) and $\mathcal{E}_{\pi}$ yields: for $t\geq C\log n$,
$d(\widehat{\pi}^{(t+1)},\pi^{*})$ $\leq d(\widehat{\bm{\pi}}(\widehat{\theta}^{(t)}),\widetilde{\pi})+d(\widetilde{\pi},\pi^{*})\lesssim{\phi(\widehat{\theta}^{(t)})\over n\Delta_{\min}\pi_{\min}^{*}}+\sqrt{\log n\over n\pi_{\min}^{*}}\lesssim\exp(-c^{\prime}\Delta_{\min})+\sqrt{\log n\over n\pi_{\min}^{*}}.$
Finally, regarding $\bm{\widehat{\Sigma}}^{(t+1)}$, using (E.1) and (E.31) as well as $dL^{2}\log(q)\log n\leq n$ yields
$d(\bm{\widehat{\Sigma}}(\widehat{\theta}^{(t)}),\Sigma^{*})$~ $\lesssim~\sqrt{d+\log n\over n}+{\phi(\widehat{\theta}^{(t)})\over n}+{dL\log(q)\over n}+\exp(-c^{\prime}\Delta_{\min})$
$\lesssim~\sqrt{d+\log n\over n}+\exp(-c^{\prime}\Delta_{\min})$
for all $t\geq C\log n$. This completes the proof. ∎
E.5 Proof of Theorem 15
Proof.
We first prove for $\Delta_{\max}\geq\log n$. For any $\theta$ and any $\ell\in[L]$ with $i\in\widehat{G}_{\ell}$, we know
$\left\{g(X_{i};\theta)\neq\ell\right\}\subseteq\left\{\gamma_{\ell}(X_{i};\theta)<{1\over 2}\right\}=\left\{\sum_{k\in[L]\setminus\{\ell\}}\gamma_{k}(X_{i};\theta)\geq{1\over 2}\right\}.$
It follows that
$\ell(\theta)={1\over n}\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}}1\{g(X_{i};\theta)\neq\ell\}$ $\leq{2\over n}\sum_{\ell=1}^{L}\sum_{i\in\widehat{G}_{\ell}}\sum_{k\in[L]\setminus\{\ell\}}\gamma_{k}(X_{i};\theta)$
[…]
Using (E.4) gives that with probability $1-n^{-1}-\exp(-c\Delta_{\min})$, for all $t\geq 1$,
$\ell(\widehat{\theta}^{(t)})\leq{2\over n\Delta_{\min}}\left(2n\exp(-c\Delta_{\min})+2^{-t}\right).$
Since $\ell(\widehat{\theta}^{(t)})$ only takes values from $\{0,1/n,\ldots,1\}$, the claim thus follows.
For $\Delta_{\max}\leq\log n$, by repeating the same arguments in the proof of Theorem 10 except adding and subtracting $J^{*}$ directly instead of $\widetilde{J}^{*}$, one can obtain that, with probability at least $1-2n^{-1}-\exp(-c\Delta_{\min})$, for all $t\geq 1$,
$\phi(\widehat{\theta}^{(t)})\leq n\exp(-c\Delta_{\min})+CnL\left\{{1\over\Delta_{\min}}\left[d(\widehat{M}^{(t)},M^{*})\right]^{2}+\left[d(\widehat{\Sigma}^{(t)},\Sigma^{*})\right]^{2}\right\}.$
The claim then follows by invoking Theorem 8 and (E.44). ∎
Appendix F Proof of Theorem 13: the minimax lower bounds
We separate the proof into $d\geq L-1$ and d<L1d<L-1.
Proof of $d\geq L-1$.
It suffices to prove
$\inf_{\widehat{\theta}}\sup_{\theta\in\Theta(\alpha,\delta)}\mathbb{P}_{\theta}\left\{d(\widehat{M},M)\geq c_{1}\sqrt{d\over n\alpha}\right\}\geq(1+c_{0})/2,$ (F.1)
$\inf_{\widehat{\theta}}\sup_{\theta\in\Theta(\alpha,\delta)}\mathbb{P}_{\theta}\left\{d(\widehat{\Sigma},\Sigma)\geq c_{1}\sqrt{d\over n}\right\}\geq(1+c_{0})/2$ (F.2)
which jointly imply the claim. We first note that when $d<16$, the lower bounds of $c/\sqrt{n\alpha}$ and $c/\sqrt{n}$ can be easily proved by applying Le Cam’s two-point argument; see, Tsybakov09. We thus focus on $d\geq 16$ in the following.
We first prove (F.1) by constructing a set of hypotheses. Let $\theta^{(0)}=(\pi,\mu_{1},\mu_{2},\ldots,\mu_{L},\mathbf{I}_{d})$ with $\alpha=\pi_{1}=\min_{\ell\in[L]}\pi_{\ell}$ and
$\|\mu_{k}-\mu_{\ell}\|_{2}=\sqrt{2\delta},\quad\forall~k\neq\ell.$
Note that the existence of such $\mu_{\ell}$’s is guaranteed by $d\geq L-1$. For all $0\leq i\leq N$, let $\theta^{(i)}=(\pi,\mu_{1}+\varepsilon\beta^{(i)},\mu_{2},\ldots,\mu_{L},\mathbf{I}_{d})$ with $\beta^{(0)}=0_{d}$ and $\beta^{(i)}\in\{0,1\}^{d}$ for $i\geq 1$, which according to the Varshamov-Gilbert bound in Tsybakov09, satisfy
$\|\beta^{(i)}-\beta^{(j)}\|_{1}\geq d/8,\qquad\forall~0\leq i<j\leq N$ (F.3)
and $\log N\geq(d/8)\log 2$. For some constant $c\in(0,1/4)$ to be determined later, by choosing
$\varepsilon=\sqrt{c\log 2\over 8n\alpha}\leq(\sqrt{2}-1)\sqrt{\delta\over d},$ (F.4)
it is easy to see that $\theta^{(i)}\in\Theta(\alpha,\delta,A)$ for all $0\leq i\leq N$. Indeed, we note that for any $k,\ell\geq 2$ with $k\neq\ell$, $\|\mu_{k}-\mu_{\ell}\|_{2}^{2}=2\delta$ by construction, and
$\|\mu_{1}^{(i)}-\mu_{\ell}\|_{2}\geq\sqrt{2\delta}-\varepsilon\|\beta^{(i)}\|_{2}\geq\sqrt{2\delta}-\sqrt{cd\log 2\over 8n\alpha}\geq\sqrt{\delta},$ (F.5)
$\|\mu_{1}^{(i)}-\mu_{\ell}\|_{2}\leq\sqrt{2\delta}+\varepsilon\|\beta^{(i)}\|_{2}\leq\sqrt{2\delta}+\sqrt{cd\log 2\over 8n\alpha}\leq 2\sqrt{\delta}.$ (F.6)
We proceed to invoke Tsybakov09 to prove (F.1), which requires to verify
  1. (i)
    ${\rm KL}(\mathbb{P}_{\theta^{(i)}},\mathbb{P}_{\theta^{(0)}})\leq c^{\prime}\log N$ for some $0<c^{\prime}<1/8$;
  2. (ii)
    $d(M^{(i)},M^{(j)})=\|\mu_{1}^{(i)}-\mu_{1}^{(j)}\|_{2}\geq 2c_{1}\sqrt{d/n\alpha}$.
Here $\mathbb{P}_{\theta^{(i)}}$ is the distribution of $n$ i.i.d. samples from (1.1) under the parametrization $\theta^{(i)}$. For (i), we find that
${\rm KL}(\mathbb{P}_{\theta^{(i)}},\mathbb{P}_{\theta^{(0)}})$ $=n~{\rm KL}\left(\sum_{\ell=1}^{L}\pi_{\ell}\mathcal{N}_{d}(\mu_{\ell}^{(i)},\mathbf{I}_{d}),\sum_{\ell=1}^{L}\pi_{\ell}\mathcal{N}_{d}(\mu_{\ell}^{(0)},\mathbf{I}_{d})\right)$
$\leq n\pi_{1}~{\rm KL}\left(\mathcal{N}_{d}(\mu_{1}^{(i)},\mathbf{I}_{d}),\mathcal{N}_{d}(\mu_{1}^{(0)},\mathbf{I}_{d})\right)$
$={n\alpha\over 2}\varepsilon^{2}\|\beta^{(i)}\|_{2}^{2}.$
By $\|\beta^{(i)}\|_{2}^{2}\leq d$, $\log N\geq(d/8)\log 2$ and (F.4) with $c^{\prime}=c/2$, we obtain (i). Since (ii) follows as
[…]
invoking Tsybakov09 proves (F.1) for $c_{1}=\sqrt{c\log 2}/16$ and
${\sqrt{N}\over 1+\sqrt{N}}\left(1-c-\sqrt{c\over\log(N)}\right)\geq(1+c_{0})/2$
by using $d\geq 16$ and choosing $c$ sufficiently small.
To establish (F.2), it suffices to prove
$\inf_{\widehat{\theta}}\sup_{\theta\in\Theta(\alpha,\delta,A),\Sigma\preceq{3\over 2}}\mathbb{P}_{\theta}\left\{\|\widehat{\Sigma}-\Sigma\|_{\rm op}\geq{3c_{1}\over 2}\sqrt{d\over n}\right\}\geq(1+c_{0})/2.$ (F.7)
For simplicity, we assume $d$ is even as the same argument can be proved for odd $d$, with $d/2$ replaced by $(d+1)/2$. Consider the following hypotheses: $\theta^{(i)}=(\pi,\mu_{1},\ldots,\mu_{L},\Sigma^{(i)})$ with $\pi_{\min}\geq\alpha$, $\|\mu_{k}-\mu_{\ell}\|_{2}^{2}=3\delta/2$ for all $k\neq\ell$, and
$\Sigma^{(i)}=\mathbf{I}_{d}+{\tau\over\sqrt{nd}}\sum_{m=1}^{d/2}\beta_{m}^{(i)}B(m)$
with $\beta_{m}^{(i)}\in\{0,1\}$ for $i=0,1,\ldots,N$, $1\leq m\leq d/2$ and some $\tau>0$ to be chosen later. Here the matrix $B(m)=(B_{ij}(m))_{d\times d}$ is defined as: for $i,j\in[d]$,
$B_{ij}(m)=1\{i=m\text{ and }m+1\leq j\leq d,\text{ or }j=m\text{ and }m+1\leq i\leq d\}.$
Then by using the inequality $\|A\|_{\rm op}\leq\max_{\ell\in[d]}\|A_{\ell\cdot}\|_{1}$ for any symmetric matrix $A$ and Weyl’s inequality, we obtain
$\lambda_{1}(\Sigma^{(i)})\leq 1+{\tau\over\sqrt{nd}}\max_{\ell\in[d]}\sum_{m=1}^{d/2}\beta_{m}^{(i)}\|B(m)_{\ell\cdot}\|_{1}\leq 1+\tau\sqrt{d\over n}\leq 3/2$ (F.8)
provided that $\tau\leq 1/2$, and similarly, we can also deduce that $\lambda_{p}(\Sigma^{(i)})\geq 1/2$. We thus obtain $\min_{k\neq\ell}\|\mu_{k}-\mu_{\ell}\|_{\Sigma^{(i)}}^{2}\geq\min_{k\neq\ell}\|\mu_{k}-\mu_{\ell}\|_{2}^{2}/\lambda_{1}(\Sigma^{(i)})\geq\delta$ as well as $\max_{k,\ell}\|\mu_{k}-\mu_{\ell}\|_{\Sigma^{(i)}}^{2}\leq\max_{k,\ell}\|\mu_{k}-\mu_{\ell}\|_{2}^{2}/\lambda_{p}(\Sigma^{(i)})\leq 3\delta$. Therefore, $\theta^{(i)}\in\Theta(\alpha,\delta,A)$ for all $0\leq i\leq N$.
By using the Varshamov-Gilbert bound again, one has $\|\beta^{(i)}-\beta^{(j)}\|_{1}\geq d/16$ for any $i\neq j$ and $\log N\geq(d/16)\log 2$. This also gives
$\|\Sigma^{(i)}-\Sigma^{(j)}\|_{\rm op}^{2}$ $=\sup_{v\neq 0}{\|(\Sigma^{(i)}-\Sigma^{(j)})v\|_{2}^{2}\over\|v\|_{2}^{2}}$
$\geq{\tau^{2}\over nd}\sum_{m=1}^{d/2}|\beta_{m}^{(i)}-\beta_{m}^{(j)}|{\|B(m)v\|_{2}^{2}\over\|v\|_{2}^{2}}$ $\text{by taking }v_{i}=1\{d/2\leq i\leq d\}$
$={\tau^{2}\over nd}\|\beta^{(i)}-\beta^{(j)}\|_{1}{d\over 2}$
$\geq{\tau^{2}d\over 32n}.$ (F.9)
On the other hand, for any $1\leq i\leq N$, we have
${\rm KL}(\mathbb{P}_{\theta^{(i)}},\mathbb{P}_{\theta^{(0)}})$ $\leq n\sum_{\ell=1}^{L}\pi_{\ell}~{\rm KL}\left(\mathcal{N}_{d}(\mu_{\ell},\Sigma^{(i)}),\mathcal{N}_{d}(\mu_{\ell},\mathbf{I}_{d})\right)$
$={n\over 2}\left({\rm tr}(\Sigma^{(i)})-d-\log{\det(\Sigma^{(i)})}\right)$
$={n\over 2}\left[{\rm tr}(\Sigma^{(i)}-\mathbf{I}_{d})-\log\det(\mathbf{I}_{d}+\Sigma^{(i)}-\mathbf{I}_{d})\right].$
Recall from (F.8) that the eigenvalues of $\Sigma^{(i)}-\mathbf{I}_{d}$ are bounded within [1/2,1/2][-1/2,1/2]. Taylor expansion gives that
$\log\det(\mathbf{I}_{d}+\Sigma^{(i)}-\mathbf{I}_{d})\geq{\rm tr}(\Sigma^{(i)}-\mathbf{I}_{d})-2\|\Sigma^{(i)}-\mathbf{I}_{d}\|_{F}^{2}$
so that, by choosing $\tau^{2}\leq c\log 2/16$,
${\rm KL}(\mathbb{P}_{\theta^{(i)}},\mathbb{P}_{\theta^{(0)}})\leq n\|\Sigma^{(i)}-\mathbf{I}_{d}\|_{F}^{2}\leq n{\tau^{2}\over nd}\sum_{m=1}^{d/2}\beta_{m}^{(i)}2d\leq\tau^{2}d\leq c\log N.$
In view of (F), choosing $c$ sufficiently small, invoking Tsybakov09 proves (F.7) for $c_{1}=\tau/(12\sqrt{2})$ and $d\geq 16$. This completes the proof. ∎
Proof of d<L1d<L-1.
The proof follows the same arguments except that: for proving (F.1), due to d<L1d<L-1, one can only construct $\mu_{1},\ldots,\mu_{L}$ such that
$\sqrt{2\delta}\leq\min_{k\neq\ell}\|\mu_{k}-\mu_{\ell}\|_{2}\leq\max_{k,\ell}\|\mu_{k}-\mu_{\ell}\|_{2}\leq L^{1/d}\sqrt{2\delta}\leq(\sqrt{A}-1)\sqrt{\delta}.$
[…]
$\|\mu_{1}^{(i)}-\mu_{\ell}\|_{2}\leq(\sqrt{A}-1)\sqrt{\delta}+\varepsilon\|\beta^{(i)}\|_{2}\leq\sqrt{A\delta},$
so that $\theta^{(i)}\in\Theta_{\alpha,\delta,A}$ for all $0\leq i\leq N$. For proving (F.2), with the choice of $\mu_{\ell}$’s such that
$\sqrt{3\delta/2}\leq\min_{k\neq\ell}\|\mu_{k}-\mu_{\ell}\|_{2}\leq\max_{k,\ell}\|\mu_{k}-\mu_{\ell}\|_{2}\leq\sqrt{A\delta/2},$
one has that for any $k\neq\ell$,
$\|\mu_{k}-\mu_{\ell}\|_{\Sigma^{(i)}}^{2}\leq\|\mu_{k}-\mu_{\ell}\|^{2}/\lambda_{p}(\Sigma^{(i)})\leq A\delta$
and
$\|\mu_{k}-\mu_{\ell}\|_{\Sigma^{(i)}}^{2}\geq\|\mu_{k}-\mu_{\ell}\|^{2}/\lambda_{1}(\Sigma^{(i)})\geq\delta$
so that the constructed hypotheses belong to the specified parameter space. The rest of the proof uses the same argument as proving Theorem 13. ∎
Appendix G Auxiliary Lemma
The following lemmas are proved in [bing2023optimal].
Lemma 29.
Assume $\pi_{\min}^{*}\geq 2\log n/n$. Then the following holds with probability at least $1-2n^{-1}$ for all $k\in[L]$,
$|\widetilde{\pi}_{k}-\pi_{k}^{*}|<\sqrt{16\pi_{k}(1-\pi_{k}^{*})\log n\over n}.$
Furthermore, if $\pi_{\min}^{*}\geq C\log n/n$ for some sufficiently large constant $C$, then
$\mathbb{P}\left\{c\pi_{k}^{*}\leq\widetilde{\pi}_{k}\leq c^{\prime}\pi_{k}^{*}\right\}\geq 1-2n^{-1}.$
The following lemma states upper bounds of the quadratic form of a sub-Gaussian random vector [hsu2012tail].
Lemma 30.
Let $\xi\in\mathbb{R}^{d}$ be a subGaussian random vector with parameter $\gamma_{\xi}$. Then, for all symmetric positive semi-definite matrices $H$, and all $t\geq 0$,
$\mathbb{P}\left\{\xi^{\top}H\xi\geq\gamma_{\xi}^{2}\left(\sqrt{\mathrm{tr}(H)}+\sqrt{2t\|H\|_{\mathrm{op}}}\right)^{2}\right\}\leq e^{-t}.$
As an application of Lemma 30, we have the following result.
Lemma 31.
Let $N_{1},\dots,N_{n}$ be i.i.d. from $\mathcal{N}_{d}(0,\mathbf{I}_{d})$. Then for any $\delta\in(0,1)$,
$\mathbb{P}\left\{\left\|{1\over n}\sum_{i=1}^{n}N_{i}\right\|_{2}\geq\sqrt{d\over n}+\sqrt{2\log(1/\delta)\over n}\right\}\leq\delta.$
Proof.
The result follows from Lemma 30 by taking $\xi=n^{-1}\sum_{i=1}^{n}N_{i}$ with $\gamma_{\xi}={1/\sqrt{n}}$. ∎
The following lemma provides an upper bound on the operator norm of $\sum_{i\in S}N_{i}N_{i}^{\top}$ for any $S\in\{0,1\}^{m}$ with any $m\leq n$.
Lemma 32.
For any $m\in\mathbb{N}$ with $m\leq n$, with probability at least $1-e^{-m}$, the following holds for all $S\subseteq[n]$ with $|S|\leq m$
$\left\|\sum_{i\in S}N_{i}N_{i}^{\top}\right\|_{\rm op}\leq~2d+12m\log\left(en\over m\right).$
Proof.
Fix any $S$ and write $N_{S}$ as the $|S|\times d$ matrix containing $\{N_{i}\}_{i\in S}$. Let $\mathcal{N}_{1/2}$ be the epsilon net of $\mathbb{S}^{|S|}$. Then
$\left\|\sum_{i\in S}N_{i}N_{i}^{\top}\right\|_{\rm op}=\left\|N_{S}N_{S}^{\top}\right\|_{\rm op}=\sup_{v\in\mathbb{S}^{|S|}}v^{\top}N_{S}N_{S}^{\top}v\leq 2\max_{v\in\mathcal{N}_{1/2}}v^{\top}N_{S}N_{S}^{\top}v.$
Since $N_{S}^{\top}v\sim\mathcal{N}_{d}(0,\mathbf{I}_{d})$, applying Lemma 30 with $\xi=N_{S}^{\top}v$ and $H=\mathbf{I}_{d}$ gives
$\mathbb{P}\left\{v^{\top}N_{S}N_{S}^{\top}v\geq\left(\sqrt{d}+\sqrt{2t}\right)^{2}\right\}\leq e^{-t},\qquad\forall~t\geq 0.$
Note that $|\mathcal{N}_{1/2}|\leq 5^{|S|}$ and there are at most ${n\choose m}\leq(en/m)^{m}$ all possible $S\in\{0,1\}^{m}$. Taking the union bounds over $v\in\mathcal{N}_{1/2}$ and $S$ and choosing $t=3m\log(en/m)$ gies the result. ∎
The following lemmas state the well-known vector-valued and matrix-valued Bernstein inequalities. See, for instance, Minsker2017.
Lemma 33 (Vector-valued Bernstein inequality).
Let $X_{1},\ldots,X_{n}$ be independent random vectors with zero mean and $\max_{i\in[n]}\|X_{i}\|_{2}\leq U$ almost surely. Denote $\sigma^{2}:=\sum_{i=1}^{n}\mathbb{E}[\|X_{i}\|_{2}^{2}]$. Then for all $t\geq{1\over 6}(U+\sqrt{U^{2}+36\sigma^{2}})$,
$\mathbb{P}\left(\left\|\sum_{i=1}^{n}X_{i}\right\|_{2}>t\right)\leq 28\exp\left(-{t^{2}/2\over\sigma^{2}+Ut/3}\right).$
Lemma 34 (Matrix-valued Bernstein inequality).
Let $X_{1},\ldots,X_{n}\in\mathbb{R}^{d\times d}$ be independent, symmetric random matrices with zero mean and $\max_{i\in[n]}\|X_{i}\|_{\rm op}\leq U$ almost surely. Denote $\sigma^{2}:=\|\sum_{i=1}^{n}\mathbb{E}[X_{i}^{2}]\|_{\rm op}$. Then for all $t\geq{1\over 6}(U+\sqrt{U^{2}+36\sigma^{2}})$,
$\mathbb{P}\left(\left\|\sum_{i=1}^{n}X_{i}\right\|_{\rm op}>t\right)\leq 14\exp\left(-{t^{2}/2\over\sigma^{2}+Ut/3}+\log(d)\right).$
Another useful concentration inequality of the operator norm of the random matrices with i.i.d. sub-Gaussian rows is stated in the following lemma [bing2020prediction, Lemma 16]. This is an immediate result of vershynin_2012.
Lemma 35.
Let $G$ be $n$ by $d$ matrix whose rows are i.i.d. $\gamma$ sub-Gaussian random vectors with covariance matrix $\Sigma_{G}$. Then, for every $t\geq 0$, with probability at least $1-2e^{-ct^{2}}$,
$\left\|{1\over n}G^{\top}G-\Sigma_{G}\right\|_{{\rm op}}\leq\max\left\{\delta,\delta^{2}\right\}\left\|\Sigma_{G}\right\|_{{\rm op}},$
with $\delta=C\sqrt{d/n}+t/\sqrt{n}$ where $c,C$ are positive constants depending on $\gamma$.